### v3

# MUBen: Benchmarking the Uncertainty of Molecular Representation Models

Abstract
Large molecular representation models pre-trained on massive unlabeled data have shown great success in predicting molecular properties. However, these models may tend to overfit the fine-tuning data, resulting in over-confident predictions on test data that fall outside of the training distribution. To address this issue, uncertainty quantification (UQ) methods can be used to improve the models’ calibration of predictions. Although many UQ approaches exist, not all of them lead to improved performance. While some studies have included UQ to improve molecular pre-trained models, the process of selecting suitable backbone and UQ methods for reliable molecular uncertainty estimation remains underexplored. To address this gap, we present MUBen, which evaluates different UQ methods for state-of-the-art backbone molecular representation models to investigate their capabilities. By fine-tuning various backbones using different molecular descriptors as inputs with UQ methods from different categories, we critically assess the influence of architectural decisions and training strategies on property prediction and uncertainty estimation. Our study offers insights for selecting UQ for backbone models, which can facilitate research on uncertainty-critical applications in fields such as materials science and drug discovery.
1 Introduction
The task of molecular representation learning is pivotal in scientific domains and bears the potential to facilitate research in fields such as chemistry, biology, and materials science (Walters & Barzilay, 2021). Nonetheless, supervised training typically requires vast quantities of data, which are challenging to acquire due to the need for expensive laboratory experiments (Eyke et al., 2020). With the advent of self-supervised Transformers (Vaswani et al., 2017) pioneered by BERT (Devlin et al., 2019) and GPT (Radford et al., 2018), there has been a surge of interest in creating large-scale pre-trained molecular representation models from unlabeled datasets (Wang et al., 2019; You et al., 2020; Rong et al., 2020; Liu et al., 2022; Zhou et al., 2023; Stärk et al., 2022). Such models have demonstrated impressive representational capabilities, achieving state-of-the-art (SOTA) performance on a variety of molecular property prediction tasks (Wu et al., 2018).
However, in many applications, there is a pressing need for reliable predictions that are not only precise but also uncertainty-aware. The provision of calibrated uncertainty estimates allows us to distinguish “noisy” predictions to improve model robustness (Geifman & El-Yaniv, 2017), or estimate data distributions to enhance downstream tasks such as high throughput screening (Noh et al., 2020; Graff et al., 2021), activity cliff identification (van Tilborg et al., 2022), or wet-lab experimental design (Eyke et al., 2020; Soleimany et al., 2021). Unfortunately, large-scale models can easily overfit the fine-tuning data and exhibit misplaced confidence in their predictions (Guo et al., 2017; Kong et al., 2020). Several works have introduced various uncertainty quantification (UQ) methods in molecular property prediction (Ryu et al., 2019; Zhang & Lee, 2019; Scalia et al., 2020; Hwang et al., 2020; Mamun et al., 2020; Hie et al., 2020; Busk et al., 2021; Soleimany et al., 2021; Deshwal et al., 2021; Gruich et al., 2023; Wollschläger et al., 2023) and in neighboring research areas such as protein engineering (Greenman et al., 2022). For instance, Busk et al. (2021) extend a message passing neural network (MPNN, Gilmer et al., 2017) with post-hoc recalibration (Guo et al., 2017) to address the overconfidence in molecular property prediction; Soleimany et al. (2021) apply evidential message passing networks (Sensoy et al., 2018; Amini et al., 2020) to quantitative structure-activity relationship regression tasks; Deshwal et al. (2021) utilized Bayesian optimization to conduct reliable predictions of Nanoporous material properties and to direct experiments.
Despite their contributions, these studies exhibit several constraints: 1) the variety of uncertainty estimation methods considered is limited with some effective strategies being overlooked; 2) each research focuses on a limited range of properties such as quantum mechanics, failing to explore a broader spectrum of tasks; and 3) none embraces the power of recent pre-trained backbone models, which have shown remarkable performance in prediction but may lead to different results in UQ. To the best of our knowledge, a comprehensive evaluation of UQ methods applied to pre-trained molecular representation models is currently lacking and warrants further investigation.
Figure 1: MUBen overview with datasets, backbones, UQ methods and metrics enumerated.
We present MUBen, a benchmark designed to assess the performance of UQ methods applied to the molecular representation models for property prediction on various metrics, as illustrated in Figure 1. It encompasses UQ methods from different categories, including deterministic prediction, Bayesian neural networks, post-hoc calibration, and ensembles (§ 4.2), on top of a set of molecular representation (backbone) models, each relies on a molecular descriptor from a distinct perspective (§ 4.1). MUBen delivers intriguing results and insights that may guide the selection of backbone model and/or UQ methods in practice (§ 5). We structure our code, available at https://github.com/Yinghao-Li/MUBen, to be user-friendly, easily transferrable, and extendable, with the hope that this work will promote the future development of UQ methods, pre-trained models, or applications within the domains of materials science and drug discovery.
2 Related Works
2.1 Pre-Training Molecular Representation Models
Existing pre-trained models for molecular representation learning can be distinguished into three categories based on the type of molecular descriptors.
  • The first category includes ChemBERTa (Chithrananda et al., 2020), SMILES-BERT (Wang et al., 2019), Molformer (Ross et al., 2022), GPT-MolBERTa (Balaji et al., 2023), etc., which leverage string-based input descriptors such as SMILES (Weininger, 1988), SELFIES (Krenn et al., 2020), Group SELFIES (Cheng et al., 2023), or SAFE (Noutahi et al., 2023). Such models are generally based on the Transformer encoder and pre-trained with masked language modeling (MLM) objectives (Devlin et al., 2019), which masks out a portion of tokens and trains the model to predict them using the context information.
  • Given that molecules can be represented as 2D topological graphs with atoms as vertices and chemical bonds as edges, the second category of pre-training strategies employs techniques such as context prediction (Hu et al., 2020; Rong et al., 2020), contrastive learning (Sun et al., 2020; You et al., 2020; Wang et al., 2022), and replaced component detection (Li et al., 2021; Kim et al., 2022) to learn molecular representation models with the awareness of the 2D atomic relationships.
  • Recent advancements in the development of 3D atomistic models have garnered significant attention, particularly due to their capability to capture the intricate three-dimensional nature of molecular structures. The body of research in this area is primarily divided into two categories: invariant models and equivariant models. Invariant models, including SchNet (Schütt et al., 2017), DimeNet (Klicpera et al., 2020), SphereNet (Liu et al., 2021), GemNet (Gasteiger et al., 2021), and ClofNet (Du et al., 2022), utilize scalar quantities that remain unchanged under Euclidean transformations such as rotations, translations, and reflections. These quantities, which include pairwise distances, angles, and dihedrals, serve as features for constructing message-passing neural networks. On the other hand, equivariant models like TFN (Thomas et al., 2018), Equiformer (Liao & Smidt, 2023), NequIP (Batzner et al., 2022), EGNN (Satorras et al., 2021), and LEFTNet (Du et al., 2023), implement message-passing mechanisms designed to maintain equivariance—ensuring that operations such as linear transformations or tensor products do not compromise the model’s ability to handle geometric changes. Further insights and comprehensive analyses of these models are discussed in (Duval et al., 2023). Additionally, the pre-training objectives of these models emphasize the importance of 3D molecular properties. For instance, GEM (Fang et al., 2022) introduces bond angles and lengths as edge attributes to enhance 3D molecular insights, while Uni-Mol (Zhou et al., 2023) employs a unique combination of a 3D position denoising loss and a masked atom prediction loss, facilitating the effective learning of 3D spatial representations of molecules.
Beyond property prediction, pre-trained molecular representation models are leveraged for additional tasks, such as chemical design (Taylor et al., 2022; Ghugare et al., 2023; Li et al., 2023a). For instance, Ghugare et al. (2023) delve into the potential of merging text-based generative models with reinforcement learning to synthesize molecules with targeted properties. This trend further emphasize the value of MUBen.
2.2 Uncertainty Quantification
Uncertainty estimation plays a crucial role in stimulating deep neural networks to acknowledge what they don’t know, which has been applied to various domains such as object detection (Chen et al., 2021; He et al., 2019) and decision making (Amiri et al., 2020; Kong et al., 2022). UQ on neural networks generally takes two directions: Bayesian and non-Bayesian. Bayesian neural networks (BNNs) quantify predictive uncertainty by imposing probability distributions over model parameters instead of using point estimates. While BNNs provide a principled way of UQ, exact inference of parameter posteriors is often intractable. To circumvent this difficulty, existing works adopt Monte Carlo dropout (Gal & Ghahramani, 2016), variational inference (Blundell et al., 2015; Kingma et al., 2015) and Markov chain Monte Carlo (MCMC) (Welling & Teh, 2011; Maddox et al., 2019) for posterior approximation. Along the other direction, non-Bayesian methods estimate uncertainty by assembling multiple networks (Lakshminarayanan et al., 2017; Fort et al., 2019), adopting specialized output distribution (Sensoy et al., 2018; Amini et al., 2020; Soleimany et al., 2021) or training scheme (Lin et al., 2017; Mukhoti et al., 2020), or calibrating model outputs through additional network layers during validation and test (Platt et al., 1999; Guo et al., 2017; Song et al., 2019; Angelopoulos & Bates, 2021; Angelopoulos et al., 2021). Recently, there have been several UQ tools (Ghosh et al., 2021; Chung et al., 2021; Detommaso et al., 2023) and benchmarks from the general domain (Nado et al., 2021; Schmähling et al., 2023), biomedical engineering (Mehrtens et al., 2023; Greenman et al., 2022) and materials science (Hirschfeld et al., 2020; Hwang et al., 2020; Scalia et al., 2020; Varivoda et al., 2023; Wollschläger et al., 2023). MUBen also targets the last domain but covers a more diverse suite of tasks, stronger pre-trained molecular representation models, and a more comprehensive set of UQ methods. We hope MUBen can serve as a standard benchmark to faithfully evaluate the ability of UQ approaches.
3 Problem Setup
Let $\bm{x}$ denote a descriptor of a molecule, such as SMILES or a 2D topological graph. We define $y$ as the label with domain $y\in\{1,\ldots,K\}$ for $K$-class classification or $y\in$ for regression. $\bm{\theta}$ denotes all parameters of the molecular representation network and the task-specific layer. We assume that a training dataset $\mathcal{D}$ consists of $N$ i.i.d. samples $\mathcal{D}=\{(\bm{x}_{n},y_{n})\}_{n=1}^{N}$. Our goal is to fine-tune the parameters $\bm{\theta}$ on $\mathcal{D}$ and learn a calibrated predictive probability model $p_{\bm{\theta}}(y|\bm{x}_{n})$. The UQ metrics are briefly illustrated below with more details being provided in appendix D.
  • Negative Log-Likelihood (NLL) is often utilized to assess the quality of model uncertainty on holdout sets for both classification and regression tasks. Despite being a proper scoring rule as per Gneiting’s framework (Gneiting & Raftery, 2007), certain limitations such as overemphasizing tail probabilities (Quiñonero-Candela et al., 2006) make it inadequate to serve as the only UQ metric.
  • Brier Score (BS, Brier, 1950) is the mean squared error between the predicted probability and the ground-truth labels, which serves as another proper scoring rule for classification tasks.
    $\mathrm{BS}=\frac{1}{N}\sum_{n=1}^{N}\sum_{y=1}^{K}(p_{\bm{\theta}}(y|\bm{x}_{n})-\bm{1}_{y=y_{n}})^{2},$
    where $\bm{1}$ is the indicator function, i.e., $\bm{1}_{a=b}=1$ if $a=b$; otherwhise $\bm{1}_{a=b}=0$.
  • Calibration Errors measure the correspondence between predicted probabilities and empirical accuracy. For classification, we use Expected Calibration Error (ECE, Naeini et al., 2015), which first divides the predicted probabilities $\{p_{\bm{\theta}}(y_{n}\mid\bm{x}_{n})\}_{n=1}^{N}$ into $S$ bins $B_{s}=\{n\in\{1,\dots,N\}\mid p_{\bm{\theta}}(y_{n}\mid\bm{x}_{n})\in(\rho_{s},\rho_{s+1}]\}$ with $s\in\{1,\dots,S\}$, each with size $N_{s}$, and then compute the $L_{1}$ loss between the accuracy and predicted probabilities within each bin:
    $\operatorname{ECE}=\sum_{s=1}^{S}\frac{N_{s}}{N}\left|\frac{1}{N_{s}}\sum_{n\in B_{s}}\bm{1}_{y_{n}=\hat{y}_{n}}-\frac{1}{N_{s}}\sum_{n\in B_{s}}p_{\bm{\theta}}(\hat{y}_{n}\mid\bm{x}_{n})\right|,$
    where $\hat{y}_{n}=\operatorname*{arg\,max}_{y}p_{\bm{\theta}}(y\mid\bm{x}_{n})$ is the $n$-th prediction.
    Similarly, for regression, we use Regression Calibration Error (CE, Kuleshov et al., 2018), which calculates the true frequency of the predicted points lying in each confidence interval against the predicted fraction of points in that interval:
    $\operatorname{CE}=\frac{1}{S}\sum_{s=1}^{S}\left(\frac{s}{S}-\frac{1}{N}\left|\{n\in\{1,\dots,N\}\mid F_{\bm{\theta}}(y_{n};\bm{x}_{n})\leq\frac{s}{S}\}\right|\right)^{2},$
    where $\frac{s}{S}$ is the expected quantile and $F_{\bm{\theta}}(y_{n};\bm{x}_{n})$ represents the predicted quantile value of $y_{n}$, e.g., Gaussian CDF $\Phi(y_{n};\hat{y}_{n},\hat{\sigma}_{n})$ parameterized by the predicted mean $\hat{y}_{n}$ and variance $\hat{\sigma}_{n}$ if we assume the labels are subject to the Gaussian distribution.
4 Experiment Setup
4.1 Backbone Models
Molecular descriptors, such as fingerprints, SMILES strings, and graphs, package the structural data of a molecule into a format recognizable for computational algorithms. When paired with model architectures carrying divergent inductive biases, these descriptors can offer varied benefits depending on the specific task. Therefore, we select $4$ primary backbone models that accept distinct descriptors as input. We also include $2$ supplementary backbones to cover other aspects such as model architecture and pre-training objectives.
For primary backbone models, we select various pre-trained Transformer encoder-based models, each is SOTA or close to SOTA for their respective input format: 1) ChemBERTa (Chithrananda et al., 2020; Ahmad et al., 2022), which accepts SMILES strings as input; 2) GROVER (Rong et al., 2020), pre-trained with 2D molecular graphs via a Transformer bolstered by a dynamic message-passing graph neural network; and 3) Uni-Mol (Zhou et al., 2023), which incorporates 3D molecular conformations into its input, and pre-trains specialized Transformer architectures to encode the invariant spatial positions of atoms and represent the edges between atom pairs. We also implement a 4) fully connected deep neural network (DNN) that uses $200$-dimensional RDKit features which are regarded as fixed outputs of a hand-crafted backbone feature generator. This simple model aims to highlight the performance difference between heuristic feature generators and the automatic counterparts that draw upon self-supervised learning processes for knowledge acquisition. In addition, we have two other backbones: 5) TorchMD-NET, as introduced by Thölke & Fabritiis (2022) and pre-trained according to Zaidi et al. (2023). This model is designed to handle 3D inputs and leverages an equivariant Transformer architecture aimed at quantum mechanical properties. 6) GIN from Xu et al. (2019), which processes 2D molecular graphs. GIN is theoretically among the most powerful graph neural networks (GNNs) available, offering an additional randomly initialized backbone baseline with less sophisticated input features. Compared with the primary one, these models are constrained in pre-training: TorchMD-NET is exclusively pre-trained on quantum mechanical data, whereas GIN does not involve any pre-processed features. Please refer to appendix B for more details.
4.2 Uncertainty Quantification Methods
In MUBen, we choose popular UQ methods from different categories. We provide a high-level sketch here of each method and leave the detailed description to appendix C.
Deterministic Uncertainty Prediction
In standard practice, deep classification networks utilize SoftMax or Sigmoid outputs to distribute prediction weights among target classes. Such weights serve as an estimate of the model’s uncertainty. In regression tasks, rather than one single output value, models often parameterize an independent Gaussian distribution with predicted means and variances for each data point, and the variance magnitude acts as an estimate of the model’s uncertainty.
An alternate loss function for classification is the Focal Loss (Lin et al., 2017; Mukhoti et al., 2020), which minimizes a regularised KL divergence between the predicted values and the true labels. The regularization increases the entropy of the predicted distribution while the KL divergence is minimized, mitigating the model overconfidence and improving the uncertainty representation.
Bayesian Learning and Inference
BNNs estimate the probability distribution over the parameters of the network. Specifically, BNNs operate on the assumption that the network layer weights follow a specific distribution that is optimized through maximum a posteriori estimation during training. During inference, multiple network instances are randomly sampled from the learned distributions. Each instance then makes independent predictions, and the prediction distribution inherently captures the desired uncertainty information. Some notable methods that employ this approach include:
  • Bayes by Backprop (BBP, Blundell et al., 2015) first introduced Monte Carlo gradients, an extension of the Gaussian reparameterization trick (Kingma & Welling, 2014), to learn the posterior distribution of network weights directly through backpropagation.
  • Stochastic Gradient Langevin Dynamics (SGLD, Welling & Teh, 2011) applies Langevin dynamics to infuse noise into the stochastic gradient descent training process, thereby mitigating the overfitting of network parameters. The samples generated via Langevin dynamics can be used to form Monte Carlo estimates of posterior expectations during the inference stage.
  • MC Dropout (Gal & Ghahramani, 2016) demonstrates that applying dropout is equivalent to a Bayesian approximation of the deep Gaussian process (Damianou & Lawrence, 2013). Uncertainty is derived from an ensemble of multiple stochastic forward passes with dropout enabled.
  • SWA-Gaussian (SWAG, Maddox et al., 2019) provides an efficient way of estimating Gaussian posteriors over network weights by utilizing stochastic weight averaging (SWA, Izmailov et al., 2018) combined with low-rank & diagonal Gaussian covariance approximations.
Post-Hoc Calibration
Post-hoc calibration is proposed to address the over-confidence issue of deterministic classification models by adjusting the output logits after training. The most popular method is Temperature Scaling (Platt et al., 1999; Guo et al., 2017), which adds a learned scaling factor to the Sigmoid or SoftMax output activation to control their “spikiness”. Guo et al. (2017) shows that this simple step improves model calibration in general, yielding better UQ performance.
Deep Ensembles
The ensemble method has long been used in machine learning to improve model performance (Dietterich, 2000) but was first adopted to estimate uncertainty in (Lakshminarayanan et al., 2017). It trains a deterministic network multiple times with different random seeds and combines their predictions at inference. The ensembles can explore training modes thoroughly in the loss landscape and thus are robust to noisy and out-of-distribution examples (Fort et al., 2019).
4.3 Datasets
We carry out our experiments on MoleculeNet (Wu et al., 2018), a collection of widely utilized datasets covering molecular properties such as quantum mechanics, solubility, and toxicity. For classification, MUBen incorporates BBBP, ClinTox, Tox21, ToxCast, SIDER, BACE, HIV, and MUV with all labels being binary; the first $5$ concern physiological properties and the last $3$ biophysics. For regression, we select ESOL, FreeSolv, Lipophilicity, QM7, QM8, and QM9, of which the first $3$ contain physical chemistry properties and the rest fall into the quantum mechanics category.
In line with previous studies (Fang et al., 2022; Zhou et al., 2023), MUBen divides all datasets with scaffold splitting to minimize the impact of dataset randomness and, consequently, enhances the reliability of the evaluation. Moreover, scaffold splitting alienates the molecular features in each dataset split, inherently creating a challenging out-of-distribution (OOD) setup that better reflects the real-world scenario. For comparison, we also briefly report the results of random splitting. Please check appendix A for detailed dataset descriptions.
4.4 Training and Evaluation Protocols
We use Sigmoid output activation and binary cross-entropy (BCE) loss function on all classification models unless the UQ method specifies a different objective. For regression, we standardize the labels by mapping the training labels to a standard Gaussian distribution before model training and evaluation to mitigate the impact of label magnitude on multi-task datasets. During inference, the predicted values are converted back to their original distribution according to the label mean and variance from the training set before computing the metrics. For datasets with multiple prediction tasks such as QM9, a classification model features one output head for each task, while regression models have two—one for predicting mean $\hat{y}$ and the other for variance $\hat{\sigma}$, which is guaranteed to be positive using a SoftPlus activation (Lakshminarayanan et al., 2017). Regression models are trained with Gaussian NLL objective. We apply validation-based early stopping for parameter selection on all experiments unless the UQ method has other requirements.
Consistent with recommendations from Wu et al. (2018) and other previous works (Fang et al., 2022; Zhou et al., 2023), we report ROC-AUC (area under the receiver operating characteristic curve) as the metric for classification prediction and RMSE (root-mean-square error) and MAE (mean absolute error) for regression. In quantifying classification uncertainty, we use ECE, NLL, and Brier Score as introduced in § 3. For regression, we compute the Gaussian NLL and regression CE. We compute the metrics for each task individually before calculating their macro average. Each reported metric is the average of $3$ individual training-test runs with random seeds 00, $1$, and $2$. This differs from Deep Ensembles, which aggregate model predictions prior to metric calculation.
Table 1: Classification results. $\uparrow$” and $\downarrow$” imply that better performance is indicated by a larger or smaller value, respectively. Text in bold signifies the top-performing UQ method within each backbone model, and cells in blue indicate the best performance across all backbone-UQ combinations. “ROC” represents ROC-AUC. Deep Ensembles consistently outperforms other UQ methods for both property prediction and uncertainty quantification; MC Dropout and Temperature Scaling also consistently improve UQ performance.
Tox21(a) ToxCast(a) Average Ranking(b)
[…]
DNN-RDKit
Deterministic 0.7386 0.0417 0.2771 0.0779 0.6222 0.1168 0.4436 0.1397 25.75 17.25 24.13 22.88
Temperature 0.7386 0.0342 0.2723 0.0773 0.6220 0.1114 0.4882 0.1398 25.75 15.38 21.25 19.88
Focal Loss 0.7374 0.1058 0.3161 0.0871 0.6289 0.1264 0.4389 0.1396 25.88 24.38 22.38 24.00
MC Dropout 0.7376 0.0356 0.2727 0.0763 0.6248 0.1093 0.4319 0.1358 26.50 13.00 19.38 18.63
SWAG 0.7364 0.0438 0.2793 0.0790 0.6207 0.1175 0.4441 0.1400 26.38 18.50 25.63 23.50
BBP 0.7243 0.0422 0.2847 0.0814 0.6020 0.1443 0.4673 0.1510 22.75 19.13 19.38 21.38
SGLD 0.7257 0.1192 0.3455 0.0978 0.5319 0.3054 0.6685 0.2378 27.75 26.00 28.88 27.88
Ensembles 0.7540 0.0344 0.2648 0.0746 0.6486 0.0900 0.4008 0.1292 20.00 7.13 11.75 13.13
ChemBERTa
Deterministic 0.7542 0.0571 0.2962 0.0812 0.6554 0.1209 0.4313 0.1330 15.63 17.38 18.88 19.38
Temperature 0.7542 0.0424 0.2744 0.0792 0.6540 0.1067 0.4817 0.1313 15.88 12.00 13.88 15.38
Focal Loss 0.7523 0.0969 0.3052 0.0845 0.6442 0.1197 0.4243 0.1346 17.63 20.13 17.88 20.63
MC Dropout 0.7641 0.0423 0.2697 0.0744 0.6624 0.1069 0.4070 0.1276 12.50 10.75 10.63 10.00
SWAG 0.7538 0.0592 0.3008 0.0818 0.6556 0.1202 0.4305 0.1327 16.13 19.50 21.38 20.38
BBP 0.7433 0.0459 0.2765 0.0780 0.5814 0.1276 0.4545 0.1469 20.88 19.00 16.00 19.50
SGLD 0.7475 0.0504 0.2784 0.0795 0.5436 0.2238 0.5602 0.1881 21.13 19.88 19.13 18.63
Ensembles 0.7681 0.0440 0.2679 0.0750 0.6733 0.1037 0.3986 0.1258 12.38 13.00 11.63 12.25
GROVER
Deterministic 0.7808 0.0358 0.2473 0.0694 0.6587 0.1043 0.4091 0.1298 11.63 11.50 8.88 9.88
Temperature 0.7810 0.0291 0.2439 0.0686 0.6496 0.1424 0.4612 0.1424 12.63 8.88 7.50 9.25
Focal Loss 0.7779 0.1148 0.3052 0.0811 0.6359 0.1221 0.4365 0.1383 15.00 23.38 21.50 22.25
MC Dropout 0.7817 0.0346 0.2455 0.0689 0.6615 0.1009 0.4042 0.1288 11.25 10.50 7.63 8.63
SWAG 0.7837 0.0360 0.2482 0.0689 0.6603 0.1060 0.4114 0.1301 9.13 11.63 8.88 8.38
BBP 0.7697 0.0438 0.2552 0.0711 0.5995 0.1731 0.5090 0.1660 16.75 22.00 14.75 15.88
SGLD 0.7635 0.0402 0.2558 0.0716 0.5542 0.2712 0.6194 0.2139 18.75 18.63 15.25 15.63
Ensembles 0.7876 0.0316 0.2411 0.0675 0.6646 0.1034 0.4061 0.1290 8.50 8.13 5.38 6.88
Uni-Mol
Deterministic 0.7895 0.0454 0.2601 0.0716 0.6734 0.1020 0.3983 0.1274 11.50 15.50 16.13 13.88
Temperature 0.7896 0.0346 0.2483 0.0704 0.7028 0.1456 0.4566 0.1355 11.00 12.00 13.13 12.75
Focal Loss 0.7904 0.0972 0.2899 0.0785 0.6934 0.1227 0.4079 0.1284 10.00 23.50 19.38 20.50
MC Dropout 0.7891 0.0480 0.2628 0.0726 0.6833 0.1074 0.4015 0.1274 12.88 19.88 19.25 17.25
SWAG 0.7842 0.0593 0.2994 0.0728 0.6870 0.1085 0.4005 0.1271 10.13 22.13 22.25 18.25
BBP 0.7932 0.0396 0.2520 0.0703 0.6273 0.1296 0.4522 0.1456 12.13 17.00 15.50 16.25
SGLD 0.7887 0.0433 0.2569 0.0684 0.5700 0.1953 0.5207 0.1717 14.75 18.13 18.88 14.50
Ensembles 0.8052 0.0332 0.2389 0.0662 0.6841 0.0953 0.3877 0.1247 5.13 8.88 7.63 6.50
TorchMD-NET(c) 0.7793 0.0409 0.2614 0.0708 0.6540 0.1546 0.4424 0.1396 - - - -
GIN(c) 0.6829 0.0634 0.3268 0.0840 0.5752 0.1381 0.4835 0.1477 - - - -
  • (a)
    The “Tox21” and “ToxCast” columns present metric scores on representative exemplar datasets, highlighting the trends observable across all datasets.
  • (b)
    The “Average Ranking” columns provide the rank of each model’s UQ metrics against all other backbone-UQ combinations averaged from all classification datasets; smaller number indicates better performance.
  • (c)
    We report the results from the best-performing UQ method—Deep Ensembles for TorchMD-NET and GIN. These backbones are not ranked together with the primary benchmark.
5 Results and Analysis
Tables 1 and 2 present the macro-averaged rankings of the primary backbone-UQ approaches across all datasets, as well as the scores on two representative datasets. Figure 2 provides further insights into primary backbone model performance by visualizing the Mean Reciprocal Ranks (MRRs) with averaged UQ reciprocal ranks. Compared to the ranks in the tables, MRRs accentuate top results, offering a complementary perspective on the backbone performance. Our analysis initiates with how the UQ methods contribute to the model prediction, followed by an exploration of the uncertainty quantification performance.
Table 2: The regression outcomes for property prediction and uncertainty quantification, where lower scores are preferable, cover two example datasets along with the average ranking. Similar to classification, Deep Ensembles consistently outperforms other methods. Specifically within the context of regression, BBP and SGLD demonstrate exceptional capabilities in estimating uncertainty, despite not consistently improving property prediction outcomes.
Lipophilicity QM9 Average Ranking
RMSE MAE NLL CE RMSE MAE NLL CE RMSE MAE NLL CE
DNN-RDKit
Deterministic 0.7575 0.5793 0.6154 0.0293 0.01511 0.01012 -3.379 0.04419 16.67 15.67 11.33 10.50
MC Dropout 0.7559 0.5773 0.9071 0.0341 0.01480 $0$3.526 0.04327 15.17 15.17 11.33 10.67
SWAG 0.7572 0.5823 0.7191 0.0308 0.01524 0.01019 -3.284 0.04495 18.00 17.67 14.00 12.50
BBP 0.7730 0.5938 0.7578 0.0305 0.01534 0.01025 -3.347 0.04452 21.17 20.50 10.33 7.33
SGLD 0.7468 0.5743 0.2152 0.0090 0.01958 0.01437 -3.335 0.00702 19.83 19.33 6.83 1.83
Ensembles 0.7172 0.5490 0.6165 0.0322 0.01430 0.00956 -3.602 0.04362 11.50 11.17 6.33 8.67
ChemBERTa
Deterministic 0.7553 0.5910 1.2368 0.0362 0.01464 0.00916 -2.410 0.05468 17.83 16.67 18.83 14.67
MC Dropout 0.7142 0.5601 0.8178 0.0349 0.01412 0.00880 -3.150 0.05133 14.33 13.83 15.00 13.67
SWAG 0.7672 0.5992 1.5809 0.0395 0.01477 0.00925 -2.170 0.05535 19.33 18.50 20.33 15.83
BBP 0.7542 0.5869 0.4419 0.0279 0.01443 0.00928 -2.593 0.05399 17.67 18.50 10.83 9.00
SGLD 0.7622 0.5982 0.8719 0.0355 0.01530 0.01012 -3.758 0.03378 19.83 20.50 9.50 8.50
Ensembles 0.7367 0.5763 0.9756 0.0360 0.01397 0.00868 -2.876 0.05425 14.83 13.83 14.67 12.67
GROVER
Deterministic 0.6316 0.4747 2.1512 0.0478 0.01148 0.00678 -0.787 0.06206 10.67 11.67 17.67 16.00
MC Dropout 0.6293 0.4740 2.0526 0.0476 0.01140 0.00676 -1.100 0.06161 9.33 11.17 16.00 14.33
SWAG 0.6317 0.4750 2.3980 0.0485 0.01156 0.00678 -0.477 0.06252 12.33 13.33 20.33 17.83
BBP 0.6481 0.5058 0.0789 0.0196 0.01179 $0$1.885 0.05909 14.17 15.67 7.67 4.00
SGLD 0.6360 0.4984 0.0544 0.0215 0.01359 0.00878 -3.785 0.02911 14.83 15.17 4.67 2.67
Ensembles 0.6250 0.4693 1.6046 0.0460 0.01143 0.00667 -1.028 0.06199 8.17 8.67 13.50 14.00
Uni-Mol
Deterministic 0.6079 0.4509 0.8975 0.0425 0.00962 0.00538 0.014 0.06637 5.83 4.83 16.67 21.17
MC Dropout 0.5983 0.4438 1.3663 0.0440 0.00961 0.00535 -0.251 0.06615 4.00 3.17 16.67 21.33
SWAG 0.6026 0.4476 1.0101 0.0453 0.00969 0.00541 -0.462 0.06597 6.33 6.17 19.67 22.67
BBP 0.6044 0.4469 0.0679 0.0306 0.00952 0.00544 -2.959 0.06179 3.17 3.00 4.33 10.33
SGLD 0.6040 0.4554 0.1565 0.0329 0.00950 0.00546 -4.209 0.04593 2.50 4.00 2.50 9.17
Ensembles 0.5809 0.4266 0.6450 0.0438 0.00948 0.00526 -0.319 0.06629 2.50 1.83 11.00 20.67
TorchMD-NET(a) 1.0313 0.8196 0.8619 0.0195 0.00860 0.00464 2.262 0.06868 - - - -
GIN(a) 0.8071 0.6515 0.3241 0.0020 0.01295 0.00814 -3.521 0.04997 - - - -
  • (a)
    We report the results from the Deep Ensembles UQ method for TorchMD-NET and GIN.
Prediction Performance
Theoretically, inserting UQ methods into the training pipeline does not guarantee better prediction on i.i.d. datasets. However, since we ensure OOD test points with scaffold splitting, UQ methods may mitigate the distribution gap, yielding better test results. The columns ROC-AUC, RMSE, and MAE in Tables 1, 2, and Figure 2 illustrate the predictive performance of each method. Examining from the lens of UQ methods, none provides a consistent guarantee of performance improvement over direct prediction, except for Deep Ensembles, and MC Dropout for regression. The randomness in the initialization and training trajectory of Deep Ensembles explores a broader range of loss landscapes, which partially addresses the distribution shift issue, as observed in previous works (Garipov et al., 2018; Fort et al., 2019; Fang et al., 2023). MC Dropout samples multiple sets of model parameters from the Gaussian distributions centered at the fine-tuned deterministic net weights, which may flatten extreme regression abnormality triggered by OOD features. This phenomenon is less pronounced for classification due to the $(0,1)$ output domain. However, other BNNs do not exhibit the same advantage. SWAG, while similar to MC Dropout w.r.t. training, might intensify training data overfitting due to the additional steps taken to fit the parameter distribution. The stochastic sampling employed by BBP and SGLD complicates network training and may impact the prediction performance as a result.
[…]
Figure 2: MRR of the backbone models, where each metric is macro-averaged from the reciprocal ranks of all corresponding UQ methods across all datasets. Uni-Mol stands out by consistently surpassing its counterparts in property prediction accuracy for both classification and regression tasks. Conversely, GROVER exhibits a more balanced performance in both prediction accuracy and UQ across various tasks.
Looking from the perspective of primary backbones, Uni-Mol secures the best prediction performance for both classification and regression. The superior molecular representation capability of Uni-Mol is attributed to the large network size, the various pre-training data and tasks, and the integration of results from different conformations of the same molecule. When contrasting DNN, ChemBERTa, and GROVER, it becomes apparent that the expressiveness of the molecular descriptors varies for different molecules/tasks. Interestingly, pre-trained models do not invariably surpass heuristic features when integrated with UQ methods. Selecting a model attuned to the specific task is advised over an indiscriminate reliance on “deep and complex” networks.
Uncertainty Quantification Performance
Tables 1, 2 and Figure 2 depict the uncertainty estimation performances via ECE, NLL, Brier score, and CE columns. One discernible trend is the consistent performance enhancement from Deep Ensembles even when the number of ensembles is limited, such as the QM9 case (appendix C). MC Dropout also exhibits a similar trend, albeit less pronounced. Despite a marginal compromise in prediction accuracy, Temperature Scaling emerges as another method that almost invariably improves calibration, as evidenced in the average ranking columns—a finding that aligns with Guo et al. (2017). Figure 3 shows that deterministic prediction tends to be over-confident, and Temperature Scaling mitigates this issue. However, it is susceptible to calibration-test alignment. If the correlation is weak, Temperature Scaling may worsen the calibration, as presented in the ToxCast dataset. In contrast, Focal Loss does not work as impressively for binary classification, which is also observed by Mukhoti et al. (2020). It is hypothesized that while the Sigmoid function sufficiently captures model confidence in binary cases, Focal Loss could over-regularize parameters, diminishing the prediction sharpness to an unreasonable value, a conjecture verified by the “S”-shaped calibration curves in Figure 3.
(a) Uni-Mol on SIDER
(b) Uni-Mol on Tox21
(c) DNN on SIDER
(d) DNN on Tox21
Figure 3: The calibration plot of UQ methods. Here “true probability” refers to the empirical prediction accuracy, and “predicted probability” refers to the average predicted probability within each bin. These results reveal that Temperature Scaling typically enhances calibration effectiveness in the majority of scenarios, whereas Focal Loss tends to lead to over-calibration of the model, producing under-confident results.
Although limited in classification efficacy, both BBP and SGLD deliver commendable performance in predicting regression uncertainty, capturing $7$ out of $8$ top ranks for $4$ backbones on $2$ metrics, narrowly missing out to Deep Ensembles for the remaining one (Table 2). Yet, their inconsistent improvement of RMSE and MAE implies a greater influence on variance prediction than the mean. Figure 4 reveals SGLD’s tendency to “play safe” by predicting larger variances, while the deterministic method is prone to over-confident by ascribing small variances even to its inaccurate predictions. In addition, we do not observe a better correlation between SGLD’s error and variance. We assume that the noisy training trajectory prevents SGLD and BBP from sufficiently minimizing the gap between the predicted mean and true labels, thus encouraging them to maintain larger variances to compensate for the error. Please refer to § C.2 and § E.1 for more analysis on UQ performance.
(a) DNN on Lipo
(b) DNN on FreeSolv
(c) ChemBERTa on Lipo
(d) Uni-Mol on FreeSolv
Figure 4: The absolute error between the model-predicted mean and true labels against the predicted standard deviation. We compare the performance of SGLD with the deterministic prediction on different backbones and datasets. The “$y=kx$” lines indicate whether the true labels lie within the $k$-std range of the predicted Gaussian. Also, a model is perfectly calibrated when its output points are arranged on an “$y=kx$” line for an arbitrary $k$. Notably, SGLD is observed to generate a larger variance for OOD samples, which tends to correspond more closely with the prediction errors.
Figure 2 indicates that Uni-Mol exhibits subpar calibration, particularly for regression. Comparing Uni-Mol, ChemBERTa, and DNN in Figure 4, we notice that larger models such as Uni-Mol are more confident in their results, as illustrated by their smaller variances and larger portion of (std, error) points exceeding $y=kx$. Apart from the inherent susceptibility of larger models to overfitting (Ying, 2019), such phenomenon for Uni-Mol could also be attributed to shared structural features in 3D conformations in the training and test molecules that remain unobserved in simpler descriptors (Zhou et al., 2023). While this similarity benefits property prediction, it also potentially misleads the model into considering data points as in-distribution, thereby erroneously assigning high confidence. Overall, our findings broadly support the notion that models with lower expressiveness tend to exhibit better calibration. Therefore, the selection of an appropriate UQ method is more critical for enhancing the calibration of larger models. This is particularly evident from the larger discrepancies in calibration errors between the Deterministic approach and the most effective UQ method for Uni-Mol compared to DNN, as demonstrated in Tables 1, 2, and the following discussion.
(a) Quantum Mechanics
(b) Physical Chemistry
(c) Biophysics
(d) Physiology
Figure 5: MRRs of TorchMDNet and Uni-Mol on datasets grouped by dataset property categories. MRR calculations are confined to results from these two backbones. Only relative values matter. TorchMDNet is comparible to Uni-Mol on Quantum Mechanics properties, where it is pre-trained on, but is outperformed by Uni-Mol on all other property categories.
TorchMD-NET and GIN
Our analysis also encompasses TorchMD-NET and GIN, two additional backbone models excluded from the primary benchmark due to their limited capabilities. As presented in the tables and Figure 5, TorchMD-NET’s performance is on par with Uni-Mol when predicting quantum mechanical properties but falls short in others. This outcome aligns with expectations, given that TorchMD-NET’s architecture is tailored specifically for predicting quantum mechanical properties (Thölke & Fabritiis, 2022). Moreover, it is pre-trained on the relatively niche dataset PCQM4Mv2 (Hu et al., 2021) with only the denoising objective, which might be suitable for molecular dynamics but limited for other properties. In contrast, Uni-Mol stands out as a versatile model, benefiting from diverse pre-training objectives that ensure superiority across various tasks. On the other hand, GIN’s performance is consistently inferior to other models including DNN, with examples presented in Tables 1 and 2, likely due to the limited expressiveness of 2D graphs and the GNN architecture when pre-training is absent. Appendix E presents the complete results.
Table 3: Performance with frozen backbone weights and on random split datasets compared with the original scores in Table 1 and 2. The result is calculated as $(\text{new}-\text{original})/\text{original}$ and is macro-averaged over all datasets, backbones and UQ methods. The performance of the frozen backbone is significantly worse than the original, while the model performs better on the randomly split (in-domain) test set.
Classification Regression
ROC-AUC (%$\uparrow$ ECE (%$\downarrow$ NLL (%$\downarrow$ Brier Score (%$\downarrow$ RMSE (%$\downarrow$ MAE (%$\downarrow$ NLL (%$\downarrow$ CE (%$\downarrow$
Frozen Backbone -24.07 145.13 53.43 88.98 78.60 92.40 58.18 -46.06
Random Split 13.25 -35.19 -33.87 -37.35 -26.34 -31.36 -53.87 19.06
Frozen Backbone and Randomly Split Datasets
Table 3 demonstrates a notable drop in prediction performance when backbone weights are fixed; and random splits outperform scaffold splits. This is consistent with intuition: if backbone models serve solely as feature extractors instead of a part of the trainable predictors, they are less expressive for downstream tasks. Additionally, in-distribution features tend to be more predictable. An interesting exception emerges in regression calibration error, where frozen backbones perform better and random splits score lower. Upon examining the predicted values, we note that predictions for random splits exhibit a sharper distribution, i.e., smaller $\hat{\sigma}$. This suggests that the models are more confident in regressing in-distribution data, aligning with our previous observation for Uni-Mol. Conversely, frozen backbones are less prone to overfitting due to their constrained expressiveness. This behavior underscores the original models’ capability to distinguish between in-distribution and OOD features and assign confidence scores with precision.
6 Conclusion, Limitation, and Future Works
We introduce MUBen, a benchmark that assesses an array of UQ methods across diverse categories, leveraging backbone models that incorporate different descriptors for multiple molecular property prediction tasks. Our results reveal that Deep Ensembles assures performance enhancement under all circumstances but with significant computation cost. Temperature scaling and MC Dropout are simple yet effective for classification, while BBP and SGLD are better suited for regression UQ. Among backbones, Uni-Mol, leveraging the expressiveness of 3D molecular conformations, is the most powerful but prone to overconfidence. Backbones with other descriptors are advantageous in different conditions and should be selected according to specific usage cases.
Given the rapid advancements in molecular representation learning and UQ methods, MUBen cannot encompass all possible combinations, forcing us to focus on a curated selection of representative methods. Furthermore, we use coarse-grained hyperparameter grids to maintain experimental feasibility, which makes MUBen, while indicative of trends, might not present the best possible results. We remain committed to refining MUBen and welcome contributions from the broader community to enhance its inclusivity and utility for research in this field and related domains.
Acknowledgments
This work was supported in part by ONR MURI N00014-17-1-2656, NSF IIS-2008334, IIS-2106961, and CAREER IIS-2144338.
Appendix A Dataset Details
Table 4: Dataset statistics in MUBen.
Property Category Dataset # Compounds # Tasks Average LIR(a) Max LIR(a)
Classification
Physiology BBBP $2039$ $1$ 0.7651 0.7651
ClinTox $1478$ $2$ 0.9303 0.9364
Tox21 $7831$ $12$ 0.9225 0.9712
ToxCast $8575$ $617$ 0.8336 0.9972
SIDER $1427$ $27$ 0.7485 0.9846
Biophysics BACE $1513$ $1$ 0.5433 0.5433
HIV $41127$ $1$ 0.9649 0.9649
MUV $93087$ $17$ 0.9980 0.9984
Regression
Physical Chemistry ESOL $1128$ $1$
FreeSolv $642$ $1$
Lipophilicity $4200$ $1$
Quantum Mechanics QM7 $7160$ $1$
QM8 $21786$ $12$
QM9 $133885$ $3$(b) - -
  • (a)
    LIR stands for “label imbalance ratio”, which is calculated as $\text{LIR}_{k}\in[0.5,1]=\max\{p_{\text{pos}},p_{\text{neg}}\}$ for a task indexed by $k\in\{1,\dots,K\}$, where $p_{\text{pos}}$ and $p_{\text{neg}}$ are the proportions of positive and negative samples in the dataset, respectively. $\text{LIR}=0.5$ indicates a balanced dataset and $\text{LIR}=1$ indicates a completely imbalanced one, with either $y=0$ or $y=1$ being absent. Average $\text{Average LIR}=\sum_{k=1}^{K}\text{LIR}_{k}$, and Max $\text{Max LIR}=\max_{k}\{\text{LIR}_{k}\}$. LIR is calculated on the entire dataset, and the number could slightly vary for each partition.
  • (b)
    QM9 dataset contains $12$ tasks, but we follow Fang et al. (2022); Zhou et al. (2023) and use only $3$ most popular ones (homo, lumo, and gap).
In this section, we provide details about the datasets used in our study. We follow the approach of previous works (Fang et al., 2022; Zhou et al., 2023) and select a subset of widely used and publicly available datasets from the MoleculeNet benchmark (Wu et al., 2018). The datasets cover both classification and regression tasks from $4$ property categories, including Physiology, Biophysics, Physical Chemistry and Quantum Mechanics. The classification datasets include:
  • BACE provides binary binding properties for a set of inhibitors of human $\beta$-secretase 1 (BACE-1) from experimental values from the published papers;
  • BBBP (Blood-Brain Barrier Penetration) studies the classification of molecules by their permeability of the blood-brain barrier;
  • ClinTox consists of two classification tasks for drugs: whether they are absent of clinical toxicity and whether they are approved by the FDA;
  • Tox21 (Toxicology in the 21st Century) consists of qualitative toxicity measurements of $8014$ compounds on $12$ different targets;
  • ToxCast provides toxicology data of $8576$ compounds on $617$ different targets;
  • SIDER (Side Effect Resource) contains marketed drugs and adverse drug reactions (ADR) extracted from package inserts;
  • HIV, introduced by the Drug Therapeutics Program (DTP) AIDS Antiviral Screen, contains compounds that are either active or inactive against HIV replication;
  • MUV (Maximum Unbiased Validation), a challenging benchmark dataset selected from PubChem BioAssay for validation of virtual screening techniques, contains $93087$ compounds tested for activity against $17$ different targets.
The regression datasets include:
  • ESOL provides experimental values for water solubility data for $1128$ compounds;
  • FreeSolv (Free Solvation Database) contains experimental and calculated hydration free energies for $643$ small molecules;
  • Lipophilicity contains experimental results of octanol/water distribution coefficient for $4200$ compounds;
  • QM7/8/9 are subsets of GDB-13 and GDB-17 databases containing quantum mechanical calculations of energies, enthalpies, and dipole moments for organic molecules with up to $7$/$8$/$9$ “heavy” atoms, respectively. For QM9, we follow previous works and select $3$ tasks that predict properties (homo, lumo, and gap) with a similar quantitative range out of the total $12$ (Fang et al., 2022; Zhou et al., 2023).
A summary of the dataset statistics is provided in Table 4. It is worth noting that some datasets, such as HIV and MUV, exhibit a high degree of class imbalance. This characteristic adds further challenges to the tasks of molecular property prediction and uncertainty quantification.
Scaffold Splitting
MUBen primarily focuses on dataset split by molecule scaffolds, which separates training, validation and test features as much as possible. This creates out-of-distribution (OOD) test sets and signifies the model’s capability of processing patterns absent in the fine-tuning process. For all datasets, we use the conventional 8:1:1 training, validation, and test ratio. The raw datasets are available on the MoleculeNet website. [^1] https://moleculenet.org/ We take advantage of the pre-processed version shared by Zhou et al. (2023) and use the same splits as in their work. [^2] Repository at https://github.com/dptech-corp/Uni-Mol/tree/main/unimol and download at https://bioos-hermite-beijing.tos-cn-beijing.volces.com/unimol_data/finetune/molecular_property_prediction.tar.gz.
Random Splitting
To explore the impact of dataset splitting techniques, we conduct experiments on four classification datasets (BACE, BBBP, ClinTox, Tox21) and four regression datasets (ESOL, FreeSolv, Lipophilicity, QM7) using random splitting. Despite their limited sizes, these datasets effectively represent the overall performance of the models. Each dataset is divided into three separate 8:1:1 training-validation-test sets using random seeds of 00, $1$, and $2$. Each backbone-UQ combination is trained once per split. We then average the results over $3$ runs from the splits to determine the final metrics for each dataset. Deep Ensembles was the only exception, being trained $3$ times for each dataset split using different seeds, totaling $9$ training cycles per dataset. All other training settings were consistent with those in the scaffold-splitting experiments.
Appendix B Backbone Models and Implementation Details
Table 5: Model statistics.
Model # Parameters (M) Average Time per Training Step (ms)(a)
DNN $0.158$ $5.39$
ChemBERTa $3.43$ $30.18$
GROVER $48.71$ $334.47$
Uni-Mol $47.59$ $392.55$
TorchMD-NET $7.23$ $217.29$
GIN $0.26$ $7.21$
  • (a)
    All models are evaluated on the BBBP dataset with a batch size of $128$. We only measure the time for forward passing, backward passing, and parameter updating. We train the model for $6$ epochs and take the average of the last $5$ to reduce the impact of GPU initialization.
Our experimental code is designed on the PyTorch framework (Paszke et al., 2019). The experiments are conducted on a single NVIDIA A100 Tensor Core GPU with a memory capacity of $80$GB. The backbone models are fine-tuned using the AdamW optimizer (Loshchilov & Hutter, 2019) with a weight decay rate of $0.01$ using full-precision floating-point numbers (FP32) for maximum compatibility. We apply different learning rates, numbers of training epochs and batch sizes for the backbone models, as specified in the following paragraphs. We adopt early stopping to select the best-performed checkpoint on the validation set, and all models have achieved their peak validation performance before the training ends. ROC-AUC is selected to assess classification validation performance. For regression, we follow Wu et al. (2018) and use RMSE for Physical Chemistry properties and MAE for Quantum Mechanics. Notice that these metrics only concern prediction, and do not take the uncertainty into account during validation steps (§ D.1). Table 5 presents the number of parameters and average time per training step for each backbone model. The training time is calculated for each model update step instead of each epoch based on our implementation, which might be slower than the models’ original realizations. Below, we offer a detailed description of each backbone model’s architecture and its implementation.
ChemBERTa
ChemBERTa (Chithrananda et al., 2020; Ahmad et al., 2022) leverages the RoBERTa model architecture and pre-training strategy (Liu et al., 2019) but with fewer layers, attention heads, and smaller hidden dimensionalities. Unlike language models which process sentences in natural language, ChemBERTa uses Simplified Molecular-Input Line-Entry System (SMILES, Weininger, 1988) strings as input. This representation is a compact and linear textual depiction of a molecule’s structure that’s frequently employed in cheminformatics. ChemBERTa pre-training adopts a corpus of $77$M SMILES strings from PubChem (Wang et al., 2009), along with the masked language modeling (MLM) objective (Devlin et al., 2019).
ChemBERTa is built with the HuggingFace Transformers library (Wolf et al., 2019), and its pre-trained parameters are shared through the Huggingface’s model hub. We retain the default model architecture and use the DeepChem/ChemBERTa-77M-MLM checkpoint for ChemBERTa’s weight initialization. [^3] Accessible at https://huggingface.co/DeepChem/ChemBERTa-77M-MLM. We employ the last-layer hidden state corresponding to the token [CLS] to represent the input SMILES sequence and attach the output heads specified by the tasks/UQ methods on top of it for property prediction. On all datasets, ChemBERTa is fine-tuned with a learning rate of $10^{-5}$, a batch size of $128$, and for $200$ epochs. A tolerance of $40$ epochs for early stopping is adopted.
GROVER
GROVER (Rong et al., 2020) is pre-trained on $11$M unlabeled molecules as 2D molecular graphs from ZINC15 (Sterling & Irwin, 2015) and ChEMBL (Gaulton et al., 2012). To enhance pre-training on molecular graphs, GROVER integrates the graph neural network (GNN)-based Message Passing Neural Networks (MPNN, Gilmer et al., 2017) into the Transformer encoder architectures. It also devises self-supervised objectives from different molecular structural granularities to better capture the rich structural and semantic information of molecules. Specifically, GROVER replaces the linear layer used for mapping the input query, key, and value vectors of each multi-head attention module into the head-dependent latent feature spaces in the Transformer encoder with a dynamic message passing network. This network aggregates the latent features from the $k$-hop neighboring nodes, where $k$ is a random integer drawn from a pre-defined Uniform of Gaussian. In other words, GROVER puts GNN layers before each head’s self-attention computation to better encode the graph structure of molecules. To facilitate learning, GROVER introduces training objectives including contextual property prediction, which predicts the statistical properties of masked subgraphs of arbitrary sizes, and graph-level motif prediction, which predicts the classes of the masked functional groups. GROVER uses a readout function to aggregate node features, which are then passed into linear layers for property prediction.
For implementation, we use the GROVER-base checkpoint for our model initialization. [^4] Repository available at https://github.com/tencent-ailab/grover, and model can be downloaded at https://ai.tencent.com/ailab/ml/ml-data/grover-models/pretrain/grover_base.tar.gz. We incorporate the “node-view” branch as discussed above and disregard the “edge-view” architecture detailed in the appendix of Rong et al. (2020), in accordance with the default settings in their GitHub repository. Under this configuration, the model generates $2$ sets of node embeddings (but no edge embeddings), one from the node hidden states and another from the edge hidden states (Rong et al., 2020). Each set of embeddings is passed into $2$ linear layers with GELU (Hendrycks & Gimpel, 2016) activation and a dropout ratio of $0.1$ post-readout layer, which simply averages the embeddings in the default configuration. Each set of embeddings corresponds to an individual output branch, predicting the properties independently. In line with the original implementation, we compute the loss for each branch individually during fine-tuning and apply a squared Euclidean distance regularization with a coefficient of $0.1$ between the two. During inference, we average the logits from the $2$ branches to generate the final output logits.
In our experiments, we configure the fine-tuning batch size at $256$, and the number of epochs at $100$ with a tolerance of $40$ epochs for early stopping. The learning rate is set at $10^{-4}$, and the entire model has a dropout ratio of $0.1$. We substitute the original Noam learning rate scheduler (Vaswani et al., 2017) with a linear learning rate scheduler with a $0.1$ warm-up ratio for easier implementation. No substantial differences in model performance were observed between the two learning rate schedulers.
Uni-Mol
Uni-Mol (Zhou et al., 2023) is a universal molecular representation framework that enhances representational capacity and broadens applicability by incorporating 3D molecular structures as model input. For the property prediction task, Uni-Mol undergoes pre-training on $209$M 3D conformations of organic molecules gathered from ZINC15 (Sterling & Irwin, 2015), ChEMBL (Gaulton et al., 2012), and a database comprising $12$M purchasable molecules (Zhou et al., 2023). It portrays atoms as tokens and utilizes pair-type aware Gaussian kernels to encode the positional information in the 3D space, thereby ensuring rotational and translational invariance. Furthermore, Uni-Mol introduces a pair-level representation by orchestrating an “atom-to-pair” communication—updating positional encodings with query-key products—and a “pair-to-atom” communication—adding pair representation as bias terms in the self-attention atom update procedure. For pre-training, Uni-Mol employs masked atom prediction, akin to BERT’s MLM, corrupting 3D positional encodings with random noise at a $15\%$ ratio. Additionally, the model is tasked with restoring the corrupted Euclidean distances between atoms and the coordinates for atoms.
Our codebase is developed atop the publicly accessible Uni-Mol repository, [^5] https://github.com/dptech-corp/Uni-Mol/tree/main. and their pre-trained checkpoint for molecular prediction serves as our model initialization. [^6] https://github.com/dptech-corp/Uni-Mol/releases/download/v0.1/mol_pre_no_h_220816.pt. During fine-tuning, Uni-Mol generates $10$ sets of 3D conformations for each molecule, supplemented with an additional 2D molecular graph. Thereafter Uni-Mol samples one from these $11$ molecular representations for each molecule at the beginning of every training epoch as the input feature. For inference, we average the logits from all $11$ representations to generate the final output. We utilize the conformations prepared by the Uni-Mol repository in our implementation.
We configure the fine-tuning batch size at $128$, the number of epochs at 100, and employ early stopping with a tolerance of $40$ epochs. We use a linear learning rate scheduler with a $0.1$ warm-up ratio and a peak learning rate of $5\times 10^{-5}$. The model is trained with a dropout ratio of $0.1$. Although the Uni-Mol repository provides a set of recommended hyperparameters, we observe no discernible improvement in model performance with these settings.
DNN
The Deep Neural Network (DNN) serves as a simple, randomly initialized baseline model designed to explore how heuristic descriptors like Morgan fingerprints (Morgan, 1965) or RDKit features perform for molecular property prediction. DNN enables us to compare the pre-trained models, which learn the molecular representation automatically through self-learning, with heuristic molecular features, which are constructed manually, and investigate whether or under what circumstances the heuristic features can achieve comparable results. For the descriptor, we adopt the approach of previous work (Wu et al., 2018; Yang et al., 2019; Rong et al., 2020) and extract $200$-dimensional molecule-level features using RDKit for each molecule, which are then used as DNN input. [^7] We use the “rdkit2dnormalized” descriptor in DescriptaStorus, available at https://github.com/bp-kelley/descriptastorus. The DNN consists of $8$ fully connected $128$-dimensional hidden layers with GELU activation and an intervening dropout ratio of $0.1$. We find no performance gain from deeper or wider DNNs and thus assume that our model is fully capable of harnessing the expressivity of RDKit features. The model is trained with a batch size of $256$ and a constant learning rate of $2\times 10^{-4}$ for $400$ epochs with an early stopping tolerance of $50$ epochs.
TorchMD-NET
The architectural design of TorchMD-NET is detailed in (Thölke & Fabritiis, 2022), while its pre-training methodology is discussed in a separate study (Zaidi et al., 2023). TorchMD-NET is an equivariant Transformer tailored for the prediction of quantum mechanical properties of 3D conformers. Unique elements of its architecture include a specialized embedding layer—encoding not just atomic numbers but also the interatomic distances influenced by neighboring atoms, a modified multi-head attention mechanism that integrates edge data, and an equivariant update layer computing atomic interactions. The model undergoes pre-training on the $3.4$M PCQM4Mv2 dataset (Nakata & Shimazaki, 2017; Hu et al., 2021), leveraging a denoising auto-encoder objective. This entails predicting Gaussian noise disturbances on atomic positions, mirroring techniques seen in prevalent diffusion models in computer vision (Song et al., 2021).
To implement TorchMD-NET, we sourced the code and model checkpoint from (Zaidi et al., 2023). [^8] The repository can be accessed athttps://github.com/shehzaidi/pre-training-via-denoising. We made minor architectural adjustments, replacing their single-head output block with our adaptive multi-head output layers. Consequently, we omitted the denoising objective during the fine-tuning process due to compatibility concerns. Our fine-tuning regimen for TorchMD-NET entails a batch size of $128$ over $100$ epochs, adopting an early stopping mechanism with the patience of $40$ epochs. The learning rate peaks at $2\times 10^{-4}$, coupled with a linear scheduler with a $0.1$ warm-up ratio, and the model trains with a dropout ratio of $0.1$.
GIN
Graph Isomorphism Network GIN (Xu et al., 2019) is a randomly initialized model with 2D graph structures as input. Compared with graph convolutional networks (GCN, Kipf & Welling, 2017), GIN mainly differs in that within the neighboring nodes message aggregation process, GIN adds a weight to each node’s self-looping, which is either trainable or pre-defined. In addition, GIN substitutes the one-layer feed-forward network within each GCN layer with a multi-layer perceptron (MLP). It has been proved in theory that these minor changes make GIN among the most powerful graph neural networks (Xu et al., 2019).
We use the Pytorch Geometric (Fey & Lenssen, 2019) to realize GIN. Our implementation contains $5$ GIN layers with $128$ hidden units and $0.1$ dropout ratio. The model is trained with a batch size of $128$ for $200$ epochs with an early stopping tolerance of $50$ epochs, at a constant learning rate of $10^{-4}$.
Appendix C Uncertainty Quantification
C.1 Method and Implementation Details
Focal Loss
First proposed by Lin et al. (2017), Focal Loss is designed to address the class imbalance issue for dense object detection in computer vision, where the number of negative samples (background) far exceeds the number of positive ones (objects). It is adopted for uncertainty estimation and model calibration later by Mukhoti et al. (2020). The idea is to add a modulating factor to the standard cross-entropy loss to down-weight the contribution from easy examples and thus focus more on hard examples. In the binary classification case, it adds a modulating factor $|y_{n}-\hat{p}_{n}|^{\gamma}$ to the standard cross-entropy loss, where $y_{n}\in\{0,1\}$ is the ground truth label, $\hat{p}_{n}\in(0,1)$ is the predicted Sigmoid probability for the $n$-th example, and $\gamma\geq 0$ is a focusing parameter:
$\mathcal{L}_{\rm focal}=-\frac{1}{N}\sum_{n=1}^{N}\left[y_{n}(1-\hat{p}_{n})^{\gamma}\log\hat{p}_{n}+(1-y_{n})\hat{p}_{n}^{\gamma}\log(1-\hat{p}_{n})\right].$ (1)
The focusing parameter $\gamma$ smoothly adjusts the rate at which easy examples are down-weighted. When $\gamma=0$, Focal Loss is equivalent to the cross-entropy loss. As $\gamma$ increases, the effect of the modulating factor increases likewise. In our implementation, we take advantage of the realization in the “torchvision” library and use their default hyper-parameters for all experiments. [^9] https://pytorch.org/vision/main/generated/torchvision.ops.sigmoid_focal_loss.html.
Bayes by Backprop
Bayes by Backprop (BBP, Blundell et al., 2015; Kingma et al., 2015) is an algorithm for training BNNs, where weights are not point estimates but distributions. The idea is to replace the deterministic network weights with Gaussian a porteriori learned from the data, which allows quantifying the uncertainty in the predictions by assembling the predictions from random networks sampled from the posterior distribution of the weights:
$\bm{w}^{\rm MAP}=\operatorname*{arg\,max}_{\bm{w}}\log p(\bm{w}|\mathcal{D})=\operatorname*{arg\,max}_{\bm{w}}\left(\log p(\mathcal{D}|\bm{w})+\log p(\bm{w})\right),$ (2)
where $p(\bm{w})$ is the prior distribution of the weights, which are also Gaussian in our realization.
However, the true posterior is generally intractable for neural networks and can only be approximated with variational inference $q(\bm{w}|\bm{\theta})$ (Kingma & Welling, 2014), where $\bm{\theta}$ are variational parameters. In our case, the variational parameters consist of the mean and standard deviation of multivariate Gaussian distribution $\bm{\theta}=\{\bm{\mu},\bm{\rho}\}$. The learning then involves finding the $\bm{\theta}$ that minimizes the Kullback-Leibler (KL) divergence between the true posterior and the variational distribution. The loss can be written as
$\mathcal{L}=\text{KL}\left[q(\bm{w}|\bm{\theta})||p(\bm{w})\right]-\mathbb{E}_{q(\bm{w}|\bm{\theta})}[\log p(\mathcal{D}|\bm{w})].$ (3)
For each training step $i$, we first draw a sample from a porteriori $\bm{w}_{i}\sim q(\bm{w}|\bm{\theta})$ and then compute the Monte Carlo estimation of the loss:
$\mathcal{L}_{i}\approx\log q(\bm{w}_{i}|\bm{\theta})-\log p(\bm{w}_{i})-\log p(\mathcal{D}|\bm{w}_{i}).$ (4)
During backpropagation, the gradient can be pushed through the sampling process with the reparameterization trick (Kingma & Welling, 2014). Specifically, we adopt the local reparameterization trick (Kingma et al., 2015), which samples the pre-activation $\bm{a}_{i}$ directly from the distribution $q(\bm{a}_{i}|\bm{x}_{i})$ parameterized by the input feature $\bm{x}_{i}$, instead of computing the is as $\bm{a}_{i}=\bm{w}_{i}\bm{x}_{i}$ using the sampled network weights. This has parameters that are deterministic functions of $\bm{x}_{i}$ and $\bm{\theta}$, which reduces the variance of the gradient estimates and can improve the efficiency of the learning process. Following the common practice (Harrison et al., 2023), we only apply BBP to the models’ output layer to reduce computational costs and optimization difficulty for large pre-trained backbone models. In addition, The last layer often directly relates to the task’s uncertainty. Modeling uncertainty in this layer can provide practical benefits in decision-making processes, allowing for an estimation of confidence in the predictions. During inference, we sample $30$ networks from the posterior distribution, generating $30$ sets of prediction results, and computing the mean of the predictions as the final output.
SGLD
Stochastic gradient Langevin dynamics (SGLD, Welling & Teh, 2011) combines the efficiency of stochastic gradient descent (SGD) with Langevin diffusion which introduces the ability to estimate parameter a posteriori. The update rule for SGLD is given by:
$\Delta\bm{\theta}=-\frac{\eta_{t}}{2}\nabla\mathcal{L}(\bm{\theta})+\sqrt{\eta_{t}}\bm{\epsilon},$ (5)
where $\bm{\theta}$ is the network parameters, $\eta_{t}$ is the learning rate at time $t$, and $\bm{\epsilon}_{t}$ is the standard Gaussian noise. With learning rate $\bm{\theta}$ or weight gradient $\nabla\mathcal{L}(\bm{\theta})$ decreasing to small values, the update rule can transit from network optimization to posterior estimation. After sufficient optimization, subsequent samples of parameters can be seen as drawing from the posterior distribution of the model parameters given the data.
In our implementation, we follow the previous implementation and use a constant learning rate. [^10] https://github.com/JavierAntoran/Bayesian-Neural-Networks/. Similar to BBP, we only apply SGLD to the last layer of the model. We first train the model until its performance has stopped improving on the validation set, and then continue training it for another $30$ epochs, resulting in $30$ networks sampled from the Langevin dynamics. This generates $30$ sets of prediction results during the test, and we compute the mean of the predictions as the final output.
MC Dropout
Compared to other Bayesian networks, Monte-Carlo Dropout (MC Dropout, Gal & Ghahramani, 2016) is a simple and efficient for modeling the network uncertainty. Dropout is proposed to prevent overfitting by randomly setting some neurons’ outputs to zero during training (Srivastava et al., 2014). At test time, dropout is deactivated and the weights are scaled down by the dropout rate to simulate the presence of all neurons. In contrast, MC Dropout proposes to keep dropout active during testing and make predictions with dropout turned on. By running several (e.g., $30$ in our experiments) forward passes with random dropout masks, we effectively obtain a Monte Carlo estimation of the predictive distribution.
SWAG
Stochastic Weight Averaging-Gaussian (SWAG, Maddox et al., 2019) is an extension of the Stochastic Weight Averaging (SWA, Izmailov et al., 2018) method, a technique used for finding wider optima in the loss landscape and leads to improved generalization. SWAG fits a Gaussian distribution with a low rank plus diagonal covariance derived from the SGD iterates to approximate the posterior distribution over neural network weights. In SWAG, the model keeps tracking the weights encountered during the last $T$ steps of the stochastic gradient descent updates and computes the Gaussian mean and covariance:
$\begin{gathered}\bm{\mu}_{\bm{\theta}}=\frac{1}{T}\sum_{t=1}^{T}\bm{\theta}_{t};\\ \bm{\Sigma}_{\bm{\theta}}=\frac{1}{T}\sum_{t=1}^{T}(\bm{\theta}_{t}-\bm{\mu}_{\bm{\theta}})(\bm{\theta}_{t}-\bm{\mu}_{\bm{\theta}})^{\mathsf{T}}\text{.}\end{gathered}$ (6)
However, such computation requires storing all the model weights in the last $T$ steps, which is expensive for large models. To address this issue, Maddox et al. (2019) propose to approximate the covariance matrix with a low-rank plus diagonal matrix, and compute the mean and covariance iteratively. Specifically, at update step $t\in\{1,\dots,T\}$,
$\bar{\bm{\theta}}_{t}=\frac{t\bar{\bm{\theta}}_{t-1}+\bm{\theta}_{t}}{t+1};\quad\overline{\bm{\theta}^{2}}_{t}=\frac{t\overline{\bm{\theta}^{2}}_{t-1}+\bm{\theta}_{t}^{2}}{t+1};\quad\widehat{D}_{:,t}=\bm{\theta}_{t}-\bar{\bm{\theta}}_{t},$ (7)
where $\bar{\bm{\theta}}_{0}$ is the best parameter weights found during training prior to the SWA session, $\overline{\bm{\theta}^{2}}_{T}-\bar{\bm{\theta}}^{2}_{T}$ is the covariance diagonal, and $\frac{1}{T-1}\widehat{D}\widehat{D}^{\mathsf{T}}\in^{d,d}$ is the low-rank approximation of the covariance matrix with $d$ being the parameter dimensionality. We can write the Gaussian weight posterior as
$\bm{\theta}_{\rm SWAG}\sim\mathcal{N}_{\rm SWAG}\left(\bar{\bm{\theta}}_{T},\frac{1}{2}\left(\overline{\bm{\theta}^{2}}_{T}-\bar{\bm{\theta}}^{2}_{T}+\frac{1}{T-1}\widehat{D}\widehat{D}^{\mathsf{T}}\right)\right).$ (8)
Notice that $\hat{D}$ has a different rank $K<=T$ in (Maddox et al., 2019), but we set $K=T<d$ for simplicity. Uncertainty in SWAG is estimated by drawing weight samples from $\mathcal{N}_{\rm SWAG}$ and running these through the network. We set $T=20$, and draw $30$ samples during the test in our experiments.
Temperature Scaling
Temperature Scaling (Platt et al., 1999; Guo et al., 2017) is a simple and effective post-hoc method for calibrating the confidence of a neural network. Post-hoc methods calibrate the output probabilities of a pre-trained model without updating the fine-tuned network parameters. The core idea behind Temperature Scaling is to add a learnable parameter $h$ (the temperature) to adjust the output probability of the model. For a trained binary classification model, Temperature Scaling scales the logits $z$ with
$z^{\prime}=\frac{z}{h}$ (9)
before feeding $z^{\prime}$ into the Sigmoid output activation function. The temperature $h$ is learned by minimizing the negative log-likelihood of the training data with other network parameters frozen. For multi-task classification such as Tox21, we assign an individual temperature to each task.
In precise terms, “Temperature Scaling” is introduced for multi-class classification utilizing SoftMax output activation (Guo et al., 2017). For binary classification in our study, we implement Platt Scaling, excluding the bias term (Platt et al., 1999). Nonetheless, we continue using “Temperature Scaling” for its widespread recognition.
Deep Ensembles
Deep Ensembles (Lakshminarayanan et al., 2017) is a technique where multiple deep learning models are independently trained from different initializations, and their predictions are combined to make a final prediction. This approach exploits the idea that different models will make different types of errors, which can be reduced by averaging model predictions, leading to better overall performance and more robust uncertainty estimates (Fort et al., 2019). Formally, given $M$ models in the ensemble, each with parameters $\bm{\theta}_{m},m\in\{1,\dots,M\}$, the ensemble prediction for an input data point $\bm{x}$ is given by:
$\hat{y}=\frac{1}{M}\sum_{m=1}^{M}\hat{y}_{m}=\frac{1}{M}\sum_{m=1}^{M}f(\bm{x};\bm{\theta}_{m})$ (10)
where $f$ represents the model architecture, and $\hat{y}_{m}$ is the post-activation result of the $m$-th model. We set $M=3$ for QM8, QM9 and MUV to reduce computational consumption, and $M=10$ for other scaffold split datasets. For random split, we uniformly use $M=3$, as mentioned in appendix A.
For regression tasks, Lakshminarayanan et al. (2017) aggregate the variances of different network predictions through parameterizing a Gaussian mixture model. In contrast, we take a simpler approach by computing the mean of the variances as the final output variance.
Table 6: Computation resources required by different uncertainty estimation methods. We assume that we have already trained a deterministic backbone model for property prediction, and would like to build up a UQ method on top of it.
UQ Method Training Starting Checkpoint Additional Cost(a)
Deterministic - 00
Temperature from fine-tuned backbone $(T_{\text{infer}}+T_{\text{train-FFN}})\times M_{\text{train-extra}}$
Focal Loss from scratch $T_{\text{train}}\times M_{\text{train}}$
MC Dropout no training $T_{\text{infer}}\times M_{\text{infer}}$
SWAG from fine-tuned backbone $T_{\text{train}}\times M_{\text{train-extra}}+T_{\text{infer}}\times M_{\text{infer}}$
BBP from scratch $T_{\text{train}}\times M_{\text{train}}+T_{\text{infer}}\times M_{\text{infer}}$
SGLD from scratch $T_{\text{train}}\times(M_{\text{train}}+M_{\text{train-extra}})+T_{\text{infer}}\times M_{\text{infer}}$
Ensembles from scratch $T_{\text{train}}\times M_{\text{train}}\times(N_{\text{ensembles}}-1)$
  • (a)
    $T_{\text{train}}$ and $T_{\text{infer}}$ are the time for one epoch of training and inference of the backbone model, respectively. In general, $T_{\text{train}}\gg T_{\text{infer}}$. $M_{\text{train}}$, $M_{\text{train-extra}}$, and $M_{\text{infer}}$ are the number of training epochs, additional training epochs, and inference epochs, respectively (§ C.1). In general, $M_{\text{train}}\gg M_{\text{train-extra}}$. Different backbones and UQ methods have different $T$s and $M$s, but we use the same symbols nonetheless for simplicity. The result is a rough estimation without considering the additional inference time or the early stopping if a model is retrained.
C.2 Resource Analysis
Table 6 summarizes the additional training cost to apply each UQ method to a backbone model already fine-tuned for property prediction. From the table, we can see that post-hoc calibration and MC Dropout are the most efficient methods, while Deep Ensembles is undoubtedly the most expensive one, even though it performs the best most of the time. Several works aim to reduce the computational cost (Wen et al., 2020; Li et al., 2023b), but we do not consider them in MUBen and leave them to future works.
Appendix D Metrics
In this section, we elaborate on the metrics utilized in our research for both property prediction and uncertainty estimation.
D.1 Property Prediction Metrics
ROC-AUC
The receiver operating characteristic area under the curve (ROC-AUC) is widely used for binary classification tasks. The ROC curve plots the true positive rate (TPR), or recall, against the false positive rate (FPR) at various decision thresholds $t\in(0,1)$. Given a set of true positives (TP), true negatives (TN), false positives (FP), and false negatives (FN), the TPR and FPR are computed as $\text{TPR}=\frac{\text{TP}}{\text{TP}+\text{FN}}$ and $\text{FPR}=\frac{\text{FP}}{\text{FP}+\text{TN}}$ respectively. The AUC signifies the likelihood of a randomly selected positive instance being ranked above a randomly chosen negative instance. This integral under the ROC curve is calculated as
[…]
and can be approximated using numerical methods. [^11] We use scikit-learn’s roc_auc_score function in our implementation.
RMSE
For regression tasks, the root mean square error (RMSE) quantifies the average discrepancy between predicted values $\hat{y}_{n}\in$ and actual values $y_{n}\in$ for $N$ data points, given by:
$\text{RMSE}=\sqrt{\frac{1}{N}\sum_{n=1}^{N}(\hat{y}_{n}-y_{n})^{2}}.$ (12)
MAE
The mean absolute error (MAE) is another regression metric, measuring the average absolute deviation between predicted and actual values. It is calculated as:
$\text{MAE}=\frac{1}{N}\sum_{i=1}^{N}|\hat{y}_{n}-y_{n}|.$ (13)
D.2 Uncertainty Quantification Metrics
NLL
Negative log-likelihood (NLL) quantifies the mean deviation between predicted and actual values in logarithmic space and is commonly used to evaluate UQ performance (Gal & Ghahramani, 2016; Lakshminarayanan et al., 2017; Gawlikowski et al., 2021).
For binary classification with Sigmoid output activation, NLL is given by:
$\text{NLL}=-\frac{1}{N}\sum_{n=1}^{N}\left[y_{n}\log\hat{p}_{n}+(1-y_{n})\log(1-\hat{p}_{n})\right],$ (14)
where $y_{n}\in\{0,1\}$ is the true label and $\hat{p}_{n}\in(0,1)=p_{\bm{\theta}}(\hat{y}_{n}\mid\bm{x}_{n})$ is the predicted probability.
For regression, the Gaussian NLL is calculated as:
$\text{NLL}=-\frac{1}{N}\sum_{n=1}^{N}\log\mathcal{N}(y_{n};\hat{y}_{n},\hat{\sigma}_{n})=\frac{1}{N}\sum_{n=1}^{N}\frac{1}{2}\left[\log(2\pi\hat{\sigma}_{n})+\frac{(y_{n}-\hat{y}_{n})^{2}}{\hat{\sigma}_{n}}\right],$ (15)
where $\hat{y}_{n}\in$ is the predicted mean and $\hat{\sigma}_{n}\in_{+}$ is the predicted variance regularized by the SoftPlus activation.
Brier Score
In classification, the Brier score (BS) is a proper scoring rule for measuring the accuracy of predicted probabilities. Similar to mean square error (MSE) in regression, it measures the mean squared difference between predicted probabilities and actual binary outcomes. For binary classification, the Brier score is calculated as
Brier $\text{Brier Score}=\frac{1}{N}\sum_{i=1}^{N}(\hat{p}_{n}-y_{n})^{2},$ (16)
where $y_{n}\in\{0,1\}$ and $\hat{p}_{n}\in(0,1)$.
[…]
ECE
Expected calibration error (ECE) measures the average difference between predicted probabilities and their corresponding empirical probabilities, whose calculation is presented in § 3 and rewritten here specifically for the binary classification case.
Binary ECE puts the predicted probabilities $\{\hat{p}_{n}\}_{n=1}^{N}$ into $S$ (which is set to $S=15$ in our experiments) equal-width bins $B_{s}=\{n\in\{1,\dots,N\}\mid\hat{p}_{n}\in(\rho_{s},\rho_{s+1}]\}$ with $s\in\{1,\dots,S\}$. For binary classification, ECE is calculated as
$\begin{gathered}\text{ECE}=\sum_{s=1}^{S}\frac{|B_{s}|}{N}\left|\text{acc}(B_{s})-\text{conf}(B_{s})\right|;\\ \text{acc}(B_{s})=\frac{1}{|B_{s}|}\sum_{n\in B_{s}}y_{n};\quad\quad\text{conf}(B_{s})=\frac{1}{|B_{s}|}\sum_{n\in B_{s}}\hat{p}_{n}.\\ \end{gathered}$ (17)
Here, $\text{acc}(\cdot)$ is the predicion accuracy and $\text{conf}(\cdot)$ is the confidence level within each bin. Notice that ||$\cdot$|| ⋅ | represents the size when used with a set, and the absolute value when its input is a real number.
Regression CE
Regression calibration is less studied than classification calibration, and regression calibration error (CE), proposed by Kuleshov et al. (2018), is adopted in only a few works (Song et al., 2019; Kamarthi et al., 2021). Similar to ECE, CE measures the average difference between the observed confidence level and the expected confidence level within each bin. Specifically, a confidence level of $\alpha$ is measured as the fraction of data points whose true value falls within $\frac{1-\alpha}{2}$th quantile and $(1-\frac{\alpha}{2})$th quantile of the predicted distribution. For example, $90\%$ confidence level contains data points with true values falling within the $5\%$th quantile and $95\%$th quantile of the predicted distribution. With this established, we can calculate the confidence levels with the predicted quantile function, in our case, the cumulative distribution function (CDF) of the predicted Gaussian. Specifically,
$\text{CE}=\frac{1}{S}\sum_{s=1}^{S}\left(\rho_{s}-\frac{1}{N}\left|\left\{n\in\{1,\dots,N\}\mid\Phi(y_{n};\hat{y}_{n},\hat{\sigma}_{n})\leq\rho_{s}\right\}\right|\right)^{2}$ (18)
where $\rho_{s}=\frac{s}{S}$ is the expected quantile, $\Phi$ is the Gaussian CDF:
[…]
$\hat{p}_{s}=\frac{1}{N}\left|\left\{n\in\{1,\dots,N\}\mid\Phi(y_{n};\hat{y}_{n},\hat{\sigma}_{n})\leq\rho_{s}\right\}\right|$ is the empirical frequency of the predicted values falling into the $\rho_{s}$th quantile. This is slightly different from the confidence level-based calculation, where
$\text{CE}=\frac{2}{S}\sum_{s=1}^{\frac{S}{2}}\left(\left(\rho_{S-s}-\rho_{s}\right)-\left(\hat{p}_{S-s}-\hat{p}_{s}\right)\right)^{2},$ (20)
but it reveals the same trend. We use $S=20$ in our experiments.
Appendix E Additional Results
In this section, we supplement the discussions from § 5 with further results and analyses. Given the extensive volume of outcomes from our experiments, we have consolidated them into tables at the end of this article. For enhanced accessibility and offline analysis, we have also made these results available as CSV files in our GitHub repository.
E.1 More Visualization of Uncertainty Estimation Performance
Table 7 displays the average rankings of UQ methods, complemented by Figure 6 which offers a visual representation. As elaborated in § 5, both Deep Ensembles and MC Dropout generally enhance prediction and uncertainty estimation across tasks. However, the performance boost is more noticeable with Deep Ensembles compared to MC Dropout, while DC Dropout is much cheaper to apply. For classification tasks, Temperature Scaling offers moderate efficacy. Given the minimal computational cost of implementing Temperature Scaling, it emerges as a favorable option for calibration enhancement when necessary. In regression tasks, both BBP and SGLD stand out as reliable choices irrespective of the backbone model, particularly for uncertainty estimation. Nonetheless, they mandate a full re-training. Conversely, SWAG and Focal Loss are underwhelming and warrant reconsideration in selection processes while dealing with molecular property prediction.
E.2 Complete Fine-Tuning Results
Tables 916 illustrate scores associated with each backbone-UQ pairing across the scaffold split classification datasets, while Tables 1722 detail the scores regarding regression metrics. Scores are represented in a “mean (standard deviation)” structure, where both mean and standard deviation values are derived from $3$ experiment runs conducted with random seeds 00, $1$, and $2$. As Deep Ensembles aggregates network predictions prior to metric calculation, standard deviation values are not generated and metrics are calculated only once.
Challenges are encountered during training on the notably imbalanced HIV and MUV dataset, depicted in Table 4. SGLD at times proves unsuccessful due to weak directional cues from the true labels, overwhelmed by the random noise embedded in the Langevin dynamics. With TorchMD-NET backbone, SGLD even fails training and constantly produces NaN values regardless of the hyper-parameters we choose. Conversely, the application of Focal Loss demonstrates effective performance on the MUV dataset, aligning with assertions in Lin et al. (2017). ECE also may fail to compute on these datasets as predicted probabilities $\hat{p}$ are excessively minimal, leading to unpopulated bins and subsequent “division by zero” errors.
In terms of the majority of datasets, the performance trends closely adhere to our previous discussions. However, an anomaly is observed with the BBBP and ClinTox datasets, where ChemBERTa produces higher performance than the more powerful models, deviating from prior findings (Ahmad et al., 2022; Zhou et al., 2023). Conceptually, the ChemBERTa checkpoint we use is solely pre-trained on the MLM objective, which is unrelated to downstream tasks and should not include label information. Consequently, we hypothesize that either the SMILES strings are occasionally more expressive than other descriptors with respect to the physiological properties, or the checkpoint may have undergone fine-tuning prior to upload, resulting in potential contamination.
Table 7: Averaged ranks of UQ methods across all datasets and backbone models. Lower value indicates better resuls. Temperature Scaling and Focal Loss are not applicable to regression tasks.
Classification Regression
ROC-AUC ECE NLL BS RMSE MAE NLL CE
Deterministic 16.12 15.41 17.03 16.53 12.75 12.21 16.12 15.58
Temperature 16.31 12.06 13.97 14.38 - - - -
Focal Loss 17.12 22.84 20.41 21.97 - - - -
MC Dropout 15.78 13.53 14.28 13.62 10.71 10.83 14.75 15.0
SWAG 15.47 17.94 19.62 17.69 14.0 13.92 18.58 17.21
BBP 18.97 20.28 17.0 18.88 14.04 14.42 8.29 7.67
SGLD 20.72 20.66 20.59 19.22 14.25 14.75 5.87 5.54
Ensembles 11.5 9.28 9.09 9.72 9.25 8.88 11.38 14.0
[…]
Figure 6: MRR of the UQ methods for each metric, each is macro-averaged from the reciprocal ranks of the results of all corresponding backbone models on all datasets.
E.3 Results on Random Split Datasets
Tables 23 to 30 illustrate the efficacy of primary backbone models across datasets with random divisions. Mirroring the discussion in § E.1, Table 8 and Figure 7 depict the average (reciprocal) rankings of UQ methods. For the most part, the comparative performance among the backbone models and UQ methods is consistent with those observed in scaffold split datasets. Nonetheless, certain exceptions emerge, especially concerning BBP, SGLD, and SWAG. On randomly split classification datasets, both BBP and SGLD continue to underperform, while a marked decline is noted in their regression NLL. This phenomenon might be attributed to the balanced nature of random split datasets, intensifying the predictive indications from actual labels. Given this environment, the noise from Langevin dynamics or Gaussian samples, which usually aids generalization, might become detrimental. This revelation highlights the capability of traditional Bayesian methods in aptly navigating OOD data (i.e., addressing epistemic uncertainty), mainly through curbing overfitting. However, they may struggle when dealing with in-distribution attributes (i.e., addressing aleatoric uncertainty) having relatively stable labels in the context of molecular property prediction. On the other hand, SWAG, another approximation of the Bayesian network that is almost useless on OOD data, delivers a much more impressive performance on random split datasets. Deep Ensembles, MC Dropout, and Temperature Scaling still emerge as the most reliable UQ methods, although Deep Ensembles no longer dominates the benchmark in the in-distribution case.
Upon comparing Figure 8 with Figure 2, it’s evident that the relative positions of the backbone models remain largely unchanged for regression datasets across both splitting methods; however, this consistency falters for classification datasets. Notably, Uni-Mol displays a marked performance decline in the randomly split datasets, positioning it below the more straightforward and cost-efficient backbones, ChemBERTa and GROVER. While definitive explanations for this observed behavior remain elusive, we hypothesize that the molecular descriptor or the pre-training data inherent to ChemBERTa and GROVER could render them better equipped to adapt to Physiological and Biophysical features. Such potential advantages might be obscured by their model architectures’ suboptimal generalization capabilities when scaffold splitting is employed.
Table 8: Averaged ranks of UQ methods across all randomly split datasets.
Classification Regression
ROC-AUC ECE NLL BS RMSE MAE NLL CE
Deterministic 16.44 16.31 17.06 16.5 13.25 12.69 12.62 15.88
Temperature 16.88 14.0 13.06 14.5 - - - -
Focal Loss 17.31 28.75 23.62 26.56 - - - -
MC Dropout 14.88 9.56 12.88 13.75 10.81 10.94 10.75 15.56
SWAG 13.81 15.0 15.06 14.0 10.44 10.31 11.19 16.44
BBP 19.94 16.12 18.12 17.31 16.56 17.5 15.94 5.44
SGLD 18.44 18.81 19.0 17.56 14.12 14.12 14.19 6.69
Ensembles 14.31 13.44 13.19 11.81 9.81 9.44 10.31 15.0
[…]
Figure 7: MRR of the UQ methods for each metric, each is macro-averaged from the reciprocal ranks of the results of all corresponding backbone models on all randomly split datasets.
[…]
Figure 8: MRR of the UQ methods for each backbone model, each is macro-averaged from the reciprocal ranks of the results of all corresponding UQ methods on all randomly split datasets.
E.4 Results with Frozen Backbone Models
Analysis of Tables 31 through 38 reveals that when the network parameters of the backbone models are fixed, all models underperform significantly. Among the three primary backbone models, Uni-Mol experiences the most drastic decline, emerging as the least effective in the majority of evaluations. These findings highlight that these backbone models are not inherently optimized for mere feature extraction. The fine-tuning process is pivotal for ensuring optimal model performance, irrespective of the use of UQ methods.
Table 9: Test results on BACE in the format of “metric mean $\pm$ standard deviation”.
ROC-AUC ECE NLL BS
DNN-rdkit Deterministic 0.8185 $\pm$ 0.0164 0.2827 $\pm$ 0.0155 0.8946 $\pm$ 0.0797 0.2810 $\pm$ 0.0125
Temperature 0.8185 $\pm$ 0.0164 0.2537 $\pm$ 0.0144 0.7412 $\pm$ 0.0472 0.2553 $\pm$ 0.0111
Focal Loss 0.8190 $\pm$ 0.0112 0.2543 $\pm$ 0.0011 0.6908 $\pm$ 0.0099 0.2488 $\pm$ 0.0027
MC Dropout 0.8168 $\pm$ 0.0121 0.2648 $\pm$ 0.0323 0.7667 $\pm$ 0.0446 0.2589 $\pm$ 0.0157
SWAG 0.8173 $\pm$ 0.0119 0.2728 $\pm$ 0.0068 0.8730 $\pm$ 0.0643 0.2730 $\pm$ 0.0085
BBP 0.8288 $\pm$ 0.0039 0.2711 $\pm$ 0.0193 0.7943 $\pm$ 0.0235 0.2535 $\pm$ 0.0075
SGLD 0.8181 $\pm$ 0.0160 0.2713 $\pm$ 0.0024 0.9023 $\pm$ 0.0416 0.2734 $\pm$ 0.0073
Ensembles 0.8207 $\pm$ - 0.2550 $\pm$ - 0.7985 $\pm$ - 0.2557 $\pm$ -
ChemBERTa Deterministic 0.7223 $\pm$ 0.0089 0.2407 $\pm$ 0.0135 0.7567 $\pm$ 0.0203 0.2720 $\pm$ 0.0081
Temperature 0.7223 $\pm$ 0.0089 0.1758 $\pm$ 0.0059 0.6937 $\pm$ 0.0118 0.2498 $\pm$ 0.0052
Focal Loss 0.6985 $\pm$ 0.0062 0.2140 $\pm$ 0.0070 0.7322 $\pm$ 0.0051 0.2679 $\pm$ 0.0023
MC Dropout 0.7613 $\pm$ 0.0216 0.2391 $\pm$ 0.0262 0.7464 $\pm$ 0.0257 0.2673 $\pm$ 0.0070
SWAG 0.7580 $\pm$ 0.0139 0.2539 $\pm$ 0.0091 0.7707 $\pm$ 0.0374 0.2654 $\pm$ 0.0139
BBP 0.7372 $\pm$ 0.0233 0.2767 $\pm$ 0.0032 0.8325 $\pm$ 0.0190 0.2878 $\pm$ 0.0046
SGLD 0.7775 $\pm$ 0.0205 0.2492 $\pm$ 0.0217 0.7922 $\pm$ 0.0283 0.2588 $\pm$ 0.0069
Ensembles 0.7350 $\pm$ - 0.2352 $\pm$ - 0.7457 $\pm$ - 0.2683 $\pm$ -
GROVER Deterministic 0.8308 $\pm$ 0.0012 0.2304 $\pm$ 0.0658 0.6921 $\pm$ 0.1114 0.2342 $\pm$ 0.0386
Temperature 0.8305 $\pm$ 0.0017 0.2083 $\pm$ 0.0521 0.6251 $\pm$ 0.0671 0.2184 $\pm$ 0.0283
Focal Loss 0.8509 $\pm$ 0.0072 0.2334 $\pm$ 0.0302 0.6397 $\pm$ 0.0300 0.2281 $\pm$ 0.0141
MC Dropout 0.8278 $\pm$ 0.0011 0.2366 $\pm$ 0.0629 0.6899 $\pm$ 0.1094 0.2368 $\pm$ 0.0394
SWAG 0.8604 $\pm$ 0.0048 0.2239 $\pm$ 0.0016 0.6656 $\pm$ 0.0169 0.2148 $\pm$ 0.0030
BBP 0.8322 $\pm$ 0.0099 0.2570 $\pm$ 0.0331 0.7001 $\pm$ 0.0583 0.2450 $\pm$ 0.0228
SGLD 0.8494 $\pm$ 0.0028 0.2281 $\pm$ 0.0066 0.6769 $\pm$ 0.0144 0.2204 $\pm$ 0.0038
Ensembles 0.8388 $\pm$ - 0.2288 $\pm$ - 0.6571 $\pm$ - 0.2252 $\pm$ -
Uni-Mol Deterministic 0.8080 $\pm$ 0.0266 0.2419 $\pm$ 0.0409 0.7928 $\pm$ 0.1639 0.2517 $\pm$ 0.0335
Temperature 0.8087 $\pm$ 0.0277 0.2125 $\pm$ 0.0259 0.6782 $\pm$ 0.0662 0.2352 $\pm$ 0.0212
Focal Loss 0.8208 $\pm$ 0.0259 0.2445 $\pm$ 0.0287 0.6721 $\pm$ 0.0482 0.2406 $\pm$ 0.0198
MC Dropout 0.8268 $\pm$ 0.0269 0.2483 $\pm$ 0.0333 0.8065 $\pm$ 0.1383 0.2512 $\pm$ 0.0331
SWAG 0.8193 $\pm$ 0.0106 0.2677 $\pm$ 0.0075 0.9415 $\pm$ 0.0623 0.2625 $\pm$ 0.0058
BBP 0.8386 $\pm$ 0.0135 0.2673 $\pm$ 0.0473 0.8158 $\pm$ 0.1584 0.2588 $\pm$ 0.0375
SGLD 0.8699 $\pm$ 0.0118 0.2437 $\pm$ 0.0131 0.7354 $\pm$ 0.0502 0.2244 $\pm$ 0.0150
Ensembles 0.8505 $\pm$ - 0.2176 $\pm$ - 0.6301 $\pm$ - 0.2184 $\pm$ -
TorchMD-NET Deterministic 0.7945 $\pm$ 0.0232 0.2193 $\pm$ 0.0374 1.1663 $\pm$ 0.6182 0.2272 $\pm$ 0.0201
Temperature 0.7945 $\pm$ 0.0231 0.1808 $\pm$ 0.0425 0.7151 $\pm$ 0.1720 0.2085 $\pm$ 0.0187
Focal Loss 0.7602 $\pm$ 0.0089 0.2141 $\pm$ 0.0249 0.7328 $\pm$ 0.0376 0.2420 $\pm$ 0.0200
MC Dropout 0.8018 $\pm$ 0.0260 0.2017 $\pm$ 0.0297 0.8454 $\pm$ 0.2767 0.2178 $\pm$ 0.0199
SWAG 0.7457 $\pm$ 0.0208 0.3287 $\pm$ 0.0164 1.9159 $\pm$ 0.3026 0.3172 $\pm$ 0.0122
BBP 0.7638 $\pm$ 0.0374 0.3130 $\pm$ 0.0457 2.8823 $\pm$ 1.2929 0.3035 $\pm$ 0.0425
SGLD 0.7787 $\pm$ 0.0173 0.2793 $\pm$ 0.0199 1.1648 $\pm$ 0.2135 0.2731 $\pm$ 0.0226
Ensembles 0.8169 $\pm$ - 0.2220 $\pm$ - 0.6703 $\pm$ - 0.2197 $\pm$ -
GIN Deterministic 0.7312 $\pm$ 0.0152 0.4024 $\pm$ 0.0362 1.5262 $\pm$ 0.4443 0.3946 $\pm$ 0.0340
Temperature 0.7312 $\pm$ 0.0152 0.3494 $\pm$ 0.0392 1.0850 $\pm$ 0.1950 0.3482 $\pm$ 0.0285
Focal Loss 0.7272 $\pm$ 0.0374 0.3034 $\pm$ 0.1213 1.2375 $\pm$ 0.6434 0.3264 $\pm$ 0.0917
MC Dropout 0.7312 $\pm$ 0.0152 0.4024 $\pm$ 0.0362 1.5262 $\pm$ 0.4443 0.3946 $\pm$ 0.0340
SWAG 0.5554 $\pm$ 0.0075 0.1929 $\pm$ 0.0326 0.7817 $\pm$ 0.0442 0.2839 $\pm$ 0.0156
BBP 0.6775 $\pm$ 0.0533 0.2567 $\pm$ 0.0644 1.3036 $\pm$ 0.7463 0.2842 $\pm$ 0.0546
SGLD 0.6682 $\pm$ 0.0302 0.2271 $\pm$ 0.1028 0.8664 $\pm$ 0.2399 0.2887 $\pm$ 0.0569
Ensembles 0.7767 $\pm$ - 0.2051 $\pm$ - 0.7157 $\pm$ - 0.2469 $\pm$ -
Table 10: Test results on BBBP in the format of “metric mean $\pm$ standard deviation”.
ROC-AUC ECE NLL BS
DNN-rdkit Deterministic 0.6721 $\pm$ 0.0129 0.3292 $\pm$ 0.0517 3.8147 $\pm$ 4.0782 0.3423 $\pm$ 0.0378
Temperature 0.6719 $\pm$ 0.0129 0.2890 $\pm$ 0.0742 2.1338 $\pm$ 1.8963 0.3186 $\pm$ 0.0489
Focal Loss 0.6693 $\pm$ 0.0047 0.1689 $\pm$ 0.1056 0.8925 $\pm$ 0.3424 0.2623 $\pm$ 0.0470
MC Dropout 0.6669 $\pm$ 0.0151 0.3287 $\pm$ 0.0402 2.1699 $\pm$ 1.7478 0.3391 $\pm$ 0.0296
SWAG 0.6777 $\pm$ 0.0136 0.3020 $\pm$ 0.0694 4.8415 $\pm$ 5.5407 0.3268 $\pm$ 0.0465
BBP 0.6856 $\pm$ 0.0075 0.3392 $\pm$ 0.0506 1.8126 $\pm$ 0.6030 0.3463 $\pm$ 0.0357
SGLD 0.6922 $\pm$ 0.0125 0.3134 $\pm$ 0.0487 1.7569 $\pm$ 1.0592 0.3247 $\pm$ 0.0396
Ensembles 0.7018 $\pm$ - 0.2615 $\pm$ - 0.9043 $\pm$ - 0.2938 $\pm$ -
ChemBERTa Deterministic 0.7407 $\pm$ 0.0101 0.2259 $\pm$ 0.0206 1.0306 $\pm$ 0.1002 0.2559 $\pm$ 0.0090
Temperature 0.7407 $\pm$ 0.0101 0.2107 $\pm$ 0.0196 0.8539 $\pm$ 0.0266 0.2453 $\pm$ 0.0064
Focal Loss 0.7410 $\pm$ 0.0122 0.1639 $\pm$ 0.0118 0.6995 $\pm$ 0.0246 0.2313 $\pm$ 0.0087
MC Dropout 0.7451 $\pm$ 0.0067 0.2082 $\pm$ 0.0035 0.8814 $\pm$ 0.0292 0.2418 $\pm$ 0.0032
SWAG 0.7408 $\pm$ 0.0112 0.2373 $\pm$ 0.0202 1.1435 $\pm$ 0.0810 0.2629 $\pm$ 0.0109
BBP 0.7245 $\pm$ 0.0112 0.2415 $\pm$ 0.0293 0.9653 $\pm$ 0.1062 0.2652 $\pm$ 0.0120
SGLD 0.7359 $\pm$ 0.0104 0.2398 $\pm$ 0.0256 1.0692 $\pm$ 0.1940 0.2580 $\pm$ 0.0056
Ensembles 0.7399 $\pm$ - 0.2290 $\pm$ - 0.9672 $\pm$ - 0.2506 $\pm$ -
GROVER Deterministic 0.7075 $\pm$ 0.0061 0.2902 $\pm$ 0.0111 0.9457 $\pm$ 0.0234 0.2998 $\pm$ 0.0060
Temperature 0.7072 $\pm$ 0.0059 0.2685 $\pm$ 0.0147 0.8607 $\pm$ 0.0258 0.2865 $\pm$ 0.0069
Focal Loss 0.7146 $\pm$ 0.0059 0.1340 $\pm$ 0.0177 0.6505 $\pm$ 0.0070 0.2284 $\pm$ 0.0040
MC Dropout 0.7066 $\pm$ 0.0071 0.2901 $\pm$ 0.0147 0.9267 $\pm$ 0.0285 0.2996 $\pm$ 0.0071
SWAG 0.7160 $\pm$ 0.0032 0.3314 $\pm$ 0.0032 1.1105 $\pm$ 0.0142 0.3251 $\pm$ 0.0025
BBP 0.6968 $\pm$ 0.0152 0.2735 $\pm$ 0.0209 0.9146 $\pm$ 0.0326 0.2986 $\pm$ 0.0055
SGLD 0.7266 $\pm$ 0.0037 0.3127 $\pm$ 0.0065 1.0187 $\pm$ 0.0202 0.3149 $\pm$ 0.0019
Ensembles 0.7083 $\pm$ - 0.2890 $\pm$ - 0.9338 $\pm$ - 0.3029 $\pm$ -
Uni-Mol Deterministic 0.7103 $\pm$ 0.0075 0.3311 $\pm$ 0.0506 1.4209 $\pm$ 0.3000 0.3354 $\pm$ 0.0397
Temperature 0.7093 $\pm$ 0.0083 0.2922 $\pm$ 0.0418 1.0023 $\pm$ 0.1064 0.3084 $\pm$ 0.0315
Focal Loss 0.7102 $\pm$ 0.0045 0.2031 $\pm$ 0.0816 0.7646 $\pm$ 0.0923 0.2632 $\pm$ 0.0275
MC Dropout 0.7029 $\pm$ 0.0142 0.3328 $\pm$ 0.0483 1.4086 $\pm$ 0.2828 0.3382 $\pm$ 0.0372
SWAG 0.7247 $\pm$ 0.0062 0.3547 $\pm$ 0.0136 1.8530 $\pm$ 0.0824 0.3456 $\pm$ 0.0078
BBP 0.6902 $\pm$ 0.0286 0.3420 $\pm$ 0.0318 1.3548 $\pm$ 0.0713 0.3463 $\pm$ 0.0275
SGLD 0.7147 $\pm$ 0.0118 0.3336 $\pm$ 0.0127 1.4174 $\pm$ 0.0475 0.3359 $\pm$ 0.0059
Ensembles 0.7253 $\pm$ - 0.3035 $\pm$ - 1.0889 $\pm$ - 0.3035 $\pm$ -
TorchMD-NET Deterministic 0.6753 $\pm$ 0.0060 0.3207 $\pm$ 0.0249 1.7352 $\pm$ 0.2714 0.3299 $\pm$ 0.0218
Temperature 0.6753 $\pm$ 0.0060 0.2739 $\pm$ 0.0251 1.1382 $\pm$ 0.1017 0.2995 $\pm$ 0.0188
Focal Loss 0.7060 $\pm$ 0.0206 0.1910 $\pm$ 0.0200 0.8688 $\pm$ 0.0498 0.2487 $\pm$ 0.0119
MC Dropout 0.6767 $\pm$ 0.0057 0.3098 $\pm$ 0.0236 1.4864 $\pm$ 0.1962 0.3190 $\pm$ 0.0177
SWAG 0.6720 $\pm$ 0.0131 0.3601 $\pm$ 0.0121 3.1976 $\pm$ 0.5027 0.3632 $\pm$ 0.0090
BBP 0.6895 $\pm$ 0.0166 0.3019 $\pm$ 0.0598 3.1956 $\pm$ 1.7451 0.3159 $\pm$ 0.0343
SGLD 0.6708 $\pm$ 0.0058 0.3629 $\pm$ 0.0083 2.3208 $\pm$ 0.5515 0.3617 $\pm$ 0.0068
Ensembles 0.6844 $\pm$ - 0.2700 $\pm$ - 1.1018 $\pm$ - 0.2934 $\pm$ -
GIN Deterministic 0.6282 $\pm$ 0.0025 0.2277 $\pm$ 0.0302 1.0566 $\pm$ 0.0130 0.2920 $\pm$ 0.0078
Temperature 0.6281 $\pm$ 0.0025 0.1987 $\pm$ 0.0164 0.9092 $\pm$ 0.0166 0.2778 $\pm$ 0.0051
Focal Loss 0.6234 $\pm$ 0.0081 0.1550 $\pm$ 0.0116 0.7581 $\pm$ 0.0143 0.2611 $\pm$ 0.0043
MC Dropout 0.6282 $\pm$ 0.0025 0.2277 $\pm$ 0.0302 1.0566 $\pm$ 0.0130 0.2920 $\pm$ 0.0078
SWAG 0.6067 $\pm$ 0.0131 0.3162 $\pm$ 0.0021 1.2223 $\pm$ 0.0235 0.3353 $\pm$ 0.0056
BBP 0.6100 $\pm$ 0.0095 0.2738 $\pm$ 0.0056 1.1617 $\pm$ 0.0529 0.3119 $\pm$ 0.0058
SGLD 0.5577 $\pm$ 0.0497 0.3493 $\pm$ 0.0217 1.4064 $\pm$ 0.1006 0.3681 $\pm$ 0.0218
Ensembles 0.6298 $\pm$ - 0.2112 $\pm$ - 0.9696 $\pm$ - 0.2802 $\pm$ -
Table 11: Test results on ClinTox in the format of “metric mean $\pm$ standard deviation”.
ROC-AUC ECE NLL BS
DNN-rdkit Deterministic 0.8301 $\pm$ 0.0226 0.0637 $\pm$ 0.0126 0.2719 $\pm$ 0.0481 0.0597 $\pm$ 0.0052
Temperature 0.8300 $\pm$ 0.0224 0.0547 $\pm$ 0.0109 0.2371 $\pm$ 0.0059 0.0572 $\pm$ 0.0039
Focal Loss 0.8507 $\pm$ 0.0165 0.0812 $\pm$ 0.0368 0.2284 $\pm$ 0.0305 0.0639 $\pm$ 0.0114
MC Dropout 0.8194 $\pm$ 0.0217 0.0445 $\pm$ 0.0055 0.2149 $\pm$ 0.0307 0.0534 $\pm$ 0.0005
SWAG 0.8290 $\pm$ 0.0163 0.0590 $\pm$ 0.0117 0.2964 $\pm$ 0.0617 0.0591 $\pm$ 0.0054
BBP 0.7541 $\pm$ 0.0102 0.0574 $\pm$ 0.0121 0.2439 $\pm$ 0.0276 0.0609 $\pm$ 0.0046
SGLD 0.8335 $\pm$ 0.0109 0.0567 $\pm$ 0.0125 0.2769 $\pm$ 0.0452 0.0553 $\pm$ 0.0063
Ensembles 0.8154 $\pm$ - 0.0283 $\pm$ - 0.1860 $\pm$ - 0.0484 $\pm$ -
ChemBERTa Deterministic 0.9856 $\pm$ 0.0028 0.0215 $\pm$ 0.0035 0.1206 $\pm$ 0.0318 0.0225 $\pm$ 0.0016
Temperature 0.9856 $\pm$ 0.0028 0.0199 $\pm$ 0.0026 0.0964 $\pm$ 0.0147 0.0222 $\pm$ 0.0013
Focal Loss 0.9832 $\pm$ 0.0027 0.0897 $\pm$ 0.0378 0.1367 $\pm$ 0.0350 0.0273 $\pm$ 0.0045
MC Dropout 0.9848 $\pm$ 0.0021 0.0212 $\pm$ 0.0019 0.1083 $\pm$ 0.0247 0.0213 $\pm$ 0.0021
SWAG 0.9850 $\pm$ 0.0031 0.0222 $\pm$ 0.0025 0.1327 $\pm$ 0.0317 0.0232 $\pm$ 0.0010
BBP 0.9816 $\pm$ 0.0079 0.0325 $\pm$ 0.0212 0.0972 $\pm$ 0.0156 0.0226 $\pm$ 0.0024
SGLD 0.9869 $\pm$ 0.0012 0.0183 $\pm$ 0.0006 0.0893 $\pm$ 0.0050 0.0214 $\pm$ 0.0005
Ensembles 0.9812 $\pm$ - 0.0157 $\pm$ - 0.0810 $\pm$ - 0.0208 $\pm$ -
GROVER Deterministic 0.9398 $\pm$ 0.0152 0.0635 $\pm$ 0.0114 0.1471 $\pm$ 0.0106 0.0367 $\pm$ 0.0031
Temperature 0.9409 $\pm$ 0.0130 0.0449 $\pm$ 0.0055 0.1260 $\pm$ 0.0045 0.0333 $\pm$ 0.0034
Focal Loss 0.9377 $\pm$ 0.0042 0.2158 $\pm$ 0.0034 0.2963 $\pm$ 0.0018 0.0731 $\pm$ 0.0011
MC Dropout 0.9347 $\pm$ 0.0152 0.0706 $\pm$ 0.0122 0.1475 $\pm$ 0.0121 0.0366 $\pm$ 0.0036
SWAG 0.9623 $\pm$ 0.0060 0.0460 $\pm$ 0.0117 0.1119 $\pm$ 0.0046 0.0280 $\pm$ 0.0022
BBP 0.9085 $\pm$ 0.0109 0.0940 $\pm$ 0.0104 0.1976 $\pm$ 0.0136 0.0488 $\pm$ 0.0033
SGLD 0.9507 $\pm$ 0.0058 0.0595 $\pm$ 0.0065 0.1287 $\pm$ 0.0063 0.0306 $\pm$ 0.0011
Ensembles 0.9429 $\pm$ - 0.0743 $\pm$ - 0.1453 $\pm$ - 0.0344 $\pm$ -
Uni-Mol Deterministic 0.8814 $\pm$ 0.0379 0.0522 $\pm$ 0.0120 0.1822 $\pm$ 0.0381 0.0484 $\pm$ 0.0124
Temperature 0.8830 $\pm$ 0.0375 0.0417 $\pm$ 0.0158 0.1754 $\pm$ 0.0358 0.0477 $\pm$ 0.0121
Focal Loss 0.8836 $\pm$ 0.0202 0.1651 $\pm$ 0.0556 0.2972 $\pm$ 0.0587 0.0827 $\pm$ 0.0207
MC Dropout 0.8745 $\pm$ 0.0438 0.0560 $\pm$ 0.0131 0.1889 $\pm$ 0.0406 0.0521 $\pm$ 0.0138
SWAG 0.9055 $\pm$ 0.0027 0.0311 $\pm$ 0.0034 0.1450 $\pm$ 0.0035 0.0337 $\pm$ 0.0009
BBP 0.9257 $\pm$ 0.0287 0.0379 $\pm$ 0.0026 0.1468 $\pm$ 0.0047 0.0398 $\pm$ 0.0032
SGLD 0.8598 $\pm$ 0.0209 0.0327 $\pm$ 0.0029 0.1619 $\pm$ 0.0059 0.0389 $\pm$ 0.0014
Ensembles 0.9088 $\pm$ - 0.0389 $\pm$ - 0.1476 $\pm$ - 0.0386 $\pm$ -
TorchMD-NET Deterministic 0.8938 $\pm$ 0.0178 0.0567 $\pm$ 0.0058 0.1905 $\pm$ 0.0083 0.0470 $\pm$ 0.0014
Temperature 0.8938 $\pm$ 0.0178 0.0608 $\pm$ 0.0100 0.1869 $\pm$ 0.0076 0.0487 $\pm$ 0.0016
Focal Loss 0.8952 $\pm$ 0.0195 0.1260 $\pm$ 0.0242 0.2582 $\pm$ 0.0218 0.0683 $\pm$ 0.0050
MC Dropout 0.8898 $\pm$ 0.0192 0.0505 $\pm$ 0.0029 0.1711 $\pm$ 0.0054 0.0460 $\pm$ 0.0013
SWAG 0.8995 $\pm$ 0.0185 0.0517 $\pm$ 0.0082 0.1969 $\pm$ 0.0065 0.0486 $\pm$ 0.0027
BBP 0.8801 $\pm$ 0.0278 0.0667 $\pm$ 0.0139 0.2483 $\pm$ 0.0463 0.0610 $\pm$ 0.0077
SGLD 0.8855 $\pm$ 0.0177 0.0475 $\pm$ 0.0045 0.1776 $\pm$ 0.0123 0.0475 $\pm$ 0.0045
Ensembles 0.8944 $\pm$ - 0.0569 $\pm$ - 0.1886 $\pm$ - 0.0485 $\pm$ -
GIN Deterministic 0.6965 $\pm$ 0.0184 0.1042 $\pm$ 0.0000 0.3398 $\pm$ 0.0260 0.0798 $\pm$ 0.0044
Temperature 0.6965 $\pm$ 0.0184 0.0876 $\pm$ 0.0000 0.3595 $\pm$ 0.0346 0.0789 $\pm$ 0.0044
Focal Loss 0.6777 $\pm$ 0.0168 0.2690 $\pm$ 0.0232 0.4610 $\pm$ 0.0314 0.1448 $\pm$ 0.0148
MC Dropout 0.6965 $\pm$ 0.0184 0.1042 $\pm$ 0.0000 0.3398 $\pm$ 0.0260 0.0798 $\pm$ 0.0044
SWAG 0.6552 $\pm$ 0.0197 $0$$\pm$ 0.0000 0.4573 $\pm$ 0.1485 0.0799 $\pm$ 0.0031
BBP 0.6122 $\pm$ 0.0078 - $\pm$ - 0.5270 $\pm$ 0.1577 0.0888 $\pm$ 0.0131
SGLD 0.6539 $\pm$ 0.0312 0.1256 $\pm$ 0.0000 0.4745 $\pm$ 0.1417 0.0827 $\pm$ 0.0046
Ensembles 0.6872 $\pm$ - 0.1170 $\pm$ - 0.3222 $\pm$ - 0.0868 $\pm$ -
Table 12: Test results on Tox21 in the format of “metric mean $\pm$ standard deviation”.
ROC-AUC ECE NLL BS
DNN-rdkit Deterministic 0.7386 $\pm$ 0.0061 0.0417 $\pm$ 0.0028 0.2771 $\pm$ 0.0024 0.0779 $\pm$ 0.0013
Temperature 0.7386 $\pm$ 0.0061 0.0342 $\pm$ 0.0027 0.2723 $\pm$ 0.0028 0.0773 $\pm$ 0.0013
Focal Loss 0.7374 $\pm$ 0.0038 0.1058 $\pm$ 0.0060 0.3161 $\pm$ 0.0078 0.0871 $\pm$ 0.0024
MC Dropout 0.7376 $\pm$ 0.0039 0.0356 $\pm$ 0.0016 0.2727 $\pm$ 0.0016 0.0763 $\pm$ 0.0013
SWAG 0.7364 $\pm$ 0.0045 0.0438 $\pm$ 0.0010 0.2793 $\pm$ 0.0037 0.0790 $\pm$ 0.0017
BBP 0.7243 $\pm$ 0.0036 0.0422 $\pm$ 0.0019 0.2847 $\pm$ 0.0010 0.0814 $\pm$ 0.0002
SGLD 0.7257 $\pm$ 0.0018 0.1192 $\pm$ 0.0332 0.3455 $\pm$ 0.0315 0.0978 $\pm$ 0.0100
Ensembles 0.7540 $\pm$ - 0.0344 $\pm$ - 0.2648 $\pm$ - 0.0746 $\pm$ -
ChemBERTa Deterministic 0.7542 $\pm$ 0.0009 0.0571 $\pm$ 0.0020 0.2962 $\pm$ 0.0067 0.0812 $\pm$ 0.0009
Temperature 0.7542 $\pm$ 0.0009 0.0424 $\pm$ 0.0040 0.2744 $\pm$ 0.0027 0.0792 $\pm$ 0.0011
Focal Loss 0.7523 $\pm$ 0.0022 0.0969 $\pm$ 0.0012 0.3052 $\pm$ 0.0018 0.0845 $\pm$ 0.0004
MC Dropout 0.7641 $\pm$ 0.0032 0.0423 $\pm$ 0.0023 0.2697 $\pm$ 0.0032 0.0744 $\pm$ 0.0005
SWAG 0.7538 $\pm$ 0.0021 0.0592 $\pm$ 0.0022 0.3008 $\pm$ 0.0079 0.0818 $\pm$ 0.0011
BBP 0.7433 $\pm$ 0.0055 0.0459 $\pm$ 0.0028 0.2765 $\pm$ 0.0035 0.0780 $\pm$ 0.0014
SGLD 0.7475 $\pm$ 0.0036 0.0504 $\pm$ 0.0038 0.2784 $\pm$ 0.0030 0.0795 $\pm$ 0.0014
Ensembles 0.7681 $\pm$ - 0.0440 $\pm$ - 0.2679 $\pm$ - 0.0750 $\pm$ -
GROVER Deterministic 0.7808 $\pm$ 0.0017 0.0358 $\pm$ 0.0038 0.2473 $\pm$ 0.0022 0.0694 $\pm$ 0.0005
Temperature 0.7810 $\pm$ 0.0016 0.0291 $\pm$ 0.0016 0.2439 $\pm$ 0.0008 0.0686 $\pm$ 0.0003
Focal Loss 0.7779 $\pm$ 0.0028 0.1148 $\pm$ 0.0102 0.3052 $\pm$ 0.0094 0.0811 $\pm$ 0.0028
MC Dropout 0.7817 $\pm$ 0.0018 0.0346 $\pm$ 0.0015 0.2455 $\pm$ 0.0016 0.0689 $\pm$ 0.0005
SWAG 0.7837 $\pm$ 0.0017 0.0359 $\pm$ 0.0009 0.2482 $\pm$ 0.0009 0.0689 $\pm$ 0.0001
BBP 0.7697 $\pm$ 0.0031 0.0438 $\pm$ 0.0017 0.2552 $\pm$ 0.0016 0.0711 $\pm$ 0.0003
SGLD 0.7635 $\pm$ 0.0009 0.0402 $\pm$ 0.0024 0.2558 $\pm$ 0.0010 0.0716 $\pm$ 0.0002
Ensembles 0.7876 $\pm$ - 0.0316 $\pm$ - 0.2411 $\pm$ - 0.0675 $\pm$ -
Uni-Mol Deterministic 0.7895 $\pm$ 0.0017 0.0454 $\pm$ 0.0056 0.2601 $\pm$ 0.0072 0.0716 $\pm$ 0.0022
Temperature 0.7896 $\pm$ 0.0016 0.0346 $\pm$ 0.0027 0.2483 $\pm$ 0.0035 0.0704 $\pm$ 0.0014
Focal Loss 0.7904 $\pm$ 0.0040 0.0972 $\pm$ 0.0045 0.2899 $\pm$ 0.0009 0.0785 $\pm$ 0.0001
MC Dropout 0.7891 $\pm$ 0.0021 0.0480 $\pm$ 0.0066 0.2628 $\pm$ 0.0065 0.0726 $\pm$ 0.0021
SWAG 0.7842 $\pm$ 0.0048 0.0593 $\pm$ 0.0025 0.2994 $\pm$ 0.0054 0.0728 $\pm$ 0.0012
BBP 0.7932 $\pm$ 0.0043 0.0396 $\pm$ 0.0048 0.2520 $\pm$ 0.0039 0.0703 $\pm$ 0.0010
SGLD 0.7887 $\pm$ 0.0059 0.0433 $\pm$ 0.0023 0.2569 $\pm$ 0.0055 0.0684 $\pm$ 0.0014
Ensembles 0.8052 $\pm$ - 0.0332 $\pm$ - 0.2389 $\pm$ - 0.0662 $\pm$ -
TorchMD-NET Deterministic 0.7342 $\pm$ 0.0143 0.0738 $\pm$ 0.0146 0.4523 $\pm$ 0.1083 0.0857 $\pm$ 0.0044
Temperature 0.7342 $\pm$ 0.0143 0.0515 $\pm$ 0.0047 0.2944 $\pm$ 0.0115 0.0801 $\pm$ 0.0014
Focal Loss 0.7403 $\pm$ 0.0032 0.0639 $\pm$ 0.0044 0.3102 $\pm$ 0.0140 0.0808 $\pm$ 0.0027
MC Dropout 0.7369 $\pm$ 0.0144 0.0685 $\pm$ 0.0131 0.3913 $\pm$ 0.0759 0.0837 $\pm$ 0.0033
SWAG 0.7347 $\pm$ 0.0093 0.0740 $\pm$ 0.0040 0.4567 $\pm$ 0.0576 0.0832 $\pm$ 0.0018
BBP 0.7459 $\pm$ 0.0067 0.0808 $\pm$ 0.0075 0.5355 $\pm$ 0.1830 0.0888 $\pm$ 0.0016
SGLD 0.7387 $\pm$ 0.0051 0.0427 $\pm$ 0.0019 0.2761 $\pm$ 0.0029 0.0757 $\pm$ 0.0004
Ensembles 0.7793 $\pm$ - 0.0409 $\pm$ - 0.2614 $\pm$ - 0.0708 $\pm$ -
GIN Deterministic 0.6706 $\pm$ 0.0032 0.0789 $\pm$ 0.0172 0.3573 $\pm$ 0.0052 0.0916 $\pm$ 0.0070
Temperature 0.6706 $\pm$ 0.0032 0.0914 $\pm$ 0.0158 0.3393 $\pm$ 0.0150 0.0941 $\pm$ 0.0068
Focal Loss 0.6571 $\pm$ 0.0061 0.1040 $\pm$ 0.0073 0.3361 $\pm$ 0.0036 0.0930 $\pm$ 0.0014
MC Dropout 0.6706 $\pm$ 0.0032 0.0789 $\pm$ 0.0172 0.3573 $\pm$ 0.0052 0.0916 $\pm$ 0.0070
SWAG 0.6619 $\pm$ 0.0103 0.0671 $\pm$ 0.0032 0.3536 $\pm$ 0.0029 0.0867 $\pm$ 0.0026
BBP 0.6425 $\pm$ 0.0034 0.0851 $\pm$ 0.0084 0.3632 $\pm$ 0.0039 0.0949 $\pm$ 0.0033
SGLD 0.6190 $\pm$ 0.0048 0.0967 $\pm$ 0.0135 0.3817 $\pm$ 0.0275 0.0995 $\pm$ 0.0056
Ensembles 0.6829 $\pm$ - 0.0634 $\pm$ - 0.3268 $\pm$ - 0.0840 $\pm$ -
Table 13: Test results on ToxCast in the format of “metric mean $\pm$ standard deviation”.
ROC-AUC ECE NLL BS
DNN-rdkit Deterministic 0.6222 $\pm$ 0.0042 0.1168 $\pm$ 0.0096 0.4436 $\pm$ 0.0146 0.1397 $\pm$ 0.0021
Temperature 0.6220 $\pm$ 0.0046 0.1114 $\pm$ 0.0050 0.4882 $\pm$ 0.0170 0.1398 $\pm$ 0.0015
Focal Loss 0.6289 $\pm$ 0.0077 0.1264 $\pm$ 0.0034 0.4389 $\pm$ 0.0084 0.1396 $\pm$ 0.0024
MC Dropout 0.6248 $\pm$ 0.0030 0.1093 $\pm$ 0.0075 0.4319 $\pm$ 0.0096 0.1358 $\pm$ 0.0015
SWAG 0.6207 $\pm$ 0.0073 0.1175 $\pm$ 0.0098 0.4440 $\pm$ 0.0150 0.1400 $\pm$ 0.0023
BBP 0.6020 $\pm$ 0.0028 0.1443 $\pm$ 0.0060 0.4673 $\pm$ 0.0059 0.1510 $\pm$ 0.0020
SGLD 0.5319 $\pm$ 0.0040 0.3054 $\pm$ 0.0034 0.6685 $\pm$ 0.0035 0.2378 $\pm$ 0.0017
Ensembles 0.6486 $\pm$ - $0$$\pm$ - 0.4008 $\pm$ - 0.1292 $\pm$ -
ChemBERTa Deterministic 0.6554 $\pm$ 0.0037 0.1209 $\pm$ 0.0026 0.4313 $\pm$ 0.0054 0.1330 $\pm$ 0.0008
Temperature 0.6540 $\pm$ 0.0042 0.1067 $\pm$ 0.0019 0.4817 $\pm$ 0.0261 0.1313 $\pm$ 0.0004
Focal Loss 0.6442 $\pm$ 0.0170 0.1197 $\pm$ 0.0037 0.4243 $\pm$ 0.0153 0.1346 $\pm$ 0.0051
MC Dropout 0.6624 $\pm$ 0.0024 0.1069 $\pm$ 0.0036 0.4070 $\pm$ 0.0052 0.1276 $\pm$ 0.0015
SWAG 0.6556 $\pm$ 0.0033 0.1202 $\pm$ 0.0036 0.4305 $\pm$ 0.0064 0.1327 $\pm$ 0.0010
BBP 0.5814 $\pm$ 0.0075 0.1276 $\pm$ 0.0015 0.4545 $\pm$ 0.0011 0.1469 $\pm$ 0.0003
SGLD 0.5436 $\pm$ 0.0056 0.2238 $\pm$ 0.0019 0.5602 $\pm$ 0.0022 0.1881 $\pm$ 0.0011
Ensembles 0.6733 $\pm$ - 0.1037 $\pm$ - 0.3986 $\pm$ - 0.1258 $\pm$ -
GROVER Deterministic 0.6587 $\pm$ 0.0018 0.1043 $\pm$ 0.0006 0.4091 $\pm$ 0.0023 0.1298 $\pm$ 0.0004
Temperature 0.6496 $\pm$ 0.0023 0.1424 $\pm$ 0.0026 0.4612 $\pm$ 0.0045 0.1424 $\pm$ 0.0013
Focal Loss 0.6359 $\pm$ 0.0026 0.1221 $\pm$ 0.0005 0.4365 $\pm$ 0.0012 0.1383 $\pm$ 0.0004
MC Dropout 0.6615 $\pm$ 0.0014 0.1009 $\pm$ 0.0010 0.4042 $\pm$ 0.0022 0.1288 $\pm$ 0.0004
SWAG 0.6603 $\pm$ 0.0020 0.1060 $\pm$ 0.0007 0.4114 $\pm$ 0.0015 0.1301 $\pm$ 0.0001
BBP 0.5995 $\pm$ 0.0026 0.1731 $\pm$ 0.0041 0.5090 $\pm$ 0.0040 0.1660 $\pm$ 0.0015
SGLD 0.5542 $\pm$ 0.0069 0.2712 $\pm$ 0.0008 0.6194 $\pm$ 0.0006 0.2139 $\pm$ 0.0003
Ensembles 0.6646 $\pm$ - 0.1034 $\pm$ - 0.4061 $\pm$ - 0.1290 $\pm$ -
Uni-Mol Deterministic 0.6734 $\pm$ 0.0101 0.1020 $\pm$ 0.0023 0.3983 $\pm$ 0.0075 0.1274 $\pm$ 0.0025
Temperature 0.7028 $\pm$ 0.0101 0.1456 $\pm$ 0.0045 0.4566 $\pm$ 0.0104 0.1355 $\pm$ 0.0029
Focal Loss 0.6934 $\pm$ 0.0035 0.1227 $\pm$ 0.0011 0.4079 $\pm$ 0.0016 0.1284 $\pm$ 0.0003
MC Dropout 0.6833 $\pm$ 0.0117 0.1074 $\pm$ 0.0090 0.4015 $\pm$ 0.0112 0.1274 $\pm$ 0.0025
SWAG 0.6870 $\pm$ 0.0075 0.1085 $\pm$ 0.0034 0.4005 $\pm$ 0.0055 0.1271 $\pm$ 0.0019
BBP 0.6273 $\pm$ 0.0066 0.1296 $\pm$ 0.0065 0.4522 $\pm$ 0.0068 0.1456 $\pm$ 0.0021
SGLD 0.5700 $\pm$ 0.0005 0.1953 $\pm$ 0.0033 0.5207 $\pm$ 0.0029 0.1717 $\pm$ 0.0012
Ensembles 0.6841 $\pm$ - 0.0953 $\pm$ - 0.3877 $\pm$ - 0.1247 $\pm$ -
TorchMD-NET Deterministic 0.6361 $\pm$ 0.0001 0.1734 $\pm$ 0.0053 0.4870 $\pm$ 0.0164 0.1486 $\pm$ 0.0027
Temperature 0.6384 $\pm$ 0.0020 - $\pm$ - 0.6053 $\pm$ 0.0984 0.1466 $\pm$ 0.0009
Focal Loss 0.6443 $\pm$ 0.0073 0.1543 $\pm$ 0.0032 0.4337 $\pm$ 0.0011 0.1395 $\pm$ 0.0006
MC Dropout 0.6339 $\pm$ 0.0042 0.1707 $\pm$ 0.0048 0.4778 $\pm$ 0.0141 0.1477 $\pm$ 0.0027
SWAG 0.6457 $\pm$ 0.0064 0.1648 $\pm$ 0.0012 0.4617 $\pm$ 0.0025 0.1446 $\pm$ 0.0004
BBP 0.6103 $\pm$ 0.0182 0.1471 $\pm$ 0.0049 0.4539 $\pm$ 0.0171 0.1434 $\pm$ 0.0026
SGLD 0.5224 $\pm$ 0.0048 0.2976 $\pm$ 0.0020 0.6678 $\pm$ 0.0020 0.2374 $\pm$ 0.0010
Ensembles 0.6540 $\pm$ - 0.1546 $\pm$ - 0.4424 $\pm$ - 0.1396 $\pm$ -
GIN Deterministic 0.5667 $\pm$ 0.0032 0.1443 $\pm$ 0.0018 0.5142 $\pm$ 0.0048 0.1506 $\pm$ 0.0015
Temperature 0.5669 $\pm$ 0.0028 0.1402 $\pm$ 0.0013 0.5333 $\pm$ 0.0192 0.1476 $\pm$ 0.0014
Focal Loss 0.5483 $\pm$ 0.0028 0.1446 $\pm$ 0.0027 0.4615 $\pm$ 0.0009 0.1468 $\pm$ 0.0006
MC Dropout 0.5667 $\pm$ 0.0032 0.1443 $\pm$ 0.0018 0.5142 $\pm$ 0.0048 0.1506 $\pm$ 0.0015
SWAG 0.5633 $\pm$ 0.0059 0.1418 $\pm$ 0.0018 0.5125 $\pm$ 0.0052 0.1495 $\pm$ 0.0014
BBP 0.5096 $\pm$ 0.0059 0.2243 $\pm$ 0.0007 0.6405 $\pm$ 0.0128 0.1935 $\pm$ 0.0014
SGLD 0.4872 $\pm$ 0.0050 0.3166 $\pm$ 0.0019 0.8083 $\pm$ 0.0245 0.2668 $\pm$ 0.0047
Ensembles 0.5752 $\pm$ - 0.1381 $\pm$ - 0.4835 $\pm$ - 0.1477 $\pm$ -
Table 14: Test results on SIDER in the format of “metric mean $\pm$ standard deviation”.
ROC-AUC ECE NLL BS
DNN-rdkit Deterministic 0.5981 $\pm$ 0.0152 0.1188 $\pm$ 0.0344 0.5462 $\pm$ 0.0469 0.1760 $\pm$ 0.0124
Temperature 0.5981 $\pm$ 0.0152 0.1033 $\pm$ 0.0271 0.5186 $\pm$ 0.0270 0.1709 $\pm$ 0.0094
Focal Loss 0.5864 $\pm$ 0.0071 0.1663 $\pm$ 0.0217 0.5605 $\pm$ 0.0193 0.1893 $\pm$ 0.0075
MC Dropout 0.5977 $\pm$ 0.0164 0.1071 $\pm$ 0.0271 0.5245 $\pm$ 0.0304 0.1705 $\pm$ 0.0084
SWAG 0.5988 $\pm$ 0.0210 0.1206 $\pm$ 0.0333 0.5477 $\pm$ 0.0491 0.1764 $\pm$ 0.0125
BBP 0.5738 $\pm$ 0.0272 0.1215 $\pm$ 0.0046 0.5241 $\pm$ 0.0077 0.1733 $\pm$ 0.0027
SGLD 0.5895 $\pm$ 0.0087 0.1559 $\pm$ 0.0119 0.5653 $\pm$ 0.0119 0.1900 $\pm$ 0.0051
Ensembles 0.6158 $\pm$ - 0.0954 $\pm$ - 0.4950 $\pm$ - 0.1631 $\pm$ -
ChemBERTa Deterministic 0.6146 $\pm$ 0.0078 0.1483 $\pm$ 0.0102 0.5740 $\pm$ 0.0217 0.1804 $\pm$ 0.0055
Temperature 0.6146 $\pm$ 0.0078 0.1275 $\pm$ 0.0104 0.5256 $\pm$ 0.0138 0.1721 $\pm$ 0.0043
Focal Loss 0.6119 $\pm$ 0.0105 0.1498 $\pm$ 0.0136 0.5346 $\pm$ 0.0129 0.1777 $\pm$ 0.0046
MC Dropout 0.6166 $\pm$ 0.0022 0.1136 $\pm$ 0.0143 0.5191 $\pm$ 0.0131 0.1672 $\pm$ 0.0037
SWAG 0.6149 $\pm$ 0.0095 0.1527 $\pm$ 0.0090 0.5836 $\pm$ 0.0230 0.1825 $\pm$ 0.0056
BBP 0.6067 $\pm$ 0.0108 0.1184 $\pm$ 0.0018 0.5127 $\pm$ 0.0043 0.1686 $\pm$ 0.0013
SGLD 0.6127 $\pm$ 0.0110 0.1299 $\pm$ 0.0118 0.5255 $\pm$ 0.0157 0.1725 $\pm$ 0.0053
Ensembles 0.6193 $\pm$ - 0.1305 $\pm$ - 0.5396 $\pm$ - 0.1718 $\pm$ -
GROVER Deterministic 0.6213 $\pm$ 0.0027 0.1034 $\pm$ 0.0061 0.5023 $\pm$ 0.0071 0.1637 $\pm$ 0.0021
Temperature 0.6209 $\pm$ 0.0030 0.0939 $\pm$ 0.0040 0.4935 $\pm$ 0.0046 0.1618 $\pm$ 0.0018
Focal Loss 0.6059 $\pm$ 0.0165 0.1817 $\pm$ 0.0019 0.5666 $\pm$ 0.0027 0.1903 $\pm$ 0.0011
MC Dropout 0.6248 $\pm$ 0.0060 0.0969 $\pm$ 0.0035 0.4942 $\pm$ 0.0057 0.1616 $\pm$ 0.0018
SWAG 0.6133 $\pm$ 0.0024 0.1183 $\pm$ 0.0016 0.5280 $\pm$ 0.0066 0.1702 $\pm$ 0.0019
BBP 0.5926 $\pm$ 0.0052 0.1347 $\pm$ 0.0014 0.5367 $\pm$ 0.0023 0.1762 $\pm$ 0.0009
SGLD 0.5989 $\pm$ 0.0129 0.1012 $\pm$ 0.0040 0.5017 $\pm$ 0.0051 0.1637 $\pm$ 0.0019
Ensembles 0.6237 $\pm$ - 0.0878 $\pm$ - 0.4917 $\pm$ - 0.1604 $\pm$ -
Uni-Mol Deterministic 0.6214 $\pm$ 0.0073 0.1054 $\pm$ 0.0130 0.5047 $\pm$ 0.0184 0.1661 $\pm$ 0.0065
Temperature 0.6210 $\pm$ 0.0075 0.0950 $\pm$ 0.0074 0.4937 $\pm$ 0.0101 0.1625 $\pm$ 0.0040
Focal Loss 0.6200 $\pm$ 0.0118 0.1660 $\pm$ 0.0096 0.5471 $\pm$ 0.0090 0.1834 $\pm$ 0.0038
MC Dropout 0.6125 $\pm$ 0.0060 0.1269 $\pm$ 0.0178 0.5324 $\pm$ 0.0270 0.1739 $\pm$ 0.0071
SWAG 0.6246 $\pm$ 0.0090 0.1327 $\pm$ 0.0084 0.5397 $\pm$ 0.0125 0.1753 $\pm$ 0.0033
BBP 0.6301 $\pm$ 0.0260 0.0991 $\pm$ 0.0130 0.4928 $\pm$ 0.0119 0.1612 $\pm$ 0.0049
SGLD 0.6323 $\pm$ 0.0108 0.1286 $\pm$ 0.0093 0.5408 $\pm$ 0.0137 0.1725 $\pm$ 0.0030
Ensembles 0.6378 $\pm$ - 0.0872 $\pm$ - 0.4836 $\pm$ - 0.1585 $\pm$ -
TorchMD-NET Deterministic 0.5881 $\pm$ 0.0060 0.1305 $\pm$ 0.0239 0.5988 $\pm$ 0.0541 0.1789 $\pm$ 0.0072
Temperature 0.5881 $\pm$ 0.0060 0.1142 $\pm$ 0.0109 0.5380 $\pm$ 0.0147 0.1721 $\pm$ 0.0038
Focal Loss 0.5909 $\pm$ 0.0048 0.1515 $\pm$ 0.0034 0.5470 $\pm$ 0.0058 0.1848 $\pm$ 0.0031
MC Dropout 0.5862 $\pm$ 0.0059 0.1230 $\pm$ 0.0205 0.5551 $\pm$ 0.0327 0.1756 $\pm$ 0.0061
SWAG 0.6081 $\pm$ 0.0085 0.1281 $\pm$ 0.0152 0.5596 $\pm$ 0.0326 0.1751 $\pm$ 0.0059
BBP 0.6072 $\pm$ 0.0150 0.1293 $\pm$ 0.0005 0.5619 $\pm$ 0.0059 0.1753 $\pm$ 0.0008
SGLD 0.6026 $\pm$ 0.0089 0.1143 $\pm$ 0.0076 0.5162 $\pm$ 0.0054 0.1693 $\pm$ 0.0017
Ensembles 0.6329 $\pm$ - 0.1093 $\pm$ - 0.5175 $\pm$ - 0.1660 $\pm$ -
GIN Deterministic 0.6159 $\pm$ 0.0016 $0$$\pm$ 0.0020 0.5385 $\pm$ 0.0084 0.1627 $\pm$ 0.0009
Temperature 0.6159 $\pm$ 0.0017 0.0916 $\pm$ 0.0031 0.5219 $\pm$ 0.0125 0.1625 $\pm$ 0.0009
Focal Loss 0.5892 $\pm$ 0.0046 0.1863 $\pm$ 0.0011 0.5764 $\pm$ 0.0034 0.1953 $\pm$ 0.0013
MC Dropout 0.6159 $\pm$ 0.0016 $0$$\pm$ 0.0020 0.5385 $\pm$ 0.0084 0.1627 $\pm$ 0.0009
SWAG 0.6054 $\pm$ 0.0034 0.0959 $\pm$ 0.0042 0.5520 $\pm$ 0.0021 0.1642 $\pm$ 0.0011
BBP 0.5499 $\pm$ 0.0132 0.1477 $\pm$ 0.0071 0.5666 $\pm$ 0.0079 0.1876 $\pm$ 0.0037
SGLD 0.5792 $\pm$ 0.0074 0.1121 $\pm$ 0.0028 0.6447 $\pm$ 0.0421 0.1723 $\pm$ 0.0018
Ensembles 0.6232 $\pm$ - 0.0874 $\pm$ - 0.5238 $\pm$ - 0.1584 $\pm$ -
Table 15: Test results on HIV in the format of “metric mean $\pm$ standard deviation”.
ROC-AUC ECE NLL BS
DNN-rdkit Deterministic 0.7345 $\pm$ 0.0065 0.0077 $\pm$ 0.0010 0.1236 $\pm$ 0.0016 0.0276 $\pm$ 0.0006
Temperature 0.7345 $\pm$ 0.0065 0.0136 $\pm$ 0.0005 0.1277 $\pm$ 0.0021 0.0277 $\pm$ 0.0007
Focal Loss 0.7338 $\pm$ 0.0078 0.1205 $\pm$ 0.0179 0.2059 $\pm$ 0.0179 0.0429 $\pm$ 0.0042
MC Dropout 0.7365 $\pm$ 0.0088 0.0066 $\pm$ 0.0010 0.1219 $\pm$ 0.0016 0.0270 $\pm$ 0.0004
SWAG 0.7439 $\pm$ 0.0014 0.0105 $\pm$ 0.0025 0.1240 $\pm$ 0.0013 0.0279 $\pm$ 0.0005
BBP 0.7541 $\pm$ 0.0120 0.0181 $\pm$ 0.0065 0.1235 $\pm$ 0.0036 0.0277 $\pm$ 0.0011
SGLD 0.7414 $\pm$ 0.0019 0.0160 $\pm$ 0.0048 0.1313 $\pm$ 0.0061 0.0281 $\pm$ 0.0007
Ensembles 0.7613 $\pm$ - 0.0074 $\pm$ - 0.1160 $\pm$ - 0.0260 $\pm$ -
ChemBERTa Deterministic 0.7681 $\pm$ 0.0061 0.0214 $\pm$ 0.0033 0.1247 $\pm$ 0.0058 0.0292 $\pm$ 0.0019
Temperature 0.7681 $\pm$ 0.0061 0.0221 $\pm$ 0.0039 0.1259 $\pm$ 0.0067 0.0289 $\pm$ 0.0018
Focal Loss 0.7509 $\pm$ 0.0106 0.0651 $\pm$ 0.0351 0.1599 $\pm$ 0.0205 0.0343 $\pm$ 0.0025
MC Dropout 0.7717 $\pm$ 0.0042 0.0079 $\pm$ 0.0014 0.1105 $\pm$ 0.0011 0.0247 $\pm$ 0.0005
SWAG 0.7474 $\pm$ 0.0123 0.0263 $\pm$ 0.0045 0.1384 $\pm$ 0.0063 0.0319 $\pm$ 0.0017
BBP 0.7672 $\pm$ 0.0031 0.0217 $\pm$ 0.0043 0.1241 $\pm$ 0.0071 0.0289 $\pm$ 0.0027
SGLD 0.7224 $\pm$ 0.0095 0.0374 $\pm$ 0.0011 0.1602 $\pm$ 0.0106 0.0376 $\pm$ 0.0002
Ensembles 0.7885 $\pm$ - 0.0205 $\pm$ - 0.1161 $\pm$ - 0.0267 $\pm$ -
GROVER Deterministic 0.7979 $\pm$ 0.0045 0.0078 $\pm$ 0.0032 0.1133 $\pm$ 0.0019 0.0263 $\pm$ 0.0005
Temperature 0.7977 $\pm$ 0.0046 0.0120 $\pm$ 0.0011 0.1154 $\pm$ 0.0022 0.0265 $\pm$ 0.0004
Focal Loss 0.7746 $\pm$ 0.0113 0.1286 $\pm$ 0.0223 0.2156 $\pm$ 0.0259 0.0444 $\pm$ 0.0066
MC Dropout 0.8002 $\pm$ 0.0018 0.0077 $\pm$ 0.0028 0.1126 $\pm$ 0.0006 0.0261 $\pm$ 0.0001
SWAG 0.8021 $\pm$ 0.0033 0.0090 $\pm$ 0.0019 0.1118 $\pm$ 0.0018 0.0258 $\pm$ 0.0005
BBP 0.7928 $\pm$ 0.0116 0.0162 $\pm$ 0.0062 0.1151 $\pm$ 0.0034 0.0258 $\pm$ 0.0002
SGLD 0.7420 $\pm$ 0.0131 $0$$\pm$ 0.0012 0.1231 $\pm$ 0.0001 0.0272 $\pm$ 0.0000
Ensembles 0.8103 $\pm$ - 0.0035 $\pm$ - 0.1081 $\pm$ - 0.0251 $\pm$ -
Uni-Mol Deterministic 0.7913 $\pm$ 0.0042 0.0189 $\pm$ 0.0063 0.1335 $\pm$ 0.0225 0.0273 $\pm$ 0.0014
Temperature 0.7915 $\pm$ 0.0041 0.0151 $\pm$ 0.0017 0.1190 $\pm$ 0.0040 0.0269 $\pm$ 0.0009
Focal Loss 0.7785 $\pm$ 0.0073 0.0713 $\pm$ 0.0251 0.1561 $\pm$ 0.0205 0.0316 $\pm$ 0.0032
MC Dropout 0.7858 $\pm$ 0.0044 0.0229 $\pm$ 0.0091 0.1557 $\pm$ 0.0339 0.0294 $\pm$ 0.0033
SWAG 0.7786 $\pm$ 0.0061 0.0264 $\pm$ 0.0015 0.1676 $\pm$ 0.0160 0.0286 $\pm$ 0.0011
BBP 0.7781 $\pm$ 0.0128 0.0191 $\pm$ 0.0071 0.1428 $\pm$ 0.0327 0.0284 $\pm$ 0.0016
SGLD 0.$7$$\pm$ 0.0121 0.0205 $\pm$ 0.0006 0.1373 $\pm$ 0.0087 0.0263 $\pm$ 0.0002
Ensembles 0.8020 $\pm$ - 0.0181 $\pm$ - 0.1181 $\pm$ - 0.0265 $\pm$ -
TorchMD-NET Deterministic 0.7221 $\pm$ 0.0133 0.0270 $\pm$ 0.0031 0.2340 $\pm$ 0.0668 0.0315 $\pm$ 0.0007
Temperature 0.7221 $\pm$ 0.0133 0.0266 $\pm$ 0.0025 0.1532 $\pm$ 0.0024 0.0317 $\pm$ 0.0008
Focal Loss 0.7350 $\pm$ 0.0170 0.0673 $\pm$ 0.0312 0.1585 $\pm$ 0.0212 0.0332 $\pm$ 0.0036
MC Dropout 0.7225 $\pm$ 0.0098 0.0265 $\pm$ 0.0036 0.2191 $\pm$ 0.0590 0.0313 $\pm$ 0.0009
SWAG 0.7555 $\pm$ 0.0132 0.0253 $\pm$ 0.0007 0.2110 $\pm$ 0.0153 0.0288 $\pm$ 0.0006
BBP 0.7521 $\pm$ 0.0078 0.0195 $\pm$ 0.0114 0.1384 $\pm$ 0.0251 0.0290 $\pm$ 0.0036
SGLD - $\pm$ - - $\pm$ - - $\pm$ - - $\pm$ -
Ensembles 0.7370 $\pm$ - 0.0218 $\pm$ - 0.1537 $\pm$ - 0.0284 $\pm$ -
GIN Deterministic 0.6893 $\pm$ 0.0214 - $\pm$ - 0.1715 $\pm$ 0.0040 0.0396 $\pm$ 0.0021
Temperature 0.6893 $\pm$ 0.0214 - $\pm$ - 0.1708 $\pm$ 0.0032 0.0378 $\pm$ 0.0024
Focal Loss 0.6703 $\pm$ 0.0204 - $\pm$ - 0.2283 $\pm$ 0.0151 0.0515 $\pm$ 0.0046
MC Dropout 0.6893 $\pm$ 0.0214 - $\pm$ - 0.1715 $\pm$ 0.0040 0.0396 $\pm$ 0.0021
SWAG 0.6863 $\pm$ 0.0305 - $\pm$ - 0.1496 $\pm$ 0.0091 0.0318 $\pm$ 0.0007
BBP 0.6947 $\pm$ 0.0106 - $\pm$ - 0.1625 $\pm$ 0.0055 0.0360 $\pm$ 0.0021
SGLD 0.6845 $\pm$ 0.0063 - $\pm$ - 0.1954 $\pm$ 0.0322 0.0362 $\pm$ 0.0026
Ensembles 0.7209 $\pm$ - - $\pm$ - 0.1492 $\pm$ - 0.0336 $\pm$ -
Table 16: Test results on MUV in the format of “metric mean $\pm$ standard deviation”.
ROC-AUC ECE NLL BS
DNN-rdkit Deterministic 0.6029 $\pm$ 0.0576 0.0015 $\pm$ 0.0003 0.0326 $\pm$ 0.0196 0.0026 $\pm$ 0.0001
Temperature 0.6059 $\pm$ 0.0538 - $\pm$ - 0.0321 $\pm$ 0.0145 0.0027 $\pm$ 0.0003
Focal Loss 0.6261 $\pm$ 0.0322 0.0780 $\pm$ 0.0033 0.0720 $\pm$ 0.0262 0.0076 $\pm$ 0.0032
MC Dropout 0.6282 $\pm$ 0.0530 0.0015 $\pm$ 0.0003 0.0256 $\pm$ 0.0114 0.0025 $\pm$ 0.0000
SWAG 0.5830 $\pm$ 0.0392 0.0025 $\pm$ 0.0006 0.0393 $\pm$ 0.0254 0.0027 $\pm$ 0.0001
BBP 0.5974 $\pm$ 0.0287 - $\pm$ - 0.0251 $\pm$ 0.0068 0.0027 $\pm$ 0.0001
SGLD 0.4955 $\pm$ 0.0509 0.2380 $\pm$ 0.0156 0.2798 $\pm$ 0.0201 0.0610 $\pm$ 0.0071
Ensembles 0.6736 $\pm$ - 0.0013 $\pm$ - 0.0175 $\pm$ - 0.0025 $\pm$ -
ChemBERTa Deterministic 0.6695 $\pm$ 0.0343 0.0030 $\pm$ 0.0001 0.0241 $\pm$ 0.0011 0.0029 $\pm$ 0.0000
Temperature 0.6695 $\pm$ 0.0343 0.0026 $\pm$ 0.0003 0.0196 $\pm$ 0.0010 0.0029 $\pm$ 0.0000
Focal Loss 0.7150 $\pm$ 0.0038 0.0208 $\pm$ 0.0119 0.0333 $\pm$ 0.0100 0.0040 $\pm$ 0.0007
MC Dropout 0.7168 $\pm$ 0.0100 0.0026 $\pm$ 0.0003 0.0230 $\pm$ 0.0027 0.0026 $\pm$ 0.0001
SWAG 0.6672 $\pm$ 0.0342 0.0031 $\pm$ 0.0000 0.0261 $\pm$ 0.0015 0.0029 $\pm$ 0.0000
BBP 0.5970 $\pm$ 0.0622 0.0026 $\pm$ 0.0010 0.0198 $\pm$ 0.0018 0.0030 $\pm$ 0.0004
SGLD 0.5030 $\pm$ 0.0346 0.0074 $\pm$ 0.0003 0.0217 $\pm$ 0.0002 0.0026 $\pm$ 0.0000
Ensembles 0.7224 $\pm$ - 0.0025 $\pm$ - 0.0197 $\pm$ - 0.0027 $\pm$ -
GROVER Deterministic 0.7157 $\pm$ 0.0330 0.0016 $\pm$ 0.0004 0.0169 $\pm$ 0.0004 0.0025 $\pm$ 0.0000
Temperature 0.7155 $\pm$ 0.0331 0.0014 $\pm$ 0.0001 0.0187 $\pm$ 0.0004 0.0025 $\pm$ 0.0000
Focal Loss 0.6825 $\pm$ 0.0127 0.0869 $\pm$ 0.0033 0.0995 $\pm$ 0.0036 0.0101 $\pm$ 0.0006
MC Dropout 0.7163 $\pm$ 0.0363 0.0016 $\pm$ 0.0004 0.0170 $\pm$ 0.0004 0.0025 $\pm$ 0.0000
SWAG 0.7281 $\pm$ 0.0226 0.0014 $\pm$ 0.0003 0.0162 $\pm$ 0.0002 0.0025 $\pm$ 0.0000
BBP 0.7490 $\pm$ 0.0068 0.0052 $\pm$ 0.0015 0.0192 $\pm$ 0.0011 0.0025 $\pm$ 0.0000
SGLD 0.5532 $\pm$ 0.0255 0.0375 $\pm$ 0.0009 0.0488 $\pm$ 0.0008 0.0039 $\pm$ 0.0001
Ensembles 0.7268 $\pm$ - 0.0014 $\pm$ - 0.0166 $\pm$ - 0.0025 $\pm$ -
Uni-Mol Deterministic 0.7718 $\pm$ 0.0200 0.0024 $\pm$ 0.0002 0.0225 $\pm$ 0.0027 0.0025 $\pm$ 0.0000
Temperature 0.7625 $\pm$ 0.0098 0.0017 $\pm$ 0.0000 0.0233 $\pm$ 0.0039 0.0025 $\pm$ 0.0001
Focal Loss 0.8239 $\pm$ 0.0140 0.0250 $\pm$ 0.0126 0.0354 $\pm$ 0.0116 0.0036 $\pm$ 0.0006
MC Dropout 0.7781 $\pm$ 0.0201 0.0026 $\pm$ 0.0001 0.0253 $\pm$ 0.0017 0.0025 $\pm$ 0.0000
SWAG 0.7429 $\pm$ 0.0442 0.0028 $\pm$ 0.0001 0.0276 $\pm$ 0.0016 0.0027 $\pm$ 0.0001
BBP 0.7394 $\pm$ 0.0140 0.0016 $\pm$ 0.0002 0.0169 $\pm$ 0.0002 0.0025 $\pm$ 0.0000
SGLD 0.5901 $\pm$ 0.0200 0.0014 $\pm$ 0.0000 0.0178 $\pm$ 0.0000 0.0025 $\pm$ 0.0000
Ensembles 0.7904 $\pm$ - 0.0023 $\pm$ - 0.0195 $\pm$ - 0.0025 $\pm$ -
TorchMD-NET Deterministic 0.7043 $\pm$ 0.0027 - $\pm$ - 0.0553 $\pm$ 0.0092 0.0027 $\pm$ 0.0001
Temperature 0.7073 $\pm$ 0.0016 - $\pm$ - 0.0318 $\pm$ 0.0059 0.0027 $\pm$ 0.0001
Focal Loss 0.7227 $\pm$ 0.0117 - $\pm$ - 0.0225 $\pm$ 0.0034 0.0028 $\pm$ 0.0002
MC Dropout 0.7003 $\pm$ 0.0057 - $\pm$ - 0.0455 $\pm$ 0.0074 0.0026 $\pm$ 0.0001
SWAG 0.7285 $\pm$ 0.0094 - $\pm$ - 0.0457 $\pm$ 0.0011 0.0029 $\pm$ 0.0001
BBP 0.7330 $\pm$ 0.0044 - $\pm$ - 0.0407 $\pm$ 0.0070 0.0029 $\pm$ 0.0002
SGLD - $\pm$ - - $\pm$ - - $\pm$ - - $\pm$ -
Ensembles 0.7045 $\pm$ - - $\pm$ - 0.0437 $\pm$ - 0.0025 $\pm$ -
GIN Deterministic 0.6996 $\pm$ 0.0094 0.0024 $\pm$ 0.0006 0.0224 $\pm$ 0.0034 0.0028 $\pm$ 0.0001
Temperature 0.6885 $\pm$ 0.0072 - $\pm$ - 0.0284 $\pm$ 0.0037 0.0028 $\pm$ 0.0002
Focal Loss 0.6964 $\pm$ 0.0121 0.0146 $\pm$ 0.0053 0.0280 $\pm$ 0.0042 0.0037 $\pm$ 0.0004
MC Dropout 0.6996 $\pm$ 0.0094 0.0024 $\pm$ 0.0006 0.0224 $\pm$ 0.0034 0.0028 $\pm$ 0.0001
SWAG 0.6898 $\pm$ 0.0132 0.0021 $\pm$ 0.0003 0.0226 $\pm$ 0.0030 0.0026 $\pm$ 0.0000
BBP 0.6425 $\pm$ 0.0114 0.0022 $\pm$ 0.0002 0.0229 $\pm$ 0.0009 0.0027 $\pm$ 0.0001
SGLD 0.5553 $\pm$ 0.0110 0.0063 $\pm$ 0.0007 $0$$\pm$ 0.0058 0.0029 $\pm$ 0.0001
Ensembles 0.7155 $\pm$ - 0.0021 $\pm$ - 0.0191 $\pm$ - 0.0026 $\pm$ -
Table 17: Test results on ESOL in the format of “metric mean $\pm$ standard deviation”.
RMSE MAE NLL CE
DNN-rdkit Deterministic 0.9048 $\pm$ 0.0289 0.6747 $\pm$ 0.0059 1.0033 $\pm$ 0.3765 0.0465 $\pm$ 0.0052
MC Dropout 0.9129 $\pm$ 0.0212 0.6853 $\pm$ 0.0049 1.0246 $\pm$ 0.3698 0.0495 $\pm$ 0.0051
SWAG 0.9165 $\pm$ 0.0244 0.6827 $\pm$ 0.0074 0.9461 $\pm$ 0.3428 0.0459 $\pm$ 0.0048
BBP 0.9557 $\pm$ 0.0298 0.7070 $\pm$ 0.0237 0.4306 $\pm$ 0.0325 0.0325 $\pm$ 0.0038
SGLD 0.9630 $\pm$ 0.0425 0.7175 $\pm$ 0.0302 0.5101 $\pm$ 0.0380 0.0162 $\pm$ 0.0017
Ensembles 0.8690 $\pm$ - 0.6455 $\pm$ - 0.5674 $\pm$ - 0.0453 $\pm$ -
ChemBERTa Deterministic 1.0044 $\pm$ 0.0358 0.7989 $\pm$ 0.0216 2.2803 $\pm$ 0.1218 $0$$\pm$ 0.0007
MC Dropout 0.9497 $\pm$ 0.0271 0.7620 $\pm$ 0.0176 1.4258 $\pm$ 0.1905 0.0481 $\pm$ 0.0023
SWAG 0.9966 $\pm$ 0.0334 0.7940 $\pm$ 0.0206 2.5815 $\pm$ 0.2470 0.0513 $\pm$ 0.0013
BBP 1.0125 $\pm$ 0.0113 0.8139 $\pm$ 0.0110 0.5990 $\pm$ 0.0499 0.0337 $\pm$ 0.0013
SGLD 0.9978 $\pm$ 0.0340 0.8083 $\pm$ 0.0225 1.2727 $\pm$ 0.0711 0.0464 $\pm$ 0.0013
Ensembles 0.9752 $\pm$ - 0.7763 $\pm$ - 1.4981 $\pm$ - 0.0470 $\pm$ -
GROVER Deterministic 0.9355 $\pm$ 0.0252 0.7396 $\pm$ 0.0193 0.9888 $\pm$ 0.0395 0.0451 $\pm$ 0.0006
MC Dropout 0.9437 $\pm$ 0.0246 0.7491 $\pm$ 0.0216 1.0065 $\pm$ 0.0351 0.0449 $\pm$ 0.0007
SWAG 0.9443 $\pm$ 0.0219 0.7420 $\pm$ 0.0189 1.2996 $\pm$ 0.1883 0.0470 $\pm$ 0.0011
BBP 0.9892 $\pm$ 0.0430 0.7821 $\pm$ 0.0327 0.5470 $\pm$ 0.0378 0.0153 $\pm$ 0.0028
SGLD 0.9955 $\pm$ 0.0104 0.7942 $\pm$ 0.0153 0.4786 $\pm$ 0.0140 0.0243 $\pm$ 0.0008
Ensembles 0.9110 $\pm$ - 0.7209 $\pm$ - 0.5854 $\pm$ - 0.0394 $\pm$ -
Uni-Mol Deterministic 0.8530 $\pm$ 0.0224 0.6809 $\pm$ 0.0164 1.0739 $\pm$ 0.1136 0.0527 $\pm$ 0.0004
MC Dropout 0.8367 $\pm$ 0.0237 0.6648 $\pm$ 0.0209 1.1098 $\pm$ 0.2625 0.0527 $\pm$ 0.0012
SWAG 0.8644 $\pm$ 0.0097 0.6950 $\pm$ 0.0042 1.3868 $\pm$ 0.1522 0.0540 $\pm$ 0.0003
BBP 0.8441 $\pm$ 0.0332 0.6719 $\pm$ 0.0279 0.3151 $\pm$ 0.0292 0.0329 $\pm$ 0.0015
SGLD 0.8037 $\pm$ 0.0121 0.6444 $\pm$ 0.0112 0.2685 $\pm$ 0.0187 0.0343 $\pm$ 0.0018
Ensembles 0.8181 $\pm$ - 0.6516 $\pm$ - 0.8372 $\pm$ - 0.0512 $\pm$ -
TorchMD-NET Deterministic 0.8772 $\pm$ 0.0638 0.6688 $\pm$ 0.0517 0.6817 $\pm$ 0.1947 0.0409 $\pm$ 0.0059
MC Dropout 0.8720 $\pm$ 0.0645 0.6660 $\pm$ 0.0510 0.6166 $\pm$ 0.1780 0.0402 $\pm$ 0.0063
SWAG 0.8569 $\pm$ 0.0581 0.6642 $\pm$ 0.0401 0.4144 $\pm$ 0.1136 0.0372 $\pm$ 0.0053
BBP 2.4684 $\pm$ 0.5232 1.9253 $\pm$ 0.3748 1.5283 $\pm$ 0.0753 0.0191 $\pm$ 0.0173
SGLD 0.9691 $\pm$ 0.0579 0.7843 $\pm$ 0.0439 0.5284 $\pm$ 0.0243 0.0143 $\pm$ 0.0016
Ensembles 1.0512 $\pm$ - 0.8394 $\pm$ - 0.9756 $\pm$ - 0.0005 $\pm$ -
GIN Deterministic 1.5513 $\pm$ 0.0639 1.2268 $\pm$ 0.0618 1.0705 $\pm$ 0.1566 0.0122 $\pm$ 0.0072
MC Dropout 1.5513 $\pm$ 0.0639 1.2268 $\pm$ 0.0618 1.0705 $\pm$ 0.1566 0.0122 $\pm$ 0.0072
SWAG 1.8390 $\pm$ 0.2814 1.4650 $\pm$ 0.2128 1.4527 $\pm$ 0.2933 0.0166 $\pm$ 0.0089
BBP 1.9791 $\pm$ 0.0920 1.5041 $\pm$ 0.0290 1.4189 $\pm$ 0.2794 0.0048 $\pm$ 0.0013
SGLD 3.3157 $\pm$ 2.1193 2.7240 $\pm$ 1.8086 1.6595 $\pm$ 0.4615 0.0424 $\pm$ 0.0276
Ensembles 1.3952 $\pm$ - 1.0846 $\pm$ - 0.9598 $\pm$ - 0.0025 $\pm$ -
Table 18: Test results on FreeSolv in the format of “metric mean $\pm$ standard deviation”.
RMSE MAE NLL CE
DNN-rdkit Deterministic 2.2403 $\pm$ 0.1366 1.6768 $\pm$ 0.1007 1.3370 $\pm$ 0.1120 0.0272 $\pm$ 0.0026
MC Dropout 2.2316 $\pm$ 0.1070 1.6469 $\pm$ 0.1122 1.3357 $\pm$ 0.1257 0.0247 $\pm$ 0.0015
SWAG 2.4516 $\pm$ 0.1188 1.7853 $\pm$ 0.1478 1.8461 $\pm$ 0.2920 0.0358 $\pm$ 0.0043
BBP 2.8456 $\pm$ 0.1416 2.1049 $\pm$ 0.1037 1.5285 $\pm$ 0.0469 0.0173 $\pm$ 0.0056
SGLD 2.5503 $\pm$ 0.1010 1.8549 $\pm$ 0.1092 1.4260 $\pm$ 0.0095 0.0137 $\pm$ 0.0033
Ensembles 1.9627 $\pm$ - 1.4359 $\pm$ - 1.0149 $\pm$ - 0.0240 $\pm$ -
ChemBERTa Deterministic 2.5267 $\pm$ 0.0864 1.7760 $\pm$ 0.0318 1.7859 $\pm$ 0.1916 0.0364 $\pm$ 0.0034
MC Dropout 2.5042 $\pm$ 0.0981 1.7804 $\pm$ 0.0456 1.7810 $\pm$ 0.1526 0.0359 $\pm$ 0.0033
SWAG 2.6269 $\pm$ 0.1627 1.8562 $\pm$ 0.1055 2.2239 $\pm$ 0.2879 0.0426 $\pm$ 0.0003
BBP 2.6340 $\pm$ 0.1368 1.9072 $\pm$ 0.1600 1.6143 $\pm$ 0.1930 0.0322 $\pm$ 0.0066
SGLD 2.6564 $\pm$ 0.0357 1.8633 $\pm$ 0.0344 1.7627 $\pm$ 0.0809 0.0347 $\pm$ 0.0053
Ensembles 2.3727 $\pm$ - 1.6904 $\pm$ - 1.5142 $\pm$ - 0.0341 $\pm$ -
GROVER Deterministic 2.1352 $\pm$ 0.0193 1.4770 $\pm$ 0.0303 2.8244 $\pm$ 0.8307 0.0428 $\pm$ 0.0034
MC Dropout 2.0981 $\pm$ 0.0069 1.4660 $\pm$ 0.0231 2.6575 $\pm$ 0.7926 0.0415 $\pm$ 0.0040
SWAG 2.2321 $\pm$ 0.0902 1.5041 $\pm$ 0.0460 3.3021 $\pm$ 0.0530 0.0469 $\pm$ 0.0008
BBP 2.6133 $\pm$ 0.0662 1.8288 $\pm$ 0.0843 1.4699 $\pm$ 0.0598 0.0126 $\pm$ 0.0014
SGLD 2.1117 $\pm$ 0.0985 1.3784 $\pm$ 0.0459 1.2791 $\pm$ 0.1411 0.0238 $\pm$ 0.0020
Ensembles 2.0238 $\pm$ - 1.3539 $\pm$ - 1.9474 $\pm$ - 0.0393 $\pm$ -
Uni-Mol Deterministic 1.7871 $\pm$ 0.1265 1.2906 $\pm$ 0.0483 2.5450 $\pm$ 1.2435 0.0521 $\pm$ 0.0018
MC Dropout 1.7585 $\pm$ 0.1322 1.2652 $\pm$ 0.0559 2.3425 $\pm$ 1.2327 0.0515 $\pm$ 0.0018
SWAG 1.8524 $\pm$ 0.1449 1.3481 $\pm$ 0.1063 2.9865 $\pm$ 1.1583 0.0545 $\pm$ 0.0013
BBP 1.7464 $\pm$ 0.0246 1.2424 $\pm$ 0.0049 0.9940 $\pm$ 0.0065 0.0289 $\pm$ 0.0027
SGLD 1.7462 $\pm$ 0.0908 1.2963 $\pm$ 0.0610 1.0358 $\pm$ 0.0881 0.0296 $\pm$ 0.0048
Ensembles 1.6716 $\pm$ - 1.1780 $\pm$ - 1.4306 $\pm$ - 0.0504 $\pm$ -
TorchMD-NET Deterministic 2.6436 $\pm$ 0.0969 1.9379 $\pm$ 0.1236 1.9251 $\pm$ 0.2705 0.0219 $\pm$ 0.0060
MC Dropout 2.6360 $\pm$ 0.0986 1.9475 $\pm$ 0.1212 1.8397 $\pm$ 0.2050 0.0212 $\pm$ 0.0052
SWAG 2.6997 $\pm$ 0.1009 1.8466 $\pm$ 0.0863 2.5681 $\pm$ 0.4816 0.0399 $\pm$ 0.0026
BBP 3.6709 $\pm$ 0.7927 2.7858 $\pm$ 0.6062 1.8484 $\pm$ 0.2682 0.0227 $\pm$ 0.0055
SGLD 2.7769 $\pm$ 0.1054 1.9696 $\pm$ 0.1725 1.5583 $\pm$ 0.1031 0.0155 $\pm$ 0.0025
Ensembles 2.9347 $\pm$ - 2.0826 $\pm$ - 1.6610 $\pm$ - 0.0138 $\pm$ -
GIN Deterministic 2.2009 $\pm$ 0.2750 1.7292 $\pm$ 0.2873 1.3182 $\pm$ 0.0981 0.0142 $\pm$ 0.0039
MC Dropout 2.2009 $\pm$ 0.2750 1.7292 $\pm$ 0.2873 1.3182 $\pm$ 0.0981 0.0142 $\pm$ 0.0039
SWAG 2.3232 $\pm$ 0.0818 1.8154 $\pm$ 0.1109 1.5179 $\pm$ 0.1044 0.0198 $\pm$ 0.0106
BBP 3.1290 $\pm$ 0.6024 2.4897 $\pm$ 0.5111 1.8594 $\pm$ 0.2348 0.0092 $\pm$ 0.0048
SGLD 2.7563 $\pm$ 0.6927 2.2125 $\pm$ 0.6421 1.6046 $\pm$ 0.2001 0.0116 $\pm$ 0.0093
Ensembles 2.0088 $\pm$ - 1.5676 $\pm$ - 1.5171 $\pm$ - 0.0005 $\pm$ -
Table 19: Test results on Lipophilicity in the format of “metric mean $\pm$ standard deviation”.
RMSE MAE NLL CE
DNN-rdkit Deterministic 0.7575 $\pm$ 0.0116 0.5793 $\pm$ 0.0138 0.6154 $\pm$ 0.2016 0.0293 $\pm$ 0.0030
MC Dropout 0.7559 $\pm$ 0.0110 0.5773 $\pm$ 0.0132 0.9071 $\pm$ 0.3721 0.0341 $\pm$ 0.0052
SWAG 0.7572 $\pm$ 0.0101 0.5823 $\pm$ 0.0129 0.7191 $\pm$ 0.2267 0.0308 $\pm$ 0.0027
BBP 0.7730 $\pm$ 0.0253 0.5938 $\pm$ 0.0245 0.7578 $\pm$ 0.1627 0.0305 $\pm$ 0.0046
SGLD 0.7468 $\pm$ 0.0018 0.5743 $\pm$ 0.0016 0.2152 $\pm$ 0.0097 0.0090 $\pm$ 0.0009
Ensembles 0.7172 $\pm$ - 0.5490 $\pm$ - 0.6165 $\pm$ - 0.0322 $\pm$ -
ChemBERTa Deterministic 0.7553 $\pm$ 0.0074 0.5910 $\pm$ 0.0065 1.2368 $\pm$ 0.3994 0.0362 $\pm$ 0.0032
MC Dropout 0.7142 $\pm$ 0.0063 0.5601 $\pm$ 0.0029 0.8178 $\pm$ 0.2701 0.0349 $\pm$ 0.0033
SWAG 0.7672 $\pm$ 0.0101 0.5992 $\pm$ 0.0080 1.5809 $\pm$ 0.4168 0.0395 $\pm$ 0.0021
BBP 0.7542 $\pm$ 0.0050 0.5869 $\pm$ 0.0087 0.4419 $\pm$ 0.0775 0.0279 $\pm$ 0.0023
SGLD 0.7622 $\pm$ 0.0045 0.5982 $\pm$ 0.0070 0.8719 $\pm$ 0.1359 0.0355 $\pm$ 0.0020
Ensembles 0.7367 $\pm$ - 0.5763 $\pm$ - 0.9756 $\pm$ - 0.0360 $\pm$ -
GROVER Deterministic 0.6316 $\pm$ 0.0066 0.4747 $\pm$ 0.0040 2.1512 $\pm$ 0.2203 0.0478 $\pm$ 0.0009
MC Dropout 0.6293 $\pm$ 0.0067 0.4740 $\pm$ 0.0028 2.0526 $\pm$ 0.2139 0.0476 $\pm$ 0.0009
SWAG 0.6317 $\pm$ 0.0075 0.4750 $\pm$ 0.0041 2.3980 $\pm$ 0.3597 0.0485 $\pm$ 0.0007
BBP 0.6481 $\pm$ 0.0063 0.5058 $\pm$ 0.0041 0.0789 $\pm$ 0.0206 0.0196 $\pm$ 0.0024
SGLD 0.6360 $\pm$ 0.0041 0.4984 $\pm$ 0.0017 0.0544 $\pm$ 0.0199 0.0215 $\pm$ 0.0010
Ensembles 0.6250 $\pm$ - 0.4693 $\pm$ - 1.6046 $\pm$ - 0.0460 $\pm$ -
Uni-Mol Deterministic 0.6079 $\pm$ 0.0032 0.4509 $\pm$ 0.0044 0.8975 $\pm$ 0.5565 0.0425 $\pm$ 0.0061
MC Dropout 0.5983 $\pm$ 0.0099 0.4438 $\pm$ 0.0117 1.3663 $\pm$ 1.2187 0.0440 $\pm$ 0.0084
SWAG 0.6026 $\pm$ 0.0004 0.4476 $\pm$ 0.0022 1.0101 $\pm$ 0.1639 0.0453 $\pm$ 0.0012
BBP 0.6044 $\pm$ 0.0016 0.4469 $\pm$ 0.0025 0.0679 $\pm$ 0.0218 0.0306 $\pm$ 0.0010
SGLD 0.6040 $\pm$ 0.0031 0.4554 $\pm$ 0.0042 0.1565 $\pm$ 0.0573 0.0329 $\pm$ 0.0017
Ensembles 0.5809 $\pm$ - 0.4266 $\pm$ - 0.6450 $\pm$ - 0.0438 $\pm$ -
TorchMD-NET Deterministic 0.8235 $\pm$ 0.2218 0.6462 $\pm$ 0.1853 4.3336 $\pm$ 2.6567 0.0447 $\pm$ 0.0127
MC Dropout 0.8887 $\pm$ 0.3142 0.7001 $\pm$ 0.2635 3.9796 $\pm$ 2.2502 0.0439 $\pm$ 0.0130
SWAG 0.6819 $\pm$ 0.0271 0.5288 $\pm$ 0.0225 2.2934 $\pm$ 1.5338 0.0394 $\pm$ 0.0144
BBP 2.8400 $\pm$ 2.2383 2.5087 $\pm$ 2.1497 1.6533 $\pm$ 0.9853 0.1008 $\pm$ 0.1219
SGLD 0.8608 $\pm$ 0.0679 0.7002 $\pm$ 0.0502 0.9818 $\pm$ 0.4045 0.0183 $\pm$ 0.0128
Ensembles 1.0313 $\pm$ - 0.8196 $\pm$ - 0.8619 $\pm$ - 0.0195 $\pm$ -
GIN Deterministic 0.9286 $\pm$ 0.1113 0.7455 $\pm$ 0.1056 0.4542 $\pm$ 0.1096 0.0048 $\pm$ 0.0055
MC Dropout 0.9286 $\pm$ 0.1113 0.7455 $\pm$ 0.1056 0.4542 $\pm$ 0.1096 0.0048 $\pm$ 0.0055
SWAG 1.1105 $\pm$ 0.2448 0.9010 $\pm$ 0.2179 0.6805 $\pm$ 0.2447 0.0147 $\pm$ 0.0090
BBP 1.0863 $\pm$ 0.0704 0.8606 $\pm$ 0.0492 0.6095 $\pm$ 0.0841 0.0018 $\pm$ 0.0010
SGLD 1.3804 $\pm$ 0.4327 1.1187 $\pm$ 0.3763 0.9884 $\pm$ 0.4025 0.0229 $\pm$ 0.0137
Ensembles 0.8071 $\pm$ - 0.6515 $\pm$ - 0.3241 $\pm$ - 0.0020 $\pm$ -
Table 20: Test results on QM7 in the format of “metric mean $\pm$ standard deviation”.
RMSE MAE NLL CE
DNN-rdkit Deterministic 142.9279 $\pm$ 2.9477 95.7200 $\pm$ 3.2512 10.2976 $\pm$ 1.0848 0.0393 $\pm$ 0.0036
MC Dropout 138.6817 $\pm$ 2.5873 90.7039 $\pm$ 3.1531 7.4627 $\pm$ 0.4138 0.0339 $\pm$ 0.0020
SWAG 143.6922 $\pm$ 1.9420 97.2590 $\pm$ 1.9750 11.3649 $\pm$ 1.4063 0.0418 $\pm$ 0.0041
BBP 141.5333 $\pm$ 4.8179 99.0430 $\pm$ 3.9287 6.8897 $\pm$ 0.8805 0.0330 $\pm$ 0.0064
SGLD 139.8857 $\pm$ 5.4764 99.4873 $\pm$ 2.0159 5.4071 $\pm$ 0.0338 0.0098 $\pm$ 0.0019
Ensembles 128.7182 $\pm$ - 84.7269 $\pm$ - 5.4568 $\pm$ - 0.0281 $\pm$ -
ChemBERTa Deterministic 114.1954 $\pm$ 1.9473 66.6504 $\pm$ 3.6637 6.9619 $\pm$ 1.4000 0.0251 $\pm$ 0.0059
MC Dropout 112.4931 $\pm$ 1.2013 64.7828 $\pm$ 3.4349 6.1942 $\pm$ 0.7998 0.0309 $\pm$ 0.0061
SWAG 114.2023 $\pm$ 1.5302 68.0992 $\pm$ 3.5122 6.9684 $\pm$ 1.3236 0.0236 $\pm$ 0.0052
BBP 116.2333 $\pm$ 1.8548 69.9783 $\pm$ 0.3732 5.5954 $\pm$ 0.2659 0.0262 $\pm$ 0.0034
SGLD 116.9667 $\pm$ 0.3664 74.6714 $\pm$ 1.3652 5.0906 $\pm$ 0.0281 0.0147 $\pm$ 0.0019
Ensembles 111.2764 $\pm$ - 63.8752 $\pm$ - 5.3896 $\pm$ - 0.0224 $\pm$ -
GROVER Deterministic 139.5969 $\pm$ 3.1922 98.6424 $\pm$ 4.8571 5.4936 $\pm$ 0.0870 0.0121 $\pm$ 0.0012
MC Dropout 136.1780 $\pm$ 2.9204 95.4940 $\pm$ 4.4117 5.4272 $\pm$ 0.0697 0.0107 $\pm$ 0.0014
SWAG 142.6412 $\pm$ 1.3203 107.1310 $\pm$ 1.6574 5.7041 $\pm$ 0.1256 0.0153 $\pm$ 0.0018
BBP 136.8576 $\pm$ 3.2036 99.8460 $\pm$ 2.2180 $5$$\pm$ 0.0135 0.0035 $\pm$ 0.0007
SGLD 159.7489 $\pm$ 2.9839 126.6406 $\pm$ 2.8867 5.5587 $\pm$ 0.0308 0.0072 $\pm$ 0.0009
Ensembles $1$$\pm$ - 89.4504 $\pm$ - 5.3241 $\pm$ - 0.0090 $\pm$ -
Uni-Mol Deterministic 103.9043 $\pm$ 1.3754 56.6287 $\pm$ 2.3360 4.6198 $\pm$ 0.2317 0.0338 $\pm$ 0.0065
MC Dropout 103.5853 $\pm$ 1.3201 54.8976 $\pm$ 1.3061 4.5158 $\pm$ 0.2036 0.0386 $\pm$ 0.0019
SWAG 107.8128 $\pm$ 0.1969 63.3542 $\pm$ 2.2114 5.5588 $\pm$ 0.6515 0.0412 $\pm$ 0.0026
BBP 101.5479 $\pm$ 1.1382 51.9550 $\pm$ 0.7905 4.5153 $\pm$ 0.0535 0.0263 $\pm$ 0.0022
SGLD 105.1573 $\pm$ 2.6264 57.0601 $\pm$ 4.5219 4.5642 $\pm$ 0.0453 0.0290 $\pm$ 0.0014
Ensembles 101.1805 $\pm$ - 52.0409 $\pm$ - 4.3732 $\pm$ - 0.0360 $\pm$ -
TorchMD-NET Deterministic 104.1467 $\pm$ 0.3144 49.6743 $\pm$ 0.7443 11.6606 $\pm$ 0.7198 0.0619 $\pm$ 0.0004
MC Dropout 104.1265 $\pm$ 0.3415 49.8274 $\pm$ 0.5136 10.8641 $\pm$ 0.6233 0.0609 $\pm$ 0.0003
SWAG 103.6753 $\pm$ 0.6387 50.1259 $\pm$ 1.4952 7.4456 $\pm$ 0.2718 0.0580 $\pm$ 0.0005
BBP 104.4771 $\pm$ 0.4558 51.7268 $\pm$ 0.6870 9.5890 $\pm$ 0.4097 0.0568 $\pm$ 0.0002
SGLD 112.7406 $\pm$ 3.9319 69.4583 $\pm$ 2.8874 5.1299 $\pm$ 0.0314 0.0212 $\pm$ 0.0022
Ensembles 102.0535 $\pm$ - 48.2234 $\pm$ - 10.6292 $\pm$ - 0.0618 $\pm$ -
GIN Deterministic 122.4303 $\pm$ 0.0690 76.2333 $\pm$ 1.0882 5.2490 $\pm$ 0.0500 0.0136 $\pm$ 0.0003
MC Dropout 122.4303 $\pm$ 0.0690 76.2333 $\pm$ 1.0882 5.2490 $\pm$ 0.0500 0.0136 $\pm$ 0.0003
SWAG 125.4810 $\pm$ 0.9210 83.3207 $\pm$ 1.3398 5.3423 $\pm$ 0.0619 0.0129 $\pm$ 0.0001
BBP 125.2453 $\pm$ 0.6326 85.2772 $\pm$ 1.4557 5.3348 $\pm$ 0.0059 0.0077 $\pm$ 0.0014
SGLD 132.1431 $\pm$ 6.9275 97.6270 $\pm$ 11.0966 5.3910 $\pm$ 0.0487 0.0126 $\pm$ 0.0011
Ensembles 118.0550 $\pm$ - 71.2196 $\pm$ - 5.1922 $\pm$ - 0.0105 $\pm$ -
Table 21: Test results on QM8 in the format of “metric mean $\pm$ standard deviation”.
RMSE MAE NLL CE
DNN-rdkit Deterministic 0.0356 $\pm$ 0.0004 0.0209 $\pm$ 0.0001 -0.7811 $\pm$ 2.1740 0.0282 $\pm$ 0.0008
MC Dropout 0.0352 $\pm$ 0.0002 0.0209 $\pm$ 0.0002 -2.7716 $\pm$ 0.1922 0.0267 $\pm$ 0.0001
SWAG 0.0360 $\pm$ 0.0002 0.0212 $\pm$ 0.0001 -0.7143 $\pm$ 1.5808 0.0289 $\pm$ 0.0005
BBP 0.0370 $\pm$ 0.0002 0.0214 $\pm$ 0.0001 -0.9807 $\pm$ 0.5485 0.0269 $\pm$ 0.0004
SGLD 0.0446 $\pm$ 0.0008 0.0294 $\pm$ 0.0005 -2.7053 $\pm$ 0.0184 0.0121 $\pm$ 0.0001
Ensembles 0.0345 $\pm$ - 0.0202 $\pm$ - -3.0118 $\pm$ - 0.0279 $\pm$ -
ChemBERTa Deterministic 0.0377 $\pm$ 0.0001 0.0208 $\pm$ 0.0000 3.8214 $\pm$ 1.1599 0.0356 $\pm$ 0.0001
MC Dropout 0.0370 $\pm$ 0.0002 0.0204 $\pm$ 0.0001 -0.6884 $\pm$ 0.1974 0.0332 $\pm$ 0.0001
SWAG 0.0379 $\pm$ 0.0002 0.0210 $\pm$ 0.0000 5.4548 $\pm$ 1.9400 0.0360 $\pm$ 0.0002
BBP 0.0358 $\pm$ 0.0003 0.0206 $\pm$ 0.0001 -1.5642 $\pm$ 0.0700 0.0316 $\pm$ 0.0005
SGLD 0.0362 $\pm$ 0.0002 0.0212 $\pm$ 0.0004 -3.1380 $\pm$ 0.0611 0.0169 $\pm$ 0.0007
Ensembles 0.0373 $\pm$ - 0.0205 $\pm$ - 0.9180 $\pm$ - 0.0349 $\pm$ -
GROVER Deterministic 0.0319 $\pm$ 0.0007 0.0174 $\pm$ 0.0003 0.0201 $\pm$ 1.4388 0.0386 $\pm$ 0.0020
MC Dropout 0.0318 $\pm$ 0.0006 0.0174 $\pm$ 0.0003 -0.8810 $\pm$ 1.1057 0.0381 $\pm$ 0.0021
SWAG 0.0323 $\pm$ 0.0004 0.0176 $\pm$ 0.0003 0.6655 $\pm$ 0.9582 0.0409 $\pm$ 0.0013
BBP 0.0324 $\pm$ 0.0004 0.0183 $\pm$ 0.0003 -3.3037 $\pm$ 0.0260 0.0199 $\pm$ 0.0005
SGLD 0.0340 $\pm$ 0.0007 0.0205 $\pm$ 0.0007 -3.1826 $\pm$ 0.0378 0.0157 $\pm$ 0.0006
Ensembles 0.0315 $\pm$ - 0.0171 $\pm$ - -1.1036 $\pm$ - 0.0383 $\pm$ -
Uni-Mol Deterministic 0.0334 $\pm$ 0.0006 0.0172 $\pm$ 0.0000 2.7759 $\pm$ 2.9972 0.0432 $\pm$ 0.0020
MC Dropout 0.0331 $\pm$ 0.0006 0.0171 $\pm$ 0.0001 0.8170 $\pm$ 2.8226 0.0412 $\pm$ 0.0023
SWAG 0.0338 $\pm$ 0.0004 0.0173 $\pm$ 0.0001 3.5568 $\pm$ 2.5309 0.0440 $\pm$ 0.0013
BBP 0.0309 $\pm$ 0.0007 0.0164 $\pm$ 0.0002 -2.4581 $\pm$ 0.4546 0.0316 $\pm$ 0.0023
SGLD $0$$\pm$ 0.0003 0.0163 $\pm$ 0.0001 -3.2399 $\pm$ 0.0908 0.0210 $\pm$ 0.0004
Ensembles 0.0331 $\pm$ - 0.0169 $\pm$ - -0.0541 $\pm$ - 0.0421 $\pm$ -
TorchMD-NET Deterministic 0.0324 $\pm$ 0.0006 0.0178 $\pm$ 0.0006 1.4661 $\pm$ 2.0676 0.0439 $\pm$ 0.0045
MC Dropout 0.0324 $\pm$ 0.0007 0.0179 $\pm$ 0.0006 0.2637 $\pm$ 1.5600 0.0419 $\pm$ 0.0042
SWAG 0.0319 $\pm$ 0.0005 0.0177 $\pm$ 0.0005 -1.1581 $\pm$ 0.9675 0.0403 $\pm$ 0.0019
BBP 0.0332 $\pm$ 0.0002 0.0176 $\pm$ 0.0001 4.4554 $\pm$ 0.7641 0.0469 $\pm$ 0.0007
SGLD 0.0354 $\pm$ 0.0003 0.0217 $\pm$ 0.0001 -2.9735 $\pm$ 0.0091 0.0091 $\pm$ 0.0003
Ensembles 0.0315 $\pm$ - 0.0173 $\pm$ - -1.9286 $\pm$ - 0.0419 $\pm$ -
GIN Deterministic 0.0344 $\pm$ 0.0004 0.0206 $\pm$ 0.0003 -3.3711 $\pm$ 0.0959 0.0271 $\pm$ 0.0008
MC Dropout 0.0344 $\pm$ 0.0004 0.0206 $\pm$ 0.0003 -3.3711 $\pm$ 0.0959 0.0271 $\pm$ 0.0008
SWAG 0.0352 $\pm$ 0.0005 0.0215 $\pm$ 0.0004 -3.3091 $\pm$ 0.0569 0.0277 $\pm$ 0.0003
BBP 0.0352 $\pm$ 0.0003 0.0220 $\pm$ 0.0001 -3.2994 $\pm$ 0.0296 0.0238 $\pm$ 0.0007
SGLD 0.0428 $\pm$ 0.0011 0.0290 $\pm$ 0.0009 -2.8866 $\pm$ 0.0191 0.0189 $\pm$ 0.0016
Ensembles 0.0341 $\pm$ - 0.0205 $\pm$ - -3.4906 $\pm$ - 0.0271 $\pm$ -
Table 22: Test results on QM9 in the format of “metric mean $\pm$ standard deviation”.
RMSE MAE NLL CE
DNN-rdkit Deterministic 0.0151 $\pm$ 0.0001 0.0101 $\pm$ 0.0001 -3.3798 $\pm$ 0.0592 0.0442 $\pm$ 0.0003
MC Dropout 0.0148 $\pm$ 0.0001 $0$$\pm$ 0.0001 -3.5269 $\pm$ 0.0258 0.0433 $\pm$ 0.0003
SWAG 0.0152 $\pm$ 0.0002 0.0102 $\pm$ 0.0001 -3.2845 $\pm$ 0.0814 0.0449 $\pm$ 0.0003
BBP 0.0153 $\pm$ 0.0002 0.0103 $\pm$ 0.0001 -3.3470 $\pm$ 0.0280 0.0445 $\pm$ 0.0004
SGLD 0.0196 $\pm$ 0.0003 0.0144 $\pm$ 0.0002 -3.3352 $\pm$ 0.0052 0.0070 $\pm$ 0.0003
Ensembles 0.0143 $\pm$ - 0.0096 $\pm$ - -3.6023 $\pm$ - 0.0436 $\pm$ -
ChemBERTa Deterministic 0.0146 $\pm$ 0.0004 0.0092 $\pm$ 0.0002 -2.4104 $\pm$ 0.0698 0.0547 $\pm$ 0.0002
MC Dropout 0.0141 $\pm$ 0.0002 0.0088 $\pm$ 0.0001 -3.1501 $\pm$ 0.1094 0.0513 $\pm$ 0.0005
SWAG 0.0148 $\pm$ 0.0003 0.0092 $\pm$ 0.0002 -2.1703 $\pm$ 0.0205 0.0553 $\pm$ 0.0001
BBP 0.0144 $\pm$ 0.0001 0.0093 $\pm$ 0.0001 -2.5932 $\pm$ 0.0665 0.0540 $\pm$ 0.0002
SGLD 0.0153 $\pm$ 0.0001 0.0101 $\pm$ 0.0002 -3.7586 $\pm$ 0.0106 0.0338 $\pm$ 0.0007
Ensembles 0.0140 $\pm$ - 0.0087 $\pm$ - -2.8766 $\pm$ - 0.0543 $\pm$ -
GROVER Deterministic 0.0115 $\pm$ 0.0000 0.0068 $\pm$ $0$0.7874 $\pm$ 0.5323 0.0621 $\pm$ 0.0008
MC Dropout 0.0114 $\pm$ 0.0000 0.0068 $\pm$ $0$1.1002 $\pm$ 0.5003 0.0616 $\pm$ 0.0008
SWAG 0.0116 $\pm$ 0.0000 0.0068 $\pm$ $0$0.4778 $\pm$ 0.1923 0.0625 $\pm$ 0.0002
BBP 0.0118 $\pm$ 0.0000 0.0070 $\pm$ $0$1.8855 $\pm$ 0.1189 0.0591 $\pm$ 0.0004
SGLD 0.0136 $\pm$ 0.0006 0.0088 $\pm$ 0.0003 -3.7855 $\pm$ 0.0411 0.0291 $\pm$ 0.0017
Ensembles 0.0114 $\pm$ - 0.0067 $\pm$ - -1.0286 $\pm$ - 0.0620 $\pm$ -
Uni-Mol Deterministic 0.0096 $\pm$ 0.0001 0.0054 $\pm$ 0.0001 0.0143 $\pm$ 0.4195 0.0664 $\pm$ 0.0003
MC Dropout 0.0096 $\pm$ 0.0001 0.0054 $\pm$ 0.0001 -0.2513 $\pm$ 0.3961 0.0661 $\pm$ 0.0003
SWAG 0.0097 $\pm$ 0.0000 0.0054 $\pm$ $0$0.4624 $\pm$ 0.0619 0.0660 $\pm$ 0.0001
BBP 0.0095 $\pm$ 0.0002 0.0054 $\pm$ $0$2.9595 $\pm$ 0.1250 0.0618 $\pm$ 0.0003
SGLD 0.0095 $\pm$ 0.0001 0.0055 $\pm$ 0.0001 -4.2093 $\pm$ 0.0049 0.0459 $\pm$ 0.0002
Ensembles 0.0095 $\pm$ - 0.0053 $\pm$ - -0.3192 $\pm$ - 0.0663 $\pm$ -
TorchMD-NET Deterministic 0.0087 $\pm$ 0.0001 0.0048 $\pm$ 0.0000 3.6334 $\pm$ 1.0190 0.0690 $\pm$ 0.0003
MC Dropout 0.0087 $\pm$ 0.0001 0.0048 $\pm$ 0.0000 3.1525 $\pm$ 1.0291 0.0687 $\pm$ 0.0003
SWAG 0.0087 $\pm$ 0.0000 0.0048 $\pm$ $0$0.7161 $\pm$ 0.1242 0.0670 $\pm$ 0.0001
BBP 0.0083 $\pm$ 0.0002 0.0046 $\pm$ 0.0001 1.7970 $\pm$ 1.8061 0.0685 $\pm$ 0.0010
SGLD 0.0113 $\pm$ 0.0006 0.0075 $\pm$ 0.0004 -3.6560 $\pm$ 0.0469 0.0132 $\pm$ 0.0013
Ensembles 0.0086 $\pm$ - 0.0046 $\pm$ - 2.2619 $\pm$ - 0.0687 $\pm$ -
GIN Deterministic 0.0132 $\pm$ 0.0001 0.0084 $\pm$ 0.0001 -3.4286 $\pm$ 0.0389 0.0497 $\pm$ 0.0003
MC Dropout 0.0132 $\pm$ 0.0001 0.0084 $\pm$ 0.0001 -3.4286 $\pm$ 0.0389 0.0497 $\pm$ 0.0003
SWAG 0.0136 $\pm$ 0.0000 0.0088 $\pm$ 0.0001 -3.3355 $\pm$ 0.0589 0.0495 $\pm$ 0.0005
BBP 0.0131 $\pm$ 0.0000 0.0084 $\pm$ $0$3.3469 $\pm$ 0.0650 0.0499 $\pm$ 0.0003
SGLD 0.0161 $\pm$ 0.0003 0.0115 $\pm$ 0.0004 -3.5925 $\pm$ 0.0198 0.0211 $\pm$ 0.0031
Ensembles 0.0130 $\pm$ - 0.0081 $\pm$ - -3.5207 $\pm$ - $0$$\pm$ -
Table 23: Test results on the randomly split BACE dataset.
ROC-AUC ECE NLL BS
DNN-rdkit Deterministic 0.8850 $\pm$ 0.0220 0.1123 $\pm$ 0.0233 0.4820 $\pm$ 0.0723 0.1380 $\pm$ 0.0266
Temperature 0.8850 $\pm$ 0.0220 0.1022 $\pm$ 0.0019 0.4475 $\pm$ 0.0624 0.1359 $\pm$ 0.0231
Focal Loss 0.8884 $\pm$ 0.0167 0.1351 $\pm$ 0.0228 0.4623 $\pm$ 0.0239 0.1471 $\pm$ 0.0072
MC Dropout 0.8910 $\pm$ 0.0245 0.0969 $\pm$ 0.0134 0.4328 $\pm$ 0.0594 0.1344 $\pm$ 0.0215
SWAG 0.8898 $\pm$ 0.0220 0.1019 $\pm$ 0.0070 0.4346 $\pm$ 0.0525 0.1346 $\pm$ 0.0211
BBP 0.8784 $\pm$ 0.0133 0.1101 $\pm$ 0.0256 0.4954 $\pm$ 0.0755 0.1414 $\pm$ 0.0156
SGLD 0.8830 $\pm$ 0.0214 0.1119 $\pm$ 0.0294 0.4706 $\pm$ 0.0793 0.1360 $\pm$ 0.0244
Ensembles 0.8872 $\pm$ 0.0223 0.0972 $\pm$ 0.0101 0.4572 $\pm$ 0.0712 0.1308 $\pm$ 0.0252
ChemBERTa Deterministic 0.8835 $\pm$ 0.0044 0.1240 $\pm$ 0.0226 0.5182 $\pm$ 0.0812 0.1492 $\pm$ 0.0072
Temperature 0.8835 $\pm$ 0.0044 0.1077 $\pm$ 0.0201 0.4566 $\pm$ 0.0274 0.1427 $\pm$ 0.0029
Focal Loss 0.8884 $\pm$ 0.0096 0.1371 $\pm$ 0.0357 0.4490 $\pm$ 0.0328 0.1437 $\pm$ 0.0110
MC Dropout 0.8825 $\pm$ 0.0124 0.0978 $\pm$ 0.0099 0.4544 $\pm$ 0.0082 0.1447 $\pm$ 0.0073
SWAG 0.8868 $\pm$ 0.0059 0.1248 $\pm$ 0.0265 0.5286 $\pm$ 0.0964 0.1479 $\pm$ 0.0103
BBP 0.8918 $\pm$ 0.0057 0.1230 $\pm$ 0.0362 0.5192 $\pm$ 0.0661 0.1468 $\pm$ 0.0117
SGLD 0.8846 $\pm$ 0.0092 0.1362 $\pm$ 0.0046 0.5406 $\pm$ 0.0774 0.1438 $\pm$ 0.0073
Ensembles 0.8959 $\pm$ 0.0058 0.1219 $\pm$ 0.0141 0.4715 $\pm$ 0.0441 0.1370 $\pm$ 0.0071
GROVER Deterministic 0.9067 $\pm$ 0.0148 0.1003 $\pm$ 0.0139 0.4230 $\pm$ 0.0372 0.1254 $\pm$ 0.0114
Temperature 0.9067 $\pm$ 0.0148 0.0884 $\pm$ 0.0130 0.4049 $\pm$ 0.0351 0.1237 $\pm$ 0.0113
Focal Loss 0.9084 $\pm$ 0.0154 0.1600 $\pm$ 0.0069 0.4651 $\pm$ 0.0323 0.1492 $\pm$ 0.0137
MC Dropout 0.9078 $\pm$ 0.0168 0.0802 $\pm$ 0.0069 0.4073 $\pm$ 0.0394 0.1238 $\pm$ 0.0119
SWAG 0.9112 $\pm$ 0.0141 0.0887 $\pm$ 0.0260 0.3916 $\pm$ 0.0351 0.1203 $\pm$ 0.0112
BBP 0.9089 $\pm$ 0.0065 0.0807 $\pm$ 0.0112 0.3946 $\pm$ 0.0178 0.1209 $\pm$ 0.0103
SGLD 0.9066 $\pm$ 0.0145 0.0852 $\pm$ 0.0101 0.4029 $\pm$ 0.0313 0.1231 $\pm$ 0.0097
Ensembles 0.9046 $\pm$ 0.0133 0.0915 $\pm$ 0.0054 0.4207 $\pm$ 0.0289 0.1248 $\pm$ 0.0104
Uni-Mol Deterministic 0.9056 $\pm$ 0.0204 0.0894 $\pm$ 0.0129 0.4081 $\pm$ 0.0529 0.1257 $\pm$ 0.0191
Temperature 0.9059 $\pm$ 0.0205 0.0854 $\pm$ 0.0082 0.4060 $\pm$ 0.0491 0.1261 $\pm$ 0.0180
Focal Loss 0.8939 $\pm$ 0.0206 0.1634 $\pm$ 0.0148 0.5027 $\pm$ 0.0139 0.1655 $\pm$ 0.0065
MC Dropout 0.9061 $\pm$ 0.0199 0.0864 $\pm$ 0.0086 0.4069 $\pm$ 0.0514 0.1256 $\pm$ 0.0188
SWAG 0.9105 $\pm$ 0.0142 0.0988 $\pm$ 0.0122 0.3822 $\pm$ 0.0340 0.1179 $\pm$ 0.0131
BBP 0.8990 $\pm$ 0.0154 0.0867 $\pm$ 0.0206 0.4303 $\pm$ 0.0513 0.1309 $\pm$ 0.0184
SGLD 0.9080 $\pm$ 0.0104 0.0834 $\pm$ 0.0139 0.3949 $\pm$ 0.0317 0.1223 $\pm$ 0.0138
Ensembles 0.9023 $\pm$ 0.0189 0.1020 $\pm$ 0.0140 0.4040 $\pm$ 0.0491 0.1272 $\pm$ 0.0220
Table 24: Test results on the randomly split BBBP dataset.
ROC-AUC ECE NLL BS
DNN-rdkit Deterministic 0.9215 $\pm$ 0.0254 0.0838 $\pm$ 0.0047 0.3722 $\pm$ 0.0275 0.0918 $\pm$ 0.0052
Temperature 0.9215 $\pm$ 0.0254 0.0711 $\pm$ 0.0035 0.3131 $\pm$ 0.0141 0.0878 $\pm$ 0.0063
Focal Loss 0.9180 $\pm$ 0.0142 0.1214 $\pm$ 0.0528 0.3506 $\pm$ 0.0599 0.1038 $\pm$ 0.0188
MC Dropout 0.9226 $\pm$ 0.0244 0.0683 $\pm$ 0.0007 0.3156 $\pm$ 0.0151 0.0864 $\pm$ 0.0079
SWAG 0.9205 $\pm$ 0.0234 0.0816 $\pm$ 0.0065 0.3358 $\pm$ 0.0247 0.0876 $\pm$ 0.0081
BBP 0.9238 $\pm$ 0.0149 0.0846 $\pm$ 0.0188 0.3752 $\pm$ 0.0675 0.0932 $\pm$ 0.0113
SGLD 0.9172 $\pm$ 0.0284 0.0865 $\pm$ 0.0125 0.3930 $\pm$ 0.0735 0.0905 $\pm$ 0.0072
Ensembles 0.9226 $\pm$ 0.0088 $0$$\pm$ 0.0120 0.3324 $\pm$ 0.0220 0.0856 $\pm$ 0.0066
ChemBERTa Deterministic 0.9605 $\pm$ 0.0034 0.0686 $\pm$ 0.0066 0.2494 $\pm$ 0.0089 0.0731 $\pm$ 0.0030
Temperature 0.9605 $\pm$ 0.0034 0.0591 $\pm$ 0.0013 0.2275 $\pm$ 0.0111 0.0697 $\pm$ 0.0040
Focal Loss 0.9587 $\pm$ 0.0027 0.0816 $\pm$ 0.0262 0.2645 $\pm$ 0.0437 0.0795 $\pm$ 0.0133
MC Dropout 0.9607 $\pm$ 0.0027 0.0615 $\pm$ 0.0094 0.2438 $\pm$ 0.0193 0.0717 $\pm$ 0.0018
SWAG 0.9588 $\pm$ 0.0018 0.0737 $\pm$ 0.0041 0.2856 $\pm$ 0.0097 0.0780 $\pm$ 0.0039
BBP 0.9590 $\pm$ 0.0011 0.0702 $\pm$ 0.0125 0.2979 $\pm$ 0.0891 0.0766 $\pm$ 0.0052
SGLD 0.9580 $\pm$ 0.0033 0.0854 $\pm$ 0.0057 0.3513 $\pm$ 0.0324 0.0826 $\pm$ 0.0055
Ensembles 0.9604 $\pm$ 0.0018 0.0691 $\pm$ 0.0124 0.2480 $\pm$ 0.0045 0.0716 $\pm$ 0.0028
GROVER Deterministic 0.9300 $\pm$ 0.0172 0.0551 $\pm$ 0.0138 0.2746 $\pm$ 0.0256 0.0742 $\pm$ 0.0075
Temperature 0.9300 $\pm$ 0.0171 0.0575 $\pm$ 0.0185 0.2689 $\pm$ 0.0252 0.0740 $\pm$ 0.0074
Focal Loss 0.9294 $\pm$ 0.0198 0.1944 $\pm$ 0.0271 0.3963 $\pm$ 0.0258 0.1153 $\pm$ 0.0102
MC Dropout 0.9311 $\pm$ 0.0171 0.0549 $\pm$ 0.0115 0.2701 $\pm$ 0.0242 0.0742 $\pm$ 0.0072
SWAG 0.9324 $\pm$ 0.0219 0.0601 $\pm$ 0.0112 0.2736 $\pm$ 0.0300 0.0741 $\pm$ 0.0097
BBP 0.9196 $\pm$ 0.0210 0.0549 $\pm$ 0.0103 0.2712 $\pm$ 0.0225 0.0751 $\pm$ 0.0084
SGLD 0.9306 $\pm$ 0.0182 0.0588 $\pm$ 0.0140 0.2750 $\pm$ 0.0304 0.0757 $\pm$ 0.0092
Ensembles 0.9320 $\pm$ 0.0141 0.0633 $\pm$ 0.0124 0.2739 $\pm$ 0.0305 0.0738 $\pm$ 0.0092
Uni-Mol Deterministic 0.9285 $\pm$ 0.0126 0.0785 $\pm$ 0.0103 0.3165 $\pm$ 0.0332 0.0810 $\pm$ 0.0108
Temperature 0.9272 $\pm$ 0.0121 0.0687 $\pm$ 0.0094 0.2838 $\pm$ 0.0231 0.0790 $\pm$ 0.0099
Focal Loss 0.9242 $\pm$ 0.0086 0.1001 $\pm$ 0.0261 0.3201 $\pm$ 0.0332 0.0936 $\pm$ 0.0109
MC Dropout 0.9284 $\pm$ 0.0125 0.0750 $\pm$ 0.0087 0.3124 $\pm$ 0.0322 0.0807 $\pm$ 0.0107
SWAG 0.9359 $\pm$ 0.0048 0.0797 $\pm$ 0.0086 0.3233 $\pm$ 0.0266 0.0792 $\pm$ 0.0091
BBP 0.9175 $\pm$ 0.0121 0.0706 $\pm$ 0.0047 0.3256 $\pm$ 0.0262 0.0809 $\pm$ 0.0094
SGLD 0.9234 $\pm$ 0.0069 0.0869 $\pm$ 0.0053 0.3671 $\pm$ 0.0262 0.0846 $\pm$ 0.0109
Ensembles 0.9341 $\pm$ 0.0117 0.0652 $\pm$ 0.0140 0.2925 $\pm$ 0.0429 0.0766 $\pm$ 0.0107
Table 25: Test results on the randomly split ClinTox dataset.
ROC-AUC ECE NLL BS
DNN-rdkit Deterministic 0.8659 $\pm$ 0.0483 0.0754 $\pm$ 0.0129 0.5589 $\pm$ 0.4190 0.0672 $\pm$ 0.0104
Temperature 0.8628 $\pm$ 0.0300 0.0719 $\pm$ 0.0087 0.4434 $\pm$ 0.1979 0.0645 $\pm$ 0.0068
Focal Loss 0.8813 $\pm$ 0.0564 0.0652 $\pm$ 0.0054 0.1980 $\pm$ 0.0696 0.0523 $\pm$ 0.0093
MC Dropout 0.8757 $\pm$ 0.0382 0.0616 $\pm$ 0.0093 0.2560 $\pm$ 0.0963 0.0566 $\pm$ 0.0067
SWAG 0.8819 $\pm$ 0.0424 0.0619 $\pm$ 0.0070 0.2574 $\pm$ 0.1058 0.0572 $\pm$ 0.0066
BBP 0.8358 $\pm$ 0.0778 0.0547 $\pm$ 0.0198 0.2679 $\pm$ 0.1254 0.0534 $\pm$ 0.0083
SGLD 0.8745 $\pm$ 0.0506 0.0666 $\pm$ 0.0085 0.2488 $\pm$ 0.0714 0.0572 $\pm$ 0.0068
Ensembles 0.8241 $\pm$ 0.0587 0.0646 $\pm$ 0.0088 0.4562 $\pm$ 0.3799 0.0561 $\pm$ 0.0077
ChemBERTa Deterministic 0.9862 $\pm$ 0.0063 0.0193 $\pm$ 0.0026 0.0703 $\pm$ 0.0063 0.0168 $\pm$ 0.0015
Temperature 0.9862 $\pm$ 0.0063 0.0181 $\pm$ 0.0045 0.0639 $\pm$ 0.0039 0.0163 $\pm$ 0.0013
Focal Loss 0.9824 $\pm$ 0.0038 0.0493 $\pm$ 0.0183 0.0907 $\pm$ 0.0132 0.0193 $\pm$ 0.0025
MC Dropout 0.9870 $\pm$ 0.0051 0.0162 $\pm$ 0.0029 0.0640 $\pm$ 0.0087 0.0148 $\pm$ 0.0025
SWAG 0.9863 $\pm$ 0.0060 0.0216 $\pm$ 0.0036 0.0751 $\pm$ 0.0084 0.0179 $\pm$ 0.0019
BBP 0.9855 $\pm$ 0.0077 0.0268 $\pm$ 0.0124 0.0687 $\pm$ 0.0095 0.0165 $\pm$ 0.0021
SGLD 0.9851 $\pm$ 0.0053 0.0201 $\pm$ 0.0058 0.0713 $\pm$ 0.0020 0.0178 $\pm$ 0.0035
Ensembles 0.9885 $\pm$ 0.0034 0.0242 $\pm$ 0.0108 0.0633 $\pm$ 0.0056 0.0154 $\pm$ 0.0015
GROVER Deterministic 0.8823 $\pm$ 0.0347 0.0392 $\pm$ 0.0059 0.1550 $\pm$ 0.0289 0.0409 $\pm$ 0.0069
Temperature 0.8818 $\pm$ 0.0341 0.0387 $\pm$ 0.0041 0.1575 $\pm$ 0.0294 0.0412 $\pm$ 0.0069
Focal Loss 0.8554 $\pm$ 0.0542 0.1687 $\pm$ 0.0024 0.2767 $\pm$ 0.0073 0.0682 $\pm$ 0.0026
MC Dropout 0.8840 $\pm$ 0.0332 0.0324 $\pm$ 0.0015 0.1505 $\pm$ 0.0242 0.0392 $\pm$ 0.0058
SWAG 0.8897 $\pm$ 0.0337 $0$$\pm$ 0.0033 0.1490 $\pm$ 0.0267 0.0397 $\pm$ 0.0060
BBP 0.8466 $\pm$ 0.0496 0.0684 $\pm$ 0.0059 0.1748 $\pm$ 0.0226 0.0419 $\pm$ 0.0064
SGLD 0.8904 $\pm$ 0.0285 0.0424 $\pm$ 0.0021 0.1545 $\pm$ 0.0297 0.0414 $\pm$ 0.0086
Ensembles 0.8854 $\pm$ 0.0371 0.0342 $\pm$ 0.0102 0.1470 $\pm$ 0.0295 0.0385 $\pm$ 0.0071
Uni-Mol Deterministic 0.8391 $\pm$ 0.0620 0.0625 $\pm$ 0.0119 0.1949 $\pm$ 0.0291 0.0531 $\pm$ 0.0078
Temperature 0.8433 $\pm$ 0.0611 0.0633 $\pm$ 0.0119 0.1880 $\pm$ 0.0236 0.0526 $\pm$ 0.0073
Focal Loss 0.8416 $\pm$ 0.0705 0.0961 $\pm$ 0.0199 0.2153 $\pm$ 0.0136 0.0586 $\pm$ 0.0050
MC Dropout 0.8394 $\pm$ 0.0616 0.0627 $\pm$ 0.0112 0.1931 $\pm$ 0.0281 0.0528 $\pm$ 0.0077
SWAG 0.8211 $\pm$ 0.0871 0.0515 $\pm$ 0.0131 0.1974 $\pm$ 0.0588 0.0484 $\pm$ 0.0097
BBP 0.8389 $\pm$ 0.0345 0.0593 $\pm$ 0.0234 0.1788 $\pm$ 0.0387 0.0472 $\pm$ 0.0132
SGLD 0.8367 $\pm$ 0.0855 0.0462 $\pm$ 0.0092 0.1867 $\pm$ 0.0655 0.0475 $\pm$ 0.0110
Ensembles 0.8387 $\pm$ 0.0463 0.0671 $\pm$ 0.0197 0.1803 $\pm$ 0.0254 0.0469 $\pm$ 0.0032
Table 26: Test results on the randomly split Tox21 dataset.
ROC-AUC ECE NLL BS
DNN-rdkit Deterministic 0.8451 $\pm$ 0.0162 0.0335 $\pm$ 0.0012 0.1999 $\pm$ 0.0041 0.0537 $\pm$ 0.0005
Temperature 0.8451 $\pm$ 0.0162 0.0338 $\pm$ 0.0023 0.1945 $\pm$ 0.0016 0.0534 $\pm$ 0.0005
Focal Loss 0.8422 $\pm$ 0.0120 0.0829 $\pm$ 0.0065 0.2265 $\pm$ 0.0055 0.0587 $\pm$ 0.0010
MC Dropout 0.8494 $\pm$ 0.0137 0.0259 $\pm$ 0.0009 0.1885 $\pm$ 0.0012 0.0510 $\pm$ 0.0002
SWAG 0.8496 $\pm$ 0.0129 0.0256 $\pm$ 0.0009 0.1882 $\pm$ 0.0008 0.0509 $\pm$ 0.0003
BBP 0.8216 $\pm$ 0.0033 0.0293 $\pm$ 0.0018 0.2052 $\pm$ 0.0054 0.0565 $\pm$ 0.0011
SGLD 0.8114 $\pm$ 0.0067 0.0864 $\pm$ 0.0038 0.2568 $\pm$ 0.0084 0.0685 $\pm$ 0.0028
Ensembles 0.8544 $\pm$ 0.0101 0.0271 $\pm$ 0.0025 0.1849 $\pm$ 0.0018 0.0506 $\pm$ 0.0005
ChemBERTa Deterministic 0.8373 $\pm$ 0.0140 0.0345 $\pm$ 0.0046 0.1993 $\pm$ 0.0110 0.0545 $\pm$ 0.0027
Temperature 0.8373 $\pm$ 0.0140 0.0360 $\pm$ 0.0020 0.1973 $\pm$ 0.0066 0.0543 $\pm$ 0.0022
Focal Loss 0.8339 $\pm$ 0.0128 0.0939 $\pm$ 0.0078 0.2361 $\pm$ 0.0024 0.0605 $\pm$ 0.0007
MC Dropout 0.8430 $\pm$ 0.0101 0.0240 $\pm$ 0.0017 0.1859 $\pm$ 0.0069 0.0506 $\pm$ 0.0019
SWAG 0.8369 $\pm$ 0.0149 0.0370 $\pm$ 0.0025 0.2027 $\pm$ 0.0110 0.0553 $\pm$ 0.0022
BBP 0.8301 $\pm$ 0.0131 0.0403 $\pm$ 0.0038 0.2052 $\pm$ 0.0077 0.0575 $\pm$ 0.0025
SGLD 0.8314 $\pm$ 0.0127 0.0376 $\pm$ 0.0018 0.2002 $\pm$ 0.0102 0.0552 $\pm$ 0.0031
Ensembles 0.8509 $\pm$ 0.0082 0.0304 $\pm$ 0.0027 0.1890 $\pm$ 0.0100 0.0519 $\pm$ 0.0025
GROVER Deterministic 0.8754 $\pm$ 0.0155 0.0259 $\pm$ 0.0018 0.1690 $\pm$ 0.0067 0.0457 $\pm$ 0.0015
Temperature 0.8754 $\pm$ 0.0154 0.0281 $\pm$ 0.0030 0.1700 $\pm$ 0.0060 0.0458 $\pm$ 0.0015
Focal Loss 0.8785 $\pm$ 0.0099 0.1337 $\pm$ 0.0012 0.2623 $\pm$ 0.0046 0.0636 $\pm$ 0.0017
MC Dropout 0.8760 $\pm$ 0.0149 0.0279 $\pm$ 0.0015 0.1697 $\pm$ 0.0062 0.0457 $\pm$ 0.0015
SWAG 0.8776 $\pm$ 0.0135 0.0247 $\pm$ 0.0014 0.1674 $\pm$ 0.0059 0.0452 $\pm$ 0.0014
BBP 0.8667 $\pm$ 0.0135 0.0439 $\pm$ 0.0038 0.1850 $\pm$ 0.0049 0.0483 $\pm$ 0.0016
SGLD 0.8724 $\pm$ 0.0123 0.0390 $\pm$ 0.0023 0.1811 $\pm$ 0.0061 0.0473 $\pm$ 0.0019
Ensembles 0.8775 $\pm$ 0.0137 0.0244 $\pm$ 0.0010 0.1680 $\pm$ 0.0055 0.0453 $\pm$ 0.0013
Uni-Mol Deterministic 0.8666 $\pm$ 0.0138 0.0302 $\pm$ 0.0016 0.1773 $\pm$ 0.0052 0.0466 $\pm$ 0.0012
Temperature 0.8661 $\pm$ 0.0136 0.0295 $\pm$ 0.0008 0.1727 $\pm$ 0.0037 0.0464 $\pm$ 0.0010
Focal Loss 0.8631 $\pm$ 0.0071 0.0824 $\pm$ 0.0093 0.2154 $\pm$ 0.0097 0.0543 $\pm$ 0.0028
MC Dropout 0.8667 $\pm$ 0.0139 0.0292 $\pm$ 0.0021 0.1758 $\pm$ 0.0050 0.0465 $\pm$ 0.0012
SWAG 0.8666 $\pm$ 0.0170 0.0308 $\pm$ 0.0018 0.1779 $\pm$ 0.0042 0.0452 $\pm$ 0.0010
BBP 0.8731 $\pm$ 0.0107 0.0259 $\pm$ 0.0031 0.1731 $\pm$ 0.0046 0.0469 $\pm$ 0.0013
SGLD 0.8663 $\pm$ 0.0149 0.0333 $\pm$ 0.0027 0.1807 $\pm$ 0.0058 0.0470 $\pm$ 0.0011
Ensembles 0.8778 $\pm$ 0.0090 0.0262 $\pm$ 0.0024 0.1660 $\pm$ 0.0043 0.0440 $\pm$ 0.0008
Table 27: Test results on the randomly split ESOL dataset.
RMSE MAE NLL CE
DNN-rdkit Deterministic 0.6530 $\pm$ 0.0615 0.4788 $\pm$ 0.0279 0.1067 $\pm$ 0.0716 0.0419 $\pm$ 0.0034
MC Dropout 0.6597 $\pm$ 0.0666 0.4820 $\pm$ 0.0351 0.1212 $\pm$ 0.0925 0.0436 $\pm$ 0.0037
SWAG 0.6542 $\pm$ 0.0766 0.4752 $\pm$ 0.0392 0.0977 $\pm$ 0.0884 0.0435 $\pm$ 0.0034
BBP 0.7060 $\pm$ 0.0604 0.5294 $\pm$ 0.0458 0.2153 $\pm$ 0.1182 0.0263 $\pm$ 0.0050
SGLD 0.7033 $\pm$ 0.0648 0.5235 $\pm$ 0.0273 0.3825 $\pm$ 0.0019 0.0163 $\pm$ 0.0019
Ensembles 0.6455 $\pm$ 0.0518 0.4735 $\pm$ 0.0310 0.0791 $\pm$ 0.0905 0.0405 $\pm$ 0.0031
ChemBERTa Deterministic 0.7760 $\pm$ 0.0921 0.5920 $\pm$ 0.0465 0.3768 $\pm$ 0.1070 0.0464 $\pm$ 0.0042
MC Dropout 0.6919 $\pm$ 0.0947 0.5192 $\pm$ 0.0619 0.1908 $\pm$ 0.1299 0.0448 $\pm$ 0.0043
SWAG 0.7704 $\pm$ 0.0784 0.5921 $\pm$ 0.0350 0.4094 $\pm$ 0.1176 0.0475 $\pm$ 0.0046
BBP 0.7648 $\pm$ 0.0328 0.5923 $\pm$ 0.0176 0.2208 $\pm$ 0.0289 0.0326 $\pm$ 0.0022
SGLD 0.7282 $\pm$ 0.0764 0.5539 $\pm$ 0.0290 0.2455 $\pm$ 0.0835 0.0456 $\pm$ 0.0030
Ensembles 0.7318 $\pm$ 0.0778 0.5641 $\pm$ 0.0416 0.3073 $\pm$ 0.1107 0.0470 $\pm$ 0.0032
GROVER Deterministic 0.6137 $\pm$ 0.0392 0.4730 $\pm$ 0.0173 0.0299 $\pm$ 0.0935 0.0377 $\pm$ 0.0063
MC Dropout 0.5997 $\pm$ 0.0256 0.4621 $\pm$ 0.0110 0.0124 $\pm$ 0.0890 0.0376 $\pm$ 0.0065
SWAG 0.5895 $\pm$ 0.0251 0.4476 $\pm$ 0.0102 -0.0342 $\pm$ 0.0624 0.0419 $\pm$ 0.0035
BBP 0.6449 $\pm$ 0.0517 0.4892 $\pm$ 0.0286 0.3780 $\pm$ 0.0474 0.0152 $\pm$ 0.0018
SGLD 0.5888 $\pm$ 0.0379 0.4542 $\pm$ 0.0252 0.0748 $\pm$ 0.0683 0.0299 $\pm$ 0.0010
Ensembles 0.5847 $\pm$ 0.0326 0.4526 $\pm$ 0.0210 -0.0377 $\pm$ 0.0725 0.0384 $\pm$ 0.0031
Uni-Mol Deterministic 0.6253 $\pm$ 0.0500 0.4691 $\pm$ 0.0291 0.0024 $\pm$ 0.0693 0.0388 $\pm$ 0.0008
MC Dropout 0.6224 $\pm$ 0.0501 0.4640 $\pm$ 0.0299 0.0058 $\pm$ 0.0647 0.0375 $\pm$ 0.0007
SWAG 0.5642 $\pm$ 0.0564 0.4030 $\pm$ 0.0342 -0.0603 $\pm$ 0.1092 0.0361 $\pm$ 0.0028
BBP 0.6762 $\pm$ 0.0766 0.4884 $\pm$ 0.0544 0.2844 $\pm$ 0.0276 0.0235 $\pm$ 0.0024
SGLD 0.6063 $\pm$ 0.0281 0.4388 $\pm$ 0.0098 0.1947 $\pm$ 0.2311 0.0270 $\pm$ 0.0096
Ensembles 0.5930 $\pm$ 0.0374 0.4311 $\pm$ 0.0129 -0.0273 $\pm$ 0.0238 0.0375 $\pm$ 0.0015
Table 28: Test results on the randomly split FreeSolv dataset.
RMSE MAE NLL CE
DNN-rdkit Deterministic 1.1867 $\pm$ 0.0369 0.8582 $\pm$ 0.0521 0.7262 $\pm$ 0.0311 0.0252 $\pm$ 0.0038
MC Dropout 1.1893 $\pm$ 0.0716 0.8576 $\pm$ 0.0283 0.7085 $\pm$ 0.0277 0.0263 $\pm$ 0.0038
SWAG 1.1183 $\pm$ 0.0960 0.7817 $\pm$ 0.0602 0.6026 $\pm$ 0.0144 0.0305 $\pm$ 0.0047
BBP 1.8405 $\pm$ 0.2306 1.2685 $\pm$ 0.0533 1.2661 $\pm$ 0.0458 0.0060 $\pm$ 0.0013
SGLD 1.5395 $\pm$ 0.1830 1.0855 $\pm$ 0.0925 1.2214 $\pm$ 0.0395 0.0045 $\pm$ 0.0020
Ensembles 1.1690 $\pm$ 0.0630 0.8500 $\pm$ 0.0239 0.7444 $\pm$ 0.0502 0.0240 $\pm$ 0.0042
ChemBERTa Deterministic 1.4125 $\pm$ 0.1890 1.0323 $\pm$ 0.1206 0.8274 $\pm$ 0.1999 0.0387 $\pm$ 0.0059
MC Dropout 1.3675 $\pm$ 0.2456 1.0278 $\pm$ 0.1732 0.6970 $\pm$ 0.1730 0.0357 $\pm$ 0.0067
SWAG 1.3789 $\pm$ 0.2128 1.0024 $\pm$ 0.1570 0.8477 $\pm$ 0.2851 0.0408 $\pm$ 0.0041
BBP 1.4355 $\pm$ 0.2128 1.0643 $\pm$ 0.1312 0.8143 $\pm$ 0.0637 0.0270 $\pm$ 0.0032
SGLD 1.3864 $\pm$ 0.2747 0.9852 $\pm$ 0.1613 0.6654 $\pm$ 0.1475 0.0361 $\pm$ 0.0023
Ensembles 1.2276 $\pm$ 0.2395 0.8745 $\pm$ 0.1727 0.5995 $\pm$ 0.1546 0.0399 $\pm$ 0.0013
GROVER Deterministic 1.3823 $\pm$ 0.1108 1.0223 $\pm$ 0.0942 0.9105 $\pm$ 0.0684 0.0188 $\pm$ 0.0018
MC Dropout 1.3570 $\pm$ 0.1082 1.0093 $\pm$ 0.0940 0.9247 $\pm$ 0.0680 0.0177 $\pm$ 0.0017
SWAG 1.3181 $\pm$ 0.0898 0.9858 $\pm$ 0.0727 0.8386 $\pm$ 0.0352 0.0218 $\pm$ 0.0042
BBP 1.4832 $\pm$ 0.1322 1.1451 $\pm$ 0.0893 1.1459 $\pm$ 0.0166 0.0060 $\pm$ 0.0019
SGLD 1.4965 $\pm$ 0.2643 1.1682 $\pm$ 0.2289 1.1154 $\pm$ 0.1317 0.0072 $\pm$ 0.0016
Ensembles 1.2826 $\pm$ 0.1700 0.9424 $\pm$ 0.1316 0.9476 $\pm$ 0.0558 0.0140 $\pm$ 0.0020
Uni-Mol Deterministic 1.0429 $\pm$ 0.0170 0.7345 $\pm$ 0.0308 0.5741 $\pm$ 0.0637 0.0322 $\pm$ 0.0054
MC Dropout 1.0381 $\pm$ 0.0249 0.7280 $\pm$ 0.0394 0.5872 $\pm$ 0.0636 0.0310 $\pm$ 0.0055
SWAG 0.9959 $\pm$ 0.0902 0.6842 $\pm$ 0.1248 0.6050 $\pm$ 0.0705 0.0271 $\pm$ 0.0051
BBP 1.1541 $\pm$ 0.1779 0.8681 $\pm$ 0.1550 0.9795 $\pm$ 0.0621 0.0154 $\pm$ 0.0027
SGLD 1.0486 $\pm$ 0.1898 0.7543 $\pm$ 0.1538 1.0065 $\pm$ 0.1283 0.0126 $\pm$ 0.0016
Ensembles 1.0009 $\pm$ 0.0416 0.7201 $\pm$ 0.0378 0.6766 $\pm$ 0.0365 0.0270 $\pm$ 0.0045
Table 29: Test results on the randomly split Lipophilicity dataset.
RMSE MAE NLL CE
DNN-rdkit Deterministic 0.6790 $\pm$ 0.0415 0.5007 $\pm$ 0.0253 0.6158 $\pm$ 0.1691 0.0365 $\pm$ 0.0049
MC Dropout 0.6732 $\pm$ 0.0398 0.4971 $\pm$ 0.0286 0.5558 $\pm$ 0.1208 0.0366 $\pm$ 0.0048
SWAG 0.6699 $\pm$ 0.0441 0.4926 $\pm$ 0.0277 0.6208 $\pm$ 0.1477 0.0370 $\pm$ 0.0052
BBP 0.6833 $\pm$ 0.0444 0.5092 $\pm$ 0.0222 0.2681 $\pm$ 0.2026 0.0281 $\pm$ 0.0040
SGLD 0.6819 $\pm$ 0.0427 0.5090 $\pm$ 0.0222 0.1218 $\pm$ 0.0554 0.0134 $\pm$ 0.0010
Ensembles 0.6399 $\pm$ 0.0402 0.4690 $\pm$ 0.0199 0.3242 $\pm$ 0.1517 0.0349 $\pm$ 0.0004
ChemBERTa Deterministic 0.7033 $\pm$ 0.0521 0.5282 $\pm$ 0.0234 1.4475 $\pm$ 1.1982 0.0418 $\pm$ 0.0064
MC Dropout 0.6583 $\pm$ 0.0330 0.4882 $\pm$ 0.0095 0.8234 $\pm$ 0.7527 0.0380 $\pm$ 0.0056
SWAG 0.7114 $\pm$ 0.0502 0.5363 $\pm$ 0.0196 1.7413 $\pm$ 1.2705 0.0441 $\pm$ 0.0059
BBP 0.7103 $\pm$ 0.0508 0.5377 $\pm$ 0.0212 0.5764 $\pm$ 0.3161 0.0368 $\pm$ 0.0043
SGLD 0.6988 $\pm$ 0.0540 0.5271 $\pm$ 0.0249 0.7978 $\pm$ 0.2380 0.0414 $\pm$ 0.0016
Ensembles 0.6763 $\pm$ 0.0534 0.5044 $\pm$ 0.0232 1.4618 $\pm$ 0.8428 0.0451 $\pm$ 0.0033
GROVER Deterministic 0.5717 $\pm$ 0.0320 0.3957 $\pm$ 0.0144 0.6373 $\pm$ 0.0281 0.0446 $\pm$ 0.0016
MC Dropout 0.5633 $\pm$ 0.0349 0.3900 $\pm$ 0.0163 0.5493 $\pm$ 0.0500 0.0441 $\pm$ 0.0018
SWAG 0.5599 $\pm$ 0.0340 0.3851 $\pm$ 0.0117 0.5192 $\pm$ 0.1369 0.0440 $\pm$ 0.0002
BBP 0.5638 $\pm$ 0.0314 0.4021 $\pm$ 0.0088 -0.0783 $\pm$ 0.0532 0.0232 $\pm$ 0.0010
SGLD 0.5597 $\pm$ 0.0409 0.3958 $\pm$ 0.0171 -0.0915 $\pm$ 0.0856 0.0271 $\pm$ 0.0006
Ensembles 0.5539 $\pm$ 0.0314 0.3833 $\pm$ 0.0134 0.4942 $\pm$ 0.1247 0.0447 $\pm$ 0.0010
Uni-Mol Deterministic 0.5387 $\pm$ 0.0439 0.3824 $\pm$ 0.0203 -0.0690 $\pm$ 0.1128 0.0362 $\pm$ 0.0024
MC Dropout 0.5367 $\pm$ 0.0437 0.3812 $\pm$ 0.0204 -0.0859 $\pm$ 0.1088 0.0355 $\pm$ 0.0024
SWAG 0.5077 $\pm$ 0.0446 0.3598 $\pm$ 0.0178 -0.1168 $\pm$ 0.1292 0.0369 $\pm$ 0.0018
BBP 0.5632 $\pm$ 0.0509 0.4139 $\pm$ 0.0207 -0.0682 $\pm$ 0.0724 0.0248 $\pm$ 0.0039
SGLD 0.5238 $\pm$ 0.0357 0.3718 $\pm$ 0.0160 -0.1525 $\pm$ 0.0618 0.0265 $\pm$ 0.0011
Ensembles 0.5255 $\pm$ 0.0434 0.3706 $\pm$ 0.0187 -0.1311 $\pm$ 0.1122 0.0347 $\pm$ 0.0010
Table 30: Test results on the randomly split QM7 dataset.
RMSE MAE NLL CE
DNN-rdkit Deterministic 107.9546 $\pm$ 8.1501 57.5611 $\pm$ 4.8507 5.5231 $\pm$ 0.0677 0.0449 $\pm$ 0.0026
MC Dropout 106.5277 $\pm$ 8.4035 56.8402 $\pm$ 4.4571 4.9801 $\pm$ 0.0805 0.0434 $\pm$ 0.0025
SWAG 105.3930 $\pm$ 8.3626 56.1532 $\pm$ 3.9650 4.9387 $\pm$ 0.1316 0.0429 $\pm$ 0.0018
BBP 111.6410 $\pm$ 8.9230 65.9097 $\pm$ 3.7377 5.1715 $\pm$ 0.2092 0.0388 $\pm$ 0.0020
SGLD 111.1149 $\pm$ 7.8189 75.5816 $\pm$ 4.4419 5.2277 $\pm$ 0.0341 0.0127 $\pm$ 0.0004
Ensembles 104.0890 $\pm$ 9.9660 55.6554 $\pm$ 4.1624 5.0123 $\pm$ 0.1890 0.0438 $\pm$ 0.0020
ChemBERTa Deterministic 97.4000 $\pm$ 8.8093 47.3906 $\pm$ 3.6007 4.5928 $\pm$ 0.1364 0.0402 $\pm$ 0.0021
MC Dropout 97.1763 $\pm$ 9.0430 43.4817 $\pm$ 4.8435 4.3224 $\pm$ 0.3558 0.0458 $\pm$ 0.0017
SWAG 97.5653 $\pm$ 8.4876 49.3942 $\pm$ 4.0995 4.7928 $\pm$ 0.1495 0.0397 $\pm$ 0.0019
BBP 100.8003 $\pm$ 8.9576 53.0419 $\pm$ 4.4522 4.6938 $\pm$ 0.2195 0.0347 $\pm$ 0.0027
SGLD 99.6487 $\pm$ 7.9623 54.8976 $\pm$ 2.4214 4.6314 $\pm$ 0.1117 0.0308 $\pm$ 0.0015
Ensembles 97.5500 $\pm$ 8.6923 47.2016 $\pm$ 3.7146 4.5436 $\pm$ 0.2688 0.0392 $\pm$ 0.0020
GROVER Deterministic 98.7533 $\pm$ 9.8780 42.7787 $\pm$ 4.8753 4.7424 $\pm$ 0.3523 0.0512 $\pm$ 0.0003
MC Dropout 96.6699 $\pm$ 9.6047 42.6411 $\pm$ 4.6449 4.7324 $\pm$ 0.3130 0.0511 $\pm$ 0.0003
SWAG 97.1293 $\pm$ 10.0641 43.3197 $\pm$ 5.1696 4.6304 $\pm$ 0.2678 0.0495 $\pm$ 0.0008
BBP 99.0196 $\pm$ 8.1011 56.8246 $\pm$ 3.9466 4.9841 $\pm$ 0.0630 0.0223 $\pm$ 0.0003
SGLD 101.0303 $\pm$ 6.6971 57.6752 $\pm$ 3.7755 4.9718 $\pm$ 0.0556 0.0237 $\pm$ 0.0009
Ensembles 98.0631 $\pm$ 9.7788 42.8305 $\pm$ 4.6214 4.6874 $\pm$ 0.3069 0.0498 $\pm$ 0.0006
Uni-Mol Deterministic 71.4475 $\pm$ 2.3163 36.7968 $\pm$ 2.3826 4.1470 $\pm$ 0.1635 0.0411 $\pm$ 0.0032
MC Dropout 70.8934 $\pm$ 2.2622 35.5291 $\pm$ 2.4115 4.1425 $\pm$ 0.1709 0.0406 $\pm$ 0.0032
SWAG 68.6536 $\pm$ 1.3320 29.1093 $\pm$ 2.1596 3.9460 $\pm$ 0.0134 0.0433 $\pm$ 0.0011
BBP 73.0523 $\pm$ 1.1094 39.2194 $\pm$ 1.8792 4.5952 $\pm$ 0.0996 0.0297 $\pm$ 0.0042
SGLD 71.2770 $\pm$ 1.6947 32.4820 $\pm$ 3.3687 4.4407 $\pm$ 0.0538 0.0321 $\pm$ 0.0017
Ensembles 69.6610 $\pm$ 0.9264 33.8478 $\pm$ 0.0577 4.0436 $\pm$ 0.0870 0.0429 $\pm$ 0.0020
Table 31: Test results with frozen backbone model weights on BACE.
ROC-AUC ECE NLL BS
ChemBERTa Deterministic 0.6289 $\pm$ 0.0090 0.2096 $\pm$ 0.0037 0.7570 $\pm$ 0.0073 0.2751 $\pm$ 0.0028
Temperature 0.6289 $\pm$ 0.0090 0.1554 $\pm$ 0.0103 0.7134 $\pm$ 0.0032 0.2587 $\pm$ 0.0014
Focal Loss 0.6474 $\pm$ 0.0051 0.1964 $\pm$ 0.0035 0.7450 $\pm$ 0.0023 0.2741 $\pm$ 0.0011
MC Dropout 0.6515 $\pm$ 0.0164 0.1931 $\pm$ 0.0045 0.7393 $\pm$ 0.0029 0.2693 $\pm$ 0.0013
SWAG 0.6305 $\pm$ 0.0081 0.2045 $\pm$ 0.0032 0.7559 $\pm$ 0.0077 0.2746 $\pm$ 0.0030
BBP 0.5454 $\pm$ 0.0619 0.0998 $\pm$ 0.0235 0.7080 $\pm$ 0.0087 0.2573 $\pm$ 0.0042
SGLD 0.5217 $\pm$ 0.0512 0.0531 $\pm$ 0.0114 0.6941 $\pm$ 0.0022 0.2505 $\pm$ 0.0011
Ensembles 0.6308 $\pm$ - 0.1415 $\pm$ - 0.7081 $\pm$ - 0.2569 $\pm$ -
GROVER Deterministic 0.7633 $\pm$ 0.0116 0.2083 $\pm$ 0.0100 0.6934 $\pm$ 0.0070 0.2514 $\pm$ 0.0030
Temperature 0.7635 $\pm$ 0.0117 0.1915 $\pm$ 0.0109 0.6826 $\pm$ 0.0067 0.2461 $\pm$ 0.0029
Focal Loss 0.7221 $\pm$ 0.0343 0.2125 $\pm$ 0.0066 0.7358 $\pm$ 0.0048 0.2710 $\pm$ 0.0020
MC Dropout 0.7672 $\pm$ 0.0081 0.2221 $\pm$ 0.0111 0.7037 $\pm$ 0.0050 0.2561 $\pm$ 0.0020
SWAG 0.7691 $\pm$ 0.0096 0.2232 $\pm$ 0.0131 0.7027 $\pm$ 0.0036 0.2559 $\pm$ 0.0013
BBP 0.7425 $\pm$ 0.0089 0.1946 $\pm$ 0.0135 0.7046 $\pm$ 0.0038 0.2565 $\pm$ 0.0015
SGLD 0.5485 $\pm$ 0.0612 0.0944 $\pm$ 0.0227 0.7007 $\pm$ 0.0135 0.2536 $\pm$ 0.0067
Ensembles 0.7826 $\pm$ - 0.2278 $\pm$ - 0.6807 $\pm$ - 0.2462 $\pm$ -
Uni-Mol Deterministic 0.3713 $\pm$ 0.1533 0.1999 $\pm$ 0.0357 0.7544 $\pm$ 0.0216 0.2795 $\pm$ 0.0101
Temperature 0.3712 $\pm$ 0.1534 0.1811 $\pm$ 0.0340 0.7317 $\pm$ 0.0149 0.2689 $\pm$ 0.0072
Focal Loss 0.3512 $\pm$ 0.1450 0.2720 $\pm$ 0.0414 0.7977 $\pm$ 0.0301 0.2985 $\pm$ 0.0131
MC Dropout 0.3698 $\pm$ 0.1569 0.2052 $\pm$ 0.0513 0.7512 $\pm$ 0.0203 0.2780 $\pm$ 0.0095
SWAG 0.4297 $\pm$ 0.1408 0.1513 $\pm$ 0.0237 0.7330 $\pm$ 0.0089 0.2695 $\pm$ 0.0042
BBP 0.5958 $\pm$ 0.0671 0.1614 $\pm$ 0.0191 0.7361 $\pm$ 0.0096 0.2709 $\pm$ 0.0045
SGLD 0.7189 $\pm$ 0.0291 0.1414 $\pm$ 0.0218 0.6737 $\pm$ 0.0091 0.2405 $\pm$ 0.0045
Ensembles 0.6799 $\pm$ - 0.1543 $\pm$ - 0.7370 $\pm$ - 0.2714 $\pm$ -
Table 32: Test results with frozen backbone model weights on BBBP.
ROC-AUC ECE NLL BS
ChemBERTa Deterministic 0.6651 $\pm$ 0.0075 0.1025 $\pm$ 0.0103 0.6534 $\pm$ 0.0072 0.2308 $\pm$ 0.0033
Temperature 0.6651 $\pm$ 0.0075 0.1681 $\pm$ 0.0189 0.7366 $\pm$ 0.0315 0.2518 $\pm$ 0.0073
Focal Loss 0.6600 $\pm$ 0.0082 0.1342 $\pm$ 0.0038 0.6782 $\pm$ 0.0017 0.2430 $\pm$ 0.0009
MC Dropout 0.6592 $\pm$ 0.0084 0.1358 $\pm$ 0.0159 0.6813 $\pm$ 0.0058 0.2428 $\pm$ 0.0030
SWAG 0.6650 $\pm$ 0.0079 0.1068 $\pm$ 0.0112 0.6537 $\pm$ 0.0074 0.2310 $\pm$ 0.0033
BBP 0.6280 $\pm$ 0.0074 0.1684 $\pm$ 0.0089 0.7250 $\pm$ 0.0082 0.2645 $\pm$ 0.0035
SGLD 0.4822 $\pm$ 0.0247 0.0346 $\pm$ 0.0091 0.6945 $\pm$ 0.0020 0.2507 $\pm$ 0.0010
Ensembles 0.6545 $\pm$ - 0.1059 $\pm$ - 0.6539 $\pm$ - 0.2320 $\pm$ -
GROVER Deterministic 0.6088 $\pm$ 0.0124 0.3048 $\pm$ 0.0071 0.9424 $\pm$ 0.0027 0.3303 $\pm$ 0.0024
Temperature 0.6083 $\pm$ 0.0122 0.2738 $\pm$ 0.0073 0.8674 $\pm$ 0.0022 0.3119 $\pm$ 0.0019
Focal Loss 0.5979 $\pm$ 0.0167 0.0867 $\pm$ 0.0234 0.6872 $\pm$ 0.0071 0.2469 $\pm$ 0.0034
MC Dropout 0.6089 $\pm$ 0.0135 0.3073 $\pm$ 0.0114 0.9372 $\pm$ 0.0028 0.3297 $\pm$ 0.0021
SWAG 0.6179 $\pm$ 0.0112 0.3087 $\pm$ 0.0038 0.9424 $\pm$ 0.0042 0.3300 $\pm$ 0.0022
BBP 0.6096 $\pm$ 0.0069 0.3023 $\pm$ 0.0065 0.9211 $\pm$ 0.0038 0.3271 $\pm$ 0.0012
SGLD 0.4628 $\pm$ 0.0682 0.0685 $\pm$ 0.0306 0.6979 $\pm$ 0.0108 0.2523 $\pm$ 0.0053
Ensembles 0.6107 $\pm$ - 0.2982 $\pm$ - 0.9338 $\pm$ - 0.3285 $\pm$ -
Uni-Mol Deterministic 0.5325 $\pm$ 0.0364 0.3534 $\pm$ 0.0074 1.0741 $\pm$ 0.0298 0.3726 $\pm$ 0.0064
Temperature 0.5326 $\pm$ 0.0366 0.2867 $\pm$ 0.0074 0.9013 $\pm$ 0.0181 0.3297 $\pm$ 0.0056
Focal Loss 0.5405 $\pm$ 0.0458 0.0522 $\pm$ 0.0042 0.6912 $\pm$ 0.0075 0.2490 $\pm$ 0.0037
MC Dropout 0.5336 $\pm$ 0.0397 0.3449 $\pm$ 0.0077 1.0457 $\pm$ 0.0290 0.3666 $\pm$ 0.0066
SWAG 0.5592 $\pm$ 0.0143 0.3008 $\pm$ 0.0042 0.9277 $\pm$ 0.0109 0.3374 $\pm$ 0.0033
BBP 0.4626 $\pm$ 0.0239 0.2631 $\pm$ 0.0447 0.8784 $\pm$ 0.0809 0.3209 $\pm$ 0.0243
SGLD 0.4590 $\pm$ 0.0456 0.1206 $\pm$ 0.0298 0.7238 $\pm$ 0.0136 0.2646 $\pm$ 0.0059
Ensembles 0.5875 $\pm$ - 0.3370 $\pm$ - 1.0186 $\pm$ - 0.3614 $\pm$ -
Table 33: Test results with frozen backbone model weights on ClinTox.
ROC-AUC ECE NLL BS
ChemBERTa Deterministic 0.6344 $\pm$ 0.2837 0.2913 $\pm$ 0.1651 0.4713 $\pm$ 0.2530 0.1618 $\pm$ 0.0952
Temperature 0.6344 $\pm$ 0.2837 0.2389 $\pm$ 0.1655 0.4157 $\pm$ 0.2533 0.1398 $\pm$ 0.0964
Focal Loss 0.6369 $\pm$ 0.2860 0.3267 $\pm$ 0.1183 0.5099 $\pm$ 0.2000 0.1688 $\pm$ 0.0859
MC Dropout 0.6689 $\pm$ 0.2643 0.2886 $\pm$ 0.1675 0.4706 $\pm$ 0.2491 0.1612 $\pm$ 0.0934
SWAG 0.6911 $\pm$ 0.2300 0.2245 $\pm$ 0.1211 0.3728 $\pm$ 0.1831 0.1142 $\pm$ 0.0607
BBP 0.7242 $\pm$ 0.0410 0.2659 $\pm$ 0.1039 0.4515 $\pm$ 0.1399 0.1379 $\pm$ 0.0638
SGLD 0.5489 $\pm$ 0.1696 0.4339 $\pm$ 0.0032 0.6898 $\pm$ 0.0048 0.2483 $\pm$ 0.0024
Ensembles 0.9734 $\pm$ - 0.2885 $\pm$ - 0.4085 $\pm$ - 0.1139 $\pm$ -
GROVER Deterministic 0.6437 $\pm$ 0.0235 0.0304 $\pm$ 0.0051 0.2357 $\pm$ 0.0041 0.0599 $\pm$ 0.0008
Temperature 0.6426 $\pm$ 0.0240 0.0179 $\pm$ 0.0032 0.2337 $\pm$ 0.0050 0.0594 $\pm$ 0.0007
Focal Loss 0.6328 $\pm$ 0.0220 0.2116 $\pm$ 0.0054 0.3897 $\pm$ 0.0064 0.1109 $\pm$ 0.0027
MC Dropout 0.6534 $\pm$ 0.0201 0.0222 $\pm$ 0.0041 0.2329 $\pm$ 0.0029 0.0594 $\pm$ 0.0005
SWAG 0.6838 $\pm$ 0.0124 0.0268 $\pm$ 0.0064 0.2306 $\pm$ 0.0031 0.0589 $\pm$ 0.0007
BBP 0.6019 $\pm$ 0.0346 0.0339 $\pm$ 0.0028 0.2422 $\pm$ 0.0032 0.0609 $\pm$ 0.0005
SGLD 0.4601 $\pm$ 0.0878 0.4365 $\pm$ 0.0355 0.7050 $\pm$ 0.0650 0.2558 $\pm$ 0.0321
Ensembles 0.7039 $\pm$ - 0.0348 $\pm$ - 0.2311 $\pm$ - 0.0591 $\pm$ -
Uni-Mol Deterministic 0.5750 $\pm$ 0.0185 0.1620 $\pm$ 0.0598 0.3423 $\pm$ 0.0569 0.0916 $\pm$ 0.0209
Temperature 0.5752 $\pm$ 0.0182 0.0713 $\pm$ 0.0499 0.2694 $\pm$ 0.0319 0.0678 $\pm$ 0.0086
Focal Loss 0.5696 $\pm$ 0.0199 0.2605 $\pm$ 0.0328 0.4397 $\pm$ 0.0359 0.1290 $\pm$ 0.0158
MC Dropout 0.5797 $\pm$ 0.0259 0.1704 $\pm$ 0.0595 0.3499 $\pm$ 0.0577 0.0944 $\pm$ 0.0216
SWAG 0.5887 $\pm$ 0.0228 0.1125 $\pm$ 0.0324 0.2948 $\pm$ 0.0250 0.0741 $\pm$ 0.0075
BBP 0.5899 $\pm$ 0.0112 0.2754 $\pm$ 0.0510 0.4608 $\pm$ 0.0683 0.1391 $\pm$ 0.0313
SGLD 0.5257 $\pm$ 0.0576 0.4285 $\pm$ 0.0442 0.7081 $\pm$ 0.0984 0.2556 $\pm$ 0.0463
Ensembles 0.5551 $\pm$ - 0.1449 $\pm$ - 0.3199 $\pm$ - 0.0811 $\pm$ -
Table 34: Test results with frozen backbone model weights on Tox21.
ROC-AUC ECE NLL BS
ChemBERTa Deterministic 0.6615 $\pm$ 0.0039 0.0402 $\pm$ 0.0008 0.2934 $\pm$ 0.0004 0.0826 $\pm$ 0.0002
Temperature 0.6615 $\pm$ 0.0039 0.0307 $\pm$ 0.0015 0.2894 $\pm$ 0.0008 0.0817 $\pm$ 0.0001
Focal Loss 0.6646 $\pm$ 0.0026 0.1183 $\pm$ 0.0025 0.3428 $\pm$ 0.0021 0.0949 $\pm$ 0.0007
MC Dropout 0.6729 $\pm$ 0.0048 0.0315 $\pm$ 0.0016 0.2859 $\pm$ 0.0007 0.0807 $\pm$ 0.0002
SWAG 0.6619 $\pm$ 0.0038 0.0397 $\pm$ 0.0012 0.2934 $\pm$ 0.0005 0.0826 $\pm$ 0.0002
BBP 0.5426 $\pm$ 0.0080 0.1150 $\pm$ 0.0079 0.3629 $\pm$ 0.0059 0.1009 $\pm$ 0.0019
SGLD 0.5062 $\pm$ 0.0212 0.3973 $\pm$ 0.0005 0.6903 $\pm$ 0.0008 0.2486 $\pm$ 0.0004
Ensembles 0.6663 $\pm$ - 0.0383 $\pm$ - 0.2916 $\pm$ - 0.0822 $\pm$ -
GROVER Deterministic 0.6897 $\pm$ 0.0102 0.0361 $\pm$ 0.0002 0.2838 $\pm$ 0.0011 0.0807 $\pm$ 0.0003
Temperature 0.6897 $\pm$ 0.0102 0.0223 $\pm$ 0.0017 0.2789 $\pm$ 0.0011 0.0799 $\pm$ 0.0003
Focal Loss 0.6632 $\pm$ 0.0139 0.1326 $\pm$ 0.0023 0.3593 $\pm$ 0.0005 0.0998 $\pm$ 0.0002
MC Dropout 0.6874 $\pm$ 0.0093 0.0343 $\pm$ 0.0004 0.2842 $\pm$ 0.0009 0.0810 $\pm$ 0.0002
SWAG 0.6923 $\pm$ 0.0075 0.0344 $\pm$ 0.0002 0.2830 $\pm$ 0.0008 0.0807 $\pm$ 0.0002
BBP 0.6496 $\pm$ 0.0040 0.0364 $\pm$ 0.0005 0.2951 $\pm$ 0.0003 0.0836 $\pm$ 0.0001
SGLD 0.5072 $\pm$ 0.0266 0.3898 $\pm$ 0.0018 0.6806 $\pm$ 0.0024 0.2438 $\pm$ 0.0012
Ensembles 0.6911 $\pm$ - 0.0365 $\pm$ - 0.2839 $\pm$ - 0.0807 $\pm$ -
Uni-Mol Deterministic 0.5702 $\pm$ 0.0065 0.0556 $\pm$ 0.0008 0.3352 $\pm$ 0.0009 0.0905 $\pm$ 0.0002
Temperature 0.5702 $\pm$ 0.0066 0.0153 $\pm$ 0.0012 0.3036 $\pm$ 0.0001 0.0862 $\pm$ 0.0000
Focal Loss 0.5620 $\pm$ 0.0067 0.0472 $\pm$ 0.0030 0.3183 $\pm$ 0.0020 0.0888 $\pm$ 0.0004
MC Dropout 0.5696 $\pm$ 0.0056 0.0519 $\pm$ 0.0004 0.3297 $\pm$ 0.0007 $0$$\pm$ 0.0002
SWAG 0.5936 $\pm$ 0.0056 0.0293 $\pm$ 0.0011 0.3070 $\pm$ 0.0005 0.0868 $\pm$ 0.0001
BBP 0.5505 $\pm$ 0.0184 0.0472 $\pm$ 0.0021 0.3249 $\pm$ 0.0021 0.0896 $\pm$ 0.0002
SGLD 0.4927 $\pm$ 0.0089 0.4022 $\pm$ 0.0118 0.7135 $\pm$ 0.0216 0.2599 $\pm$ 0.0104
Ensembles 0.5866 $\pm$ - 0.0546 $\pm$ - 0.3335 $\pm$ - 0.0901 $\pm$ -
Table 35: Test results with frozen backbone model weights on ESOL.
RMSE MAE NLL CE
ChemBERTa Deterministic 1.9543 $\pm$ 0.0830 1.4348 $\pm$ 0.0895 1.2260 $\pm$ 0.0454 0.0090 $\pm$ 0.0010
MC Dropout 1.8954 $\pm$ 0.1132 1.4051 $\pm$ 0.1047 1.1497 $\pm$ 0.0815 0.0080 $\pm$ 0.0006
SWAG 1.9708 $\pm$ 0.0865 1.4372 $\pm$ 0.0905 1.2561 $\pm$ 0.0495 0.0105 $\pm$ 0.0011
BBP 2.2450 $\pm$ 0.0440 1.7152 $\pm$ 0.0346 1.3560 $\pm$ 0.0285 0.0155 $\pm$ 0.0026
SGLD 2.3491 $\pm$ 0.0542 1.7899 $\pm$ 0.0472 1.4684 $\pm$ 0.0373 0.0180 $\pm$ 0.0055
Ensembles 1.8844 $\pm$ - 1.3699 $\pm$ - 1.1914 $\pm$ - 0.0082 $\pm$ -
GROVER Deterministic 1.8533 $\pm$ 0.1590 1.4171 $\pm$ 0.1281 1.1335 $\pm$ 0.0833 0.0097 $\pm$ 0.0024
MC Dropout 1.8629 $\pm$ 0.1562 1.4131 $\pm$ 0.1234 1.1444 $\pm$ 0.0836 0.0110 $\pm$ 0.0031
SWAG 1.8059 $\pm$ 0.1521 1.3651 $\pm$ 0.1102 1.0930 $\pm$ 0.0857 0.0120 $\pm$ 0.0029
BBP 1.9032 $\pm$ 0.1318 1.4502 $\pm$ 0.1186 1.3243 $\pm$ 0.0225 0.0088 $\pm$ 0.0014
SGLD 2.4699 $\pm$ 0.1371 1.8858 $\pm$ 0.1161 1.3697 $\pm$ 0.0510 0.0285 $\pm$ 0.0050
Ensembles 1.8185 $\pm$ - 1.3161 $\pm$ - 1.1063 $\pm$ - 0.0098 $\pm$ -
Uni-Mol Deterministic 2.2126 $\pm$ 0.0258 1.7134 $\pm$ 0.0482 1.3168 $\pm$ 0.0112 0.0078 $\pm$ 0.0054
MC Dropout 2.2155 $\pm$ 0.0223 1.7118 $\pm$ 0.0426 1.3169 $\pm$ 0.0106 0.0078 $\pm$ 0.0053
SWAG 2.1809 $\pm$ 0.0258 1.6573 $\pm$ 0.0294 1.2883 $\pm$ 0.0210 0.0085 $\pm$ 0.0035
BBP 2.1964 $\pm$ 0.1226 1.7223 $\pm$ 0.0996 1.3008 $\pm$ 0.0625 0.0053 $\pm$ 0.0042
SGLD 2.3596 $\pm$ 0.0599 1.8831 $\pm$ 0.0734 1.4158 $\pm$ 0.0808 0.0139 $\pm$ 0.0037
Ensembles 2.1476 $\pm$ - 1.6524 $\pm$ - 1.2669 $\pm$ - 0.0053 $\pm$ -
Table 36: Test results with frozen backbone model weights on FreeSolv.
RMSE MAE NLL CE
ChemBERTa Deterministic 3.7564 $\pm$ 0.0520 2.7933 $\pm$ 0.0770 2.0303 $\pm$ 0.0828 0.0378 $\pm$ 0.0078
MC Dropout 3.7861 $\pm$ 0.0542 2.8493 $\pm$ 0.0801 2.0393 $\pm$ 0.0773 0.0429 $\pm$ 0.0065
SWAG 3.7629 $\pm$ 0.0297 2.8017 $\pm$ 0.0563 2.0520 $\pm$ 0.0679 0.0395 $\pm$ 0.0062
BBP 4.5822 $\pm$ 0.0444 3.4921 $\pm$ 0.0636 2.3843 $\pm$ 0.0507 0.0669 $\pm$ 0.0037
SGLD 4.3996 $\pm$ 0.1248 3.3132 $\pm$ 0.1317 2.3018 $\pm$ 0.1113 0.0546 $\pm$ 0.0088
Ensembles 3.7094 $\pm$ - 2.7309 $\pm$ - 2.0012 $\pm$ - 0.0370 $\pm$ -
GROVER Deterministic 4.2407 $\pm$ 0.3174 3.1254 $\pm$ 0.4023 2.0701 $\pm$ 0.1023 0.0233 $\pm$ 0.0141
MC Dropout 4.2045 $\pm$ 0.3199 3.1098 $\pm$ 0.3737 2.0422 $\pm$ 0.0929 0.0199 $\pm$ 0.0130
SWAG 4.1854 $\pm$ 0.2899 3.1039 $\pm$ 0.3888 2.0604 $\pm$ 0.1230 0.0251 $\pm$ 0.0127
BBP 4.4898 $\pm$ 0.2067 $3$$\pm$ 0.2186 2.0452 $\pm$ 0.0530 0.0378 $\pm$ 0.0136
SGLD 4.6731 $\pm$ 0.5142 3.5858 $\pm$ 0.5165 2.1826 $\pm$ 0.0714 0.0529 $\pm$ 0.0178
Ensembles 4.0551 $\pm$ - 3.0166 $\pm$ - 1.9352 $\pm$ - 0.0245 $\pm$ -
Uni-Mol Deterministic 3.9034 $\pm$ 0.4092 2.8791 $\pm$ 0.3542 1.8779 $\pm$ 0.0965 0.0207 $\pm$ 0.0150
MC Dropout 3.9103 $\pm$ 0.4142 2.8836 $\pm$ 0.3612 1.8813 $\pm$ 0.0990 0.0209 $\pm$ 0.0151
SWAG 4.1746 $\pm$ 0.2749 3.1272 $\pm$ 0.2718 1.9582 $\pm$ 0.0786 0.0360 $\pm$ 0.0128
BBP 4.0001 $\pm$ 0.2701 3.0110 $\pm$ 0.2507 1.9486 $\pm$ 0.1219 0.0276 $\pm$ 0.0196
SGLD 4.1575 $\pm$ 0.0222 3.1164 $\pm$ 0.0627 2.0574 $\pm$ 0.1045 0.0321 $\pm$ 0.0066
Ensembles 4.0416 $\pm$ - 2.9823 $\pm$ - 1.9718 $\pm$ - 0.0287 $\pm$ -
Table 37: Test results with frozen backbone model weights on Lipophilicity.
RMSE MAE NLL CE
ChemBERTa Deterministic 0.9958 $\pm$ 0.0033 0.8083 $\pm$ 0.0047 0.4837 $\pm$ 0.0039 0.0016 $\pm$ 0.0001
MC Dropout 0.9985 $\pm$ 0.0064 0.8185 $\pm$ 0.0061 0.4885 $\pm$ 0.0063 0.0041 $\pm$ 0.0001
SWAG 0.9948 $\pm$ 0.0029 0.8070 $\pm$ 0.0043 0.4823 $\pm$ 0.0031 0.0016 $\pm$ 0.0001
BBP 1.0477 $\pm$ 0.0017 0.8578 $\pm$ 0.0018 0.5537 $\pm$ 0.0033 0.0025 $\pm$ 0.0001
SGLD 1.1189 $\pm$ 0.0033 0.9266 $\pm$ 0.0036 0.6225 $\pm$ 0.0038 0.0074 $\pm$ 0.0012
Ensembles 0.9837 $\pm$ - 0.7981 $\pm$ - 0.4729 $\pm$ - 0.0016 $\pm$ -
GROVER Deterministic 0.9297 $\pm$ 0.0182 0.7568 $\pm$ 0.0165 0.4238 $\pm$ 0.0178 0.0019 $\pm$ 0.0002
MC Dropout 0.9366 $\pm$ 0.0146 0.7683 $\pm$ 0.0113 0.4348 $\pm$ 0.0148 0.0033 $\pm$ 0.0004
SWAG 0.9290 $\pm$ 0.0148 0.7621 $\pm$ 0.0104 0.4223 $\pm$ 0.0144 0.0034 $\pm$ 0.0003
BBP 0.9753 $\pm$ 0.0229 0.7950 $\pm$ 0.0192 0.6023 $\pm$ 0.0346 0.0025 $\pm$ 0.0007
SGLD 1.1203 $\pm$ 0.0238 0.9301 $\pm$ 0.0243 0.6134 $\pm$ 0.0206 0.0052 $\pm$ 0.0031
Ensembles 0.9160 $\pm$ - 0.7455 $\pm$ - 0.4134 $\pm$ - 0.0019 $\pm$ -
Uni-Mol Deterministic 1.0917 $\pm$ 0.0210 0.9069 $\pm$ 0.0203 0.6446 $\pm$ 0.0189 0.0055 $\pm$ 0.0014
MC Dropout 1.0929 $\pm$ 0.0216 0.9080 $\pm$ 0.0208 0.6442 $\pm$ 0.0192 0.0054 $\pm$ 0.0014
SWAG 1.0822 $\pm$ 0.0114 0.8986 $\pm$ 0.0115 0.6144 $\pm$ 0.0121 0.0045 $\pm$ 0.0004
BBP 1.1003 $\pm$ 0.0198 0.9112 $\pm$ 0.0160 0.8454 $\pm$ 0.0580 0.0124 $\pm$ 0.0022
SGLD 1.1796 $\pm$ 0.0644 0.9567 $\pm$ 0.0190 0.6812 $\pm$ 0.0469 0.0135 $\pm$ 0.0051
Ensembles 1.0863 $\pm$ - 0.9019 $\pm$ - 0.6421 $\pm$ - 0.0052 $\pm$ -
Table 38: Test results with frozen backbone model weights on QM7.
RMSE MAE NLL CE
ChemBERTa Deterministic 137.8736 $\pm$ 2.2380 98.5749 $\pm$ 2.9618 5.4092 $\pm$ 0.0182 0.0049 $\pm$ 0.0001
MC Dropout 133.8345 $\pm$ 1.8327 97.7798 $\pm$ 2.3986 5.4049 $\pm$ 0.0117 0.0028 $\pm$ 0.0001
SWAG 138.2471 $\pm$ 2.2020 99.0168 $\pm$ 2.7806 5.4107 $\pm$ 0.0177 0.0048 $\pm$ 0.0001
BBP 141.0033 $\pm$ 1.2990 104.8384 $\pm$ 1.5366 5.4635 $\pm$ 0.0100 0.0026 $\pm$ 0.0005
SGLD 202.5163 $\pm$ 1.4235 165.5756 $\pm$ 1.1589 5.8131 $\pm$ 0.0087 0.0208 $\pm$ 0.0015
Ensembles 136.6747 $\pm$ - 97.2405 $\pm$ - 5.4066 $\pm$ - 0.0045 $\pm$ -
GROVER Deterministic 134.4406 $\pm$ 3.7901 95.5997 $\pm$ 4.3313 5.4077 $\pm$ 0.0717 0.0058 $\pm$ 0.0023
MC Dropout 131.7638 $\pm$ 3.5591 93.2436 $\pm$ 4.3578 5.4005 $\pm$ 0.0761 0.0042 $\pm$ 0.0021
SWAG 130.2656 $\pm$ 2.2246 92.1086 $\pm$ 3.1537 5.3670 $\pm$ 0.0446 0.0047 $\pm$ 0.0015
BBP 137.9643 $\pm$ 6.4603 100.5829 $\pm$ 7.7898 5.8332 $\pm$ 0.1322 0.0055 $\pm$ 0.0030
SGLD 191.3884 $\pm$ 3.4342 154.4454 $\pm$ 1.8662 5.8488 $\pm$ 0.0227 0.0158 $\pm$ 0.0033
Ensembles 130.9773 $\pm$ - 93.3082 $\pm$ - 5.4027 $\pm$ - 0.0051 $\pm$ -
Uni-Mol Deterministic 183.5096 $\pm$ 9.5881 148.3822 $\pm$ 9.8276 5.7908 $\pm$ 0.0353 0.0180 $\pm$ 0.0051
MC Dropout 183.7110 $\pm$ 9.1987 148.5660 $\pm$ 9.3875 5.7908 $\pm$ 0.0336 0.0178 $\pm$ 0.0048
SWAG 171.6691 $\pm$ 1.7586 137.1572 $\pm$ 2.1770 5.7147 $\pm$ 0.0078 0.0111 $\pm$ 0.0008
BBP 178.0057 $\pm$ 3.9910 142.9159 $\pm$ 3.6271 6.0873 $\pm$ 0.0082 0.0193 $\pm$ 0.0007
SGLD 226.7307 $\pm$ 21.2619 188.4545 $\pm$ 19.4483 6.0090 $\pm$ 0.1956 0.0458 $\pm$ 0.0322
Ensembles 175.6095 $\pm$ - 140.6629 $\pm$ - 5.7658 $\pm$ - 0.0129 $\pm$ -